### v2

Abstract
Over 40% of computational power in Large Language Model (LLM) serving systems can be systematically wasted—not from hardware limits, but from load imbalance in barrier-synchronized parallel processing. When progress is gated by the slowest worker at each step, heterogeneous and evolving workloads create persistent stragglers; faster workers idle while drawing power, producing nothing. In large language model inference alone, this translates to gigawatt-hours of wasted electricity daily. Here we develop a universal load-balancing principle for barrier-synchronized systems with non-migratable state. We prove worst-case theoretical guarantees: imbalance reduction grows with system scale, and the resulting energy savings can exceed 52% for modern hardware at fleet scale. Experiments corroborate the theory, demonstrating 28% energy reduction alongside substantial throughput and latency improvements. Formulated as an online integer optimization with provable guarantees, the principle extends beyond LLM serving to broad classes of barrier-synchronized parallel systems, establishing a theoretical foundation for sustainable high-performance computing.
1 Introduction
Load balancing—the problem of distributing work across parallel resources to minimize latency, energy, or cost—is a central challenge across the natural and engineering sciences and in large-scale service systems. In large-scale molecular dynamics and biomolecular simulations, dynamic load-balancing schemes are essential to maintain scalability as particles move and local density fluctuates [hess2008gromacs4, deng2000adaptive]. In climate and weather models, specialized load-balancing strategies are required to handle localized, high-intensity physics kernels and heterogeneous meshes, improving time-to-solution for global simulations and regional storm-surge predictions [foster1994load, rodrigues2010comparative]. Similar issues arise in cloud computing platforms [mishra2020load, khiyaita2012load], manufacturing systems [chen1987task, li2017big], and emergency medical management [kolomvatsos2015load, friesen2011load], where workloads and resource availability fluctuate over time. In this work, we study a new instance of this broad class: load balancing in large-scale LLM (Large Language Model) serving. This setting adds stringent constraints—jobs whose processing times are unknown and evolve as computation proceeds, per-request state that cannot be migrated between workers without prohibitive cost, and step-wise synchronization barriers across workers—that make it substantially more challenging than many classical load-balancing problems. Our goal is to develop a universal and scientifically sound load balancing principle that not only resolves this emerging bottleneck in LLM serving, but also applies, in appropriately simplified form, to more traditional load-balancing tasks.
Modern generative LLMs [brown2020language, chowdhery2023palm, openai2023gpt, kaplan2020scaling, wei2022emergent] have transformed artificial intelligence by achieving unprecedented performance in natural-language generation, reasoning, and task generalization across diverse domains. LLM serving refers to the process of hosting these models in production environments to process user prompts and generate responses in real time. Meeting this demand requires vast networks of AI datacenters that deliver timely and reliable inference at massive scale [kim2025inquiry, Milmo2025]. Even considering OpenAI alone, its LLM serving services handle at least $2.5$ billion user prompts per day [chatgptdemand], corresponding to roughly on the order of gigawatt-hour of electricity consumption daily [chatgptpower]. As LLM usage continues to accelerate, the energy and cooling resources required for such inference are approaching practical and environmental limits—often referred to as the emerging energy wall [energywall]. Despite remarkable advances in model architecture and hardware acceleration, the overall efficiency and energy use of LLM serving remains constrained by how computational workloads are balanced across parallel workers.
In LLM serving, inference unfolds in two stages: a prefill stage that encodes the input prompt and initializes internal state, followed by a decode stage that generates the response one token at a time. In modern deployments, decoding is executed in parallel across many computational workers, and each token step typically includes a coordinated stage across workers (e.g., collective communication in model-parallel execution), which induces barrier synchronization: the step completes when the slowest worker finishes. As active requests evolve and complete at different times, per-worker workloads drift, creating persistent stragglers and forcing faster workers to idle. Figure 1 quantifies the resulting inefficiency in a real industrial trace: both the mean and median barrier-induced idle time exceed 40% per decode step, wasting more than two-fifths of aggregate compute on average. This inefficiency scales directly into energy cost at fleet scale. Because GPUs draw substantial power even while waiting at synchronization barriers—idle time consumes energy [ozcan2025quantifying] without producing useful computation—the 40% barrier-induced idle in Figure 1 represents not merely wasted time but wasted watts. Figure 2 demonstrates this concretely: reducing decode-stage imbalance via the approach introduced in this paper yields a 28.2% reduction in total energy consumption relative to the default serving policy over this trace, with the improvement growing as system scale increases. At the scale of modern LLM deployments—where inference alone can consume gigawatt-hours of electricity daily—such reductions translate to megawatt-hour savings with proportionate reductions in cooling load and carbon footprint. We next unpack the decode-side mechanics that create this imbalance and clarify why routing decisions are both critical and difficult.
Figure 1: Workload imbalance and idle time in 436436 decoding steps over 33 minutes from a real industrial trace. (Left): Dynamic total workload across 32 GPUs. At each step, the completion time for all GPUs is dictated by the heaviest workload unit, demonstrating significant and persistent workload imbalance. (Right): Per-step idle time resulting from this imbalance. The yellow bars represent the average percentage of time that GPUs spend waiting during each decode step. With a mean (and median) idle time of 40% (and 41%), more than two-fifths of the aggregate computational resources are wasted on average per step.
Figure 2: Comparison of energy and power consumption over the default policy and the approach introduced in this paper (BF-IO). (Left): Instantaneous GPU power consumption over time under the default policy (orange curve) and BF-IO (blue curve). Total energy, computed as the time integral of instantaneous power, is 29.1 MJ for the default and 20.9 MJ for BF-IO —a 28.2% reduction. BF-IO draws slightly higher instantaneous power (reflecting improved GPU utilization from reduced barrier idling), but completes the same workload in less time; the resulting shorter integration window yields lower total energy consumption. (Right): Energy consumption over increasing system scale (number of GPUs). The energy reduction percentage keeps increasing when the system becomes larger.
To make the bottleneck precise, we outline the decode-side mechanics relevant to load balance. Prefill–decode (PD) disaggregation is the most common structure in LLM serving, where prompt encoding (prefill) and token-by-token generation (decode) run on distinct workers. When a request finishes prefill, the router assigns it to a decode worker; from that handoff onward the request remains on that decode worker. During decoding, each newly generated token appends one unit of key–value (KV) vectors to the per-request KV cache, so the cache grows linearly with response length. In practice, the decode stage is organized via data parallelism (DP): each worker maintains a batch of active requests and their KV caches; migrating a request would require transferring its entire cache and is therefore impractical—assignments are effectively sticky, making the assignment decision crucial. At each decode step $t$, worker $g$ performs the local attention compute for its request batch; the runtime $T_{\mathrm{local}}^{(g)}(t)$ is approximately linear with the aggregate resident KV it must read and carry over. Once the local computation finishes, the system executes a model-parallel stage—such as expert parallelism (EP) or tensor parallelism (TP)—in which the forward pass is split across workers because the model (or its expert set) is partitioned across devices. This stage requires collective communication and is therefore synchronous across workers, taking time $T_{\mathrm{sync}}(t)$. The per-step wall-clock time is thus
$T_{\mathrm{step}}(t)=\max_{g}T_{\mathrm{local}}^{(g)}(t)+T_{\mathrm{sync}}(t).$
As response lengths diverge across requests, these local times diverge, so the maximization term governs progress and induces persistent idle time for lightly loaded workers. Figure 3 illustrates the LLM serving system architecture and the load balancing barrier.
Figure 3: Prefill–decode serving and the decode-stage load-balancing barrier. Left (Decode Step t): Prefill workers build each request’s KV cache and hand the request to the scheduler, which assigns it to a decode worker; the assignment is sticky because migrating the KV cache is impractical. Each decode worker processes a batch of requests during the local compute stage (Multi-head Latent Attention). Gray squares indicate the resident KV cache per request. After local compute, all workers must begin the EP stage synchronously. The step time is therefore governed by the slowest local stage: here, Worker 2 must wait for Worker 1, exposing the load-balancing bottleneck. Right (Decode Step t+1): Every active request produces one token, appending to its KV cache (red square). Completed requests leave the batch (e.g., Request 4), and newly prefetched requests may be admitted by the scheduler to a worker (e.g., Request 7 with green square). Because request lengths differ and assignments are sticky, per-worker loads drift from step to step, creating persistent imbalance that limits throughput and energy efficiency. Legend: gray = existing KV cache; red = KV added by the current step; green = newly admitted request after prefill.
Balancing decode-time work is intrinsically difficult for three coupled reasons. (i) Unknown and evolving workloads. The response length of a request is unknown at assignment. Although several studies predict decode length from prefill features [zheng2023response, qiu2024efficient, fu2024efficient, chen2024kvdirect, shahout2024don], the inherent randomness of generation and the limited information in these features result in low prediction accuracy. (ii) Online, non-stationary arrivals. Requests finish prefill and arrive to the router sequentially, so decisions must be made without knowledge of the future mix of lengths. Moreover, the frequency of arrival varies over time and is hard to model, so even an arrival distribution for decode is unavailable. (iii) Sticky assignments with drifting load. Once placed, a request cannot be migrated without moving its entire KV cache. Each request’s decode is processed on a single decode worker for its entire lifetime. Meanwhile, per-worker load changes at every step (Figure 3): continuing requests add one token’s workload, completed requests remove their entire workload, and newly admitted requests add their KV cache from the prefill step. Early placement errors therefore persist and compound.
Together, these constraints define an online routing problem with unknown, time-varying job sizes under hard synchronization barriers, in which classical policies such as round-robin or join-shortest-queue can be systematically misaligned with the true objective and perform poorly. Because the barrier-induced bottleneck is architectural and the workload mix is inherently non-stationary, improvements that depend on a particular trace or dataset are insufficient; what is needed is a universal load-balancing principle that is robust across deployments and adversarial arrival patterns. Appendix A.1 reviews related work, including the limitations of classical load-balancing policies in this setting.
2 Results
We summarize the main findings of this work in two parts. First, we introduce a universal load-balancing principle for sticky assignment systems with barrier synchronization and show that it admits a clean optimization formulation and rigorous worst-case guarantees. Second, we instantiate the principle in the decode-stage DP bottleneck of LLM serving, design a sustainable routing policy, and demonstrate substantial improvements on both public and proprietary workloads.
2.1 A universal load-balancing principle: Balance a Short Future
Scientific structure: a step-wise integer-optimization formulation.
At each assignment step, the Balance Future with Integer Optimization principle (BF-IO) chooses assignments that minimize a short-horizon prediction of future imbalance (Section 4). We write this decision as a finite-horizon integer optimization over binary assignment variables, subject to per-worker capacity constraints and full utilization of available slots, in the spirit of classical linear- [luenberger1984linear, dantzig2002linear] and integer-optimization [schrijver1998theory] formulations used in scientific and engineering applications. Concretely, the step-$k$ decision is characterized by the integer program (IO), whose objective is the accumulated predicted imbalance over a short window and whose feasible set encodes the sticky assignment rule and slot constraints. This formulation makes explicit that decode-time load balancing is not merely a heuristic dispatch problem: it is an online sequence of structured combinatorial optimization problems with a barrier-driven objective.
A counterintuitive information suggested: short lookahead, not full-job prediction.
A central insight is that effective load balancing in barrier-synchronized systems does not require accurate prediction of the total remaining workload of newly arriving jobs—a task that is notoriously unreliable in LLM inference and many other domains. Instead, our BF-IO principle relies on a weaker and more robust signal: a short lookahead description of the near-future evolution of the currently active jobs, not the newly arriving jobs. Intuitively, the barrier cost at time $k$ is dominated by which workers are likely to become stragglers in the next few steps; decisions can be revisited at subsequent steps as new arrivals are observed, so errors about the distant future can be corrected by future routing choices. This is precisely the regime where short-horizon prediction is feasible: while forecasting the full completion time of a long job may be difficult, predicting whether an ongoing job will finish within a small window is often considerably easier (and can be performed using lightweight signals, or even manual rules in some settings).
2.2 Provable theoretical guarantees under worst-case settings
Guaranteed load balancing improvement in the LLM serving model.
We build a mathematical model in Section 3 for the LLM DP decode load imbalance problem, where each active request contributes an increasing KV workload until termination. Then, we provide worst-case theoretical guarantees, where arrivals may be chosen adversarially in Section 5. We rigorously prove that even under this worst-case arrival model, BF-IO provably reduces long-run imbalance relative to the real baseline algorithm by a factor $\Omega\!\bigl(\sqrt{B\log G}\bigr)$ that grows with system scale (Theorem 1, Theorem 2), where $B$ is the per-worker batch size and $G$ is the number of workers. This scaling is significant: it shows that the benefit of principled balancing increases with cluster size and batching, precisely the regime in which industrial huge serving systems operate.
From imbalance reduction to energy savings.
Beyond efficiency metrics, we establish a rigorous theoretical connection between load balancing and energy consumption. Theorem 4 proves that any guaranteed improvement in imbalance translates into an explicit guaranteed improvement in total synchronized-phase energy, with constants depending on GPU power model parameters and request distribution. Instantiating this result with our imbalance guarantees for BF-IO yields Corollary 1: as system scale $G\to\infty$, the energy saving fraction approaches a hardware-dependent constant that exceeds 52% for modern GPUs. This provides, to our knowledge, the first worst-case theoretical guarantee linking scheduling policy to energy efficiency in LLM serving, establishing a principled foundation for sustainable inference at scale.
Universality beyond LLMs: a broad class of drifting workloads.
The LLM analysis exploits the fact that per-request workload drifts by one unit during processing, but the same phenomenon occurs far more broadly. We therefore extend the theory to a general non-decreasing drift model in which all active jobs share a bounded, time-varying workload increment sequence $(\delta_{k})_{k\geq 1}$, encompassing the classical constant-workload case ($\delta_{k}\equiv 0$), standard LLM decoding with unit-step KV growth ($\delta_{k}\equiv 1$), speculative decoding where multiple tokens may be accepted per step ($\delta_{k}\geq 1$) and and memory-efficient architectures employing cache compression or sparse attention (Section 5.1). Theorem 3 shows that the same scale-sensitive improvement persists in this general setting: BF-IO achieves a worst-case imbalance-reduction factor of order $\Omega(\sqrt{B\log G})$ relative to the baseline under the worst-case arrival sequences.
Beyond its immediate implication for system design, the mathematical proof introduces a technically involved analytical route for studying online load balancing under stickiness and barrier coupling; we view this as a novel and reusable perspective for future theoretical work on load balancing in modern, stateful parallel systems.
2.3 Resolving a decode-stage load balancing bottleneck in LLM serving
Decode-stage load imbalance as a dominant efficiency bottleneck
We establish that decode-stage load imbalance constitutes a dominant efficiency bottleneck in production-scale LLM serving with large-scale EP/TP synchronization. In such deployments, each token step is barrier-gated by the slowest data-parallel worker; heterogeneity in resident KV footprints therefore translates directly into idle time on faster workers while they await synchronization. Our analysis of real industrial traces quantifies this inefficiency: Figure 1 shows that the average fraction of compute time lost per decode step exceeds 40%, representing more than two-fifths of aggregate GPU resources wasted at barriers, which results in a 28.2% energy waste. Crucially, the magnitude of this waste varies sharply with arrival patterns and request characteristics, indicating that improvements tied to any single trace or dataset are insufficient—a finding that motivates our focus on a universal principle with worst-case guarantees.
Empirical results on open and proprietary workloads.
We evaluate BF-IO through extensive numerical simulations to assess whether the theoretical improvements translate into practical efficiency gains. We evaluate BF-IO on (i) public datasets and (ii) proprietary industrial workloads under matched serving configurations (Section 6). Across these settings, BF-IO consistently reduces imbalance and improves system efficiency relative to the default baselines used in real industries. Concretely, on the public workload, BF-IO improves throughput by 90%, reduces latency by 44%, and reduces energy consumption by 28%. The proprietary workload has been verified by a major LLM developer industry in China.
3 Model and Baseline Policy
We consider a general online assignment problem with $G$ computational workers. Each worker $g\in[G]$ has a maximum concurrency of $B\in\mathbb{N}$, meaning that at most $B$ requests can be processed simultaneously on $g$. The discrete time horizon is indexed by $k\in\{0,1,\ldots,K\}$, where $K$ can be taken large (and later set to infinity in asymptotic analyses).
Arrival instance and workload profiles.
Let $\mathcal{I}$ denote the set of all requests in a given arrival instance. Each request $i\in\mathcal{I}$ arrives at some step $k_{i}\in\{0,\ldots,K\}$ and may be assigned to a worker at any step $x_{i}\geq k_{i}$. Once assigned, the request is processed on a single worker until completion; we do not allow migration between workers (assignments are sticky) and we do not preempt processing.
The processing requirement of request $i$ is represented by a workload profile
$W_{i}=\bigl(w_{i}^{(1)},w_{i}^{(2)},\ldots,w_{i}^{(o_{i})}\bigr),$
where $o_{i}\in\mathbb{N}$ is the total number of processing steps for request $i$, and $w_{i}^{(j)}\geq 0$ denotes the workload contributed by $i$ in its $j$-th processing step. Thus, once $i$ starts processing at time $x_{i}$, it occupies exactly $o_{i}$ consecutive steps $x_{i},x_{i}+1,\ldots,x_{i}+o_{i}-1$, and its workload at time $k$ (while active) is $w_{i}^{(k-x_{i}+1)}$.
Two simple examples illustrate this notation. If a request $i$ has constant workload $5$ in each of $3$ steps, then $W_{i}=(5,5,5)$. In LLM inference, if a request $i$ has a KV-cache load that grows linearly with the number of generated tokens and starts with prefill size $3$, then a run of $4$ decoding steps corresponds to $W_{i}=(3,4,5,6)$. In general, the full profile $W_{i}$ is unknown to the scheduler at arrival time; it is treated as fixed but unobserved.
Policies and per-step workloads.
A non-anticipative policy $\pi$ specifies, for each request $i$ and each step $k\geq k_{i}$, whether to assign $i$ at time $k$, and if so to which worker $g\in[G]$, subject to the capacity constraint that at most $B$ requests are actively processed on each worker at every step. Let $g(i)$ denote the worker to which request $i$ is assigned under policy $\pi$, and $x_{i}$ its assignment time.
For each worker $g$ and step $k$, define the active set
$\mathcal{A}_{g}(k)=\bigl\{i\in\mathcal{I}:g(i)=g,x_{i}\leq k<x_{i}+o_{i}\,\bigr\},$
i.e., the set of requests that are currently being processed on worker $g$ at time $k$. The instantaneous workload of worker $g$ at time $k$ is then
$L_{g}(k)=\sum_{i\in\mathcal{A}_{g}(k)}w_{i}^{(k-x_{i}+1)}.$ (1)
We denote by
$L_{g^{*}}(k)\;=\;\max_{g\in[G]}L_{g}(k)$
the maximum workload across all workers at time $k$.
Imbalance and objective.
In many parallel systems, including data-parallel LLM decoding, the processing time of a step on a worker is linear in its workload, and all workers must wait for the slowest worker at each step (a barrier). Thus, the idle work at step kk—the total amount of work that other workers could have done if they were as heavily loaded as the busiest worker—is proportional to
$\text{Imbalance}(k;\pi)\;=\;\sum_{g=1}^{G}\bigl(L_{g^{*}}(k)-L_{g}(k)\bigr)\;=\;G\cdot L_{g^{*}}(k)-\sum_{g=1}^{G}L_{g}(k),$ (2)
where we make the dependence on the policy $\pi$ explicit. Over a finite horizon $\{1,\ldots,K\}$, we define the average imbalance of a policy $\pi$ as
$\text{AvgImbalance}(\pi)\;=\;\frac{1}{K}\sum_{k=1}^{K}\text{Imbalance}(k;\pi).$
Our goal is to design an assignment policy $\pi$ that makes $\text{AvgImbalance}(\pi)$ as small as possible. Intuitively, this corresponds to maximizing the amount of “useful” work performed each step and minimizing the total idle time induced by waiting for the slowest worker at the barrier.
4 Method: A Universal Load Balance Principle
In the fully agnostic model of Section 3, where the scheduler has no information about $W_{i}$ beyond arrivals and completions, no policy can differentiate between two requests with the same arrival time if their future workloads are indistinguishable. In practice, however, many systems expose partial information about the near future. Our goal in this section is to formalize a minimal and broadly feasible form of such information, and then to develop a load-balancing principle that uses it in a disciplined way.
Short lookahead workload information.
We introduce a short lookahead window of length $H\in\mathbb{N}$ and assume that, at each step $k$, the scheduler may observe for every active request $i$ an estimate of its contributions to future workloads over the next $H$ steps after the current one. Formally, let
$\widehat{W}_{i}^{H}(k)\;=\;\bigl(\widehat{w}_{i}^{(1)}(k),\widehat{w}_{i}^{(2)}(k),\ldots,\widehat{w}_{i}^{(H)}(k)\bigr)$
denote any prediction of the workloads that request $i$ will contribute in its next $H$ processing steps, i.e., at times $k+1,\ldots,k+H$, conditional on the information available at time $k$. When $i$ finishes before $H$ further steps, we interpret the remaining entries as zeros. The special case $H=0$ corresponds to having no lookahead beyond the current step: the scheduler only knows the current workloads $\{L_{g}(k)\}$ and has no predictive signal about $k+1,k+2,\ldots$
This interface captures situations where the total workload of a job is hard to predict at arrival, but its short-term behavior is easier to infer once the job has progressed. In LLM serving, for example, the total number of decode steps for a request is difficult to predict accurately from the prompt alone, yet once a response is near completion (e.g., exhibits “in conclusion”, closing punctuation, or termination tokens), it becomes much easier to estimate whether the request will complete within the next $H$ steps. In the LLM serving setting, the per-step workload is driven by the KV cache, which grows by one token per decode step until termination, so $\widehat{W}_{i}^{H}(k)$ effectively reduces to a prediction of whether and when a request will finish within the lookahead window. As another example, consider a multi-stage document-processing pipeline where early stages have highly variable runtimes, but once a document reaches the final stage, sensors or partial results reveal that only a short, nearly deterministic tail of work remains; here, $\widehat{W}_{i}^{H}(k)$ summarizes this remaining tail, even if the total processing time from arrival was unpredictable.
The universal Balance Future with Integer Optimization principle (BF-IO).
Given access to short lookahead workload information, we now state our universal load-balancing principle. At a high level, the scheduler should choose assignments that make the future workloads across workers as balanced as possible, rather than greedily equalizing only the current step. To make this precise, fix a window length $H\geq 0$ and consider step $k$. Let $R_{\text{wait}}(k)$ be the set of waiting requests, and let $\texttt{cap}[g](k)$ be the number of free slots on worker $g$ at that time. An allocation at step $k$ consists of disjoint subsets $\{S_{g}(k)\}_{g\in[G]}$ of $R_{\text{wait}}(k)$, with
$S_{g}(k)\subseteq R_{\text{wait}}(k),\quad S_{g}(k)\cap S_{g^{\prime}}(k)=\emptyset\ \text{for }g\neq g^{\prime},\quad|S_{g}(k)|\leq\texttt{cap}[g](k)\ \text{for all }g.$
We write
$S(k)\;=\;\bigcup_{g\in[G]}S_{g}(k)$
for the set of all requests admitted at step $k$. For any such allocation $S(k)$, the induced per-worker sets $\{S_{g}(k)\}$ determine an updated active set $\{\mathcal{A}_{g}(k)\}$ and hence a predicted load trajectory over the next $H$ steps, using the profile estimates $\{\widehat{W}_{i}^{H}(k)\}$ for all active and newly admitted jobs.
Let $L_{g}(k)$ denote the current workload on worker $g$ at time $k$ and $L_{g}(k+h)$ the predicted workload at time $k+h$ under allocation $S(k)$ for $h\geq 1$, and let $\text{Imbalance}(k+h)$ be the corresponding imbalance as in Equation (2). The Balance–Future principle is then to choose, at each step $k$, an allocation $S(k)$ that minimizes the accumulated predicted imbalance over the window:
$J\bigl(S(k)\bigr)\;=\;\sum_{h=0}^{H}\text{Imbalance}(k+h),$
where $h=0$ uses the current workloads and $h\geq 1$ uses the predicted workloads based on $\{\widehat{W}_{i}^{H}(k)\}$.
This minimization can be achieved via an integer optimization. Introduce binary decision variables $x_{ig}\in\{0,1\}$ indicating whether request $i\in R_{\text{wait}}(k)$ is assigned to worker $g$ at step $k$, and define
$S_{g}(k;x)\;:=\;\bigl\{\,i\in R_{\!\text{wait}}(k):x_{ig}=1\,\bigr\},\qquad S(k;x)\;:=\;\bigcup_{g\in[G]}S_{g}(k;x),$
so that $S(k;x)$ is the admitted set induced by the assignment matrix $x=\{x_{ig}\}$. We write $J_{k}(x):=J\bigl(S(k;x)\bigr)$ for the accumulated predicted imbalance cost at step $k$. Let
$U(k)\;:=\;\min\Bigl\{\lvert R_{\!\text{wait}}(k)\rvert,\ \sum_{g=1}^{G}\texttt{cap}[g](k)\Bigr\}$
denote the number of slots that can be filled at step $k$ (either all waiting requests, or all free capacity, whichever is smaller). Then the optimal minimization choice at step $k$ is the solution of the following integer optimization:
$\min_{\{x_{ig}\}}$ $J_{k}(x)\;=\;J\bigl(S(k;x)\bigr)$ (IO)
s.t. $\sum_{g=1}^{G}x_{ig}\;\leq 1,$ $\forall\,i\in R_{\!\text{wait}}(k)\quad\text{(each request assigned to at most one worker)},$
$\sum_{i\in R_{\!\text{wait}}(k)}x_{ig}\;\leq\;\texttt{cap}[g](k),$ $\forall\,g\in[G]\quad\text{(per-worker capacity constraints)},$
$\sum_{g=1}^{G}\sum_{i\in R_{\!\text{wait}}(k)}x_{ig}\;=\;U(k),$ $\text{(full utilization of available slots or waiting requests)},$
$x_{ig}\in\{0,1\},$ $\forall\,i\in R_{\!\text{wait}}(k),\ \forall\,g\in[G].$
An optimal solution $x^{\star}$ to (IO) yields the selected allocation $\{S_{g}^{\star}(k)\}_{g\in[G]}$ with
$S_{g}^{\star}(k)\;:=\;S_{g}\bigl(k;x^{\star}\bigr)\;=\;\bigl\{\,i\in R_{\!\text{wait}}(k):x_{ig}^{\star}=1\,\bigr\}.$
In words, at each step BF-IO solves an integer optimization that simultaneously assigns requests to workers and minimizes a finite-horizon prediction of future imbalance.
We capture this procedure in the following step-wise pseudocode. Conceptually, the “inner loop” that enumerates feasible allocations and chooses the one minimizing $J$ corresponds exactly to solving the integer optimization (IO).
Input: Waiting set $R_{\!\text{wait}}(k)$, active sets $\{\mathcal{A}_{g}(k)\}$, free slots $\{\texttt{cap}[g](k)\}$, lookahead window length $H$
Output: Assignments $\{S_{g}(k)\}_{g\in[G]}$ at step $k$
1exCompute the number of admissible assignments:
$U(k)\leftarrow\min\Bigl\{\lvert R_{\!\text{wait}}(k)\rvert,\ \sum_{g}\texttt{cap}[g](k)\Bigr\}\,;$
1exEnumerate all feasible allocations $\{S_{g}(k)\}$ with $S_{g}(k)\subseteq R_{\!\text{wait}}(k)$, $|S_{g}(k)|\leq\texttt{cap}[g](k)$ for all $g$, $S_{g}(k)\cap S_{g^{\prime}}(k)=\emptyset$ for $g\neq g^{\prime}$, and $\bigl|\bigcup_{g}S_{g}(k)\bigr|=U(k)$;
for each feasible allocation $\{S_{g}(k)\}$ do
   Form $S(k)\leftarrow\bigcup_{g\in[G]}S_{g}(k)$;
   Simulate the predicted load trajectory $\{L_{g}(k+h)\}_{h=0}^{H}$ for this allocation using $\{\widehat{W}_{i}^{H}(k)\}$;
   Compute the imbalance cost
$J\bigl(S(k)\bigr)\;=\;\sum_{h=0}^{H}\text{Imbalance}(k+h)\,;$
Select an allocation $\{S_{g}^{\star}(k)\}$ that minimizes $J\bigl(S(k)\bigr)$ (equivalently, solves the integer optimization (IO));
Assign requests according to $\{S_{g}^{\star}(k)\}$;
Algorithm 1 Balance–Future Algorithm (BF-IO)
Although BF-IO (Algorithm 1) is not globally optimal over the entire time horizon $\{0,\ldots,K\}$, it explicitly balances across the near future rather than acting myopically. When $H=0$, the window collapses to the current step and $J\bigl(S(k)\bigr)$ reduces to $\text{Imbalance}(k)$; the algorithm selects the assignment that best equalizes the next workloads $\{L_{g}(k)\}$, using no lookahead beyond the current step. Larger $H$ allows the scheduler to anticipate near-term completions and avoid assignments that look good now but create pronounced imbalance a few steps later.
Intuition and choice of window length $H$.
The Balance–Future principle can be understood as a compromise between two extremes. When $H$ is very small (e.g., $H=0$), the scheduler only equalizes the current-step workloads; this is cheap and prediction-free but can be short-sighted, because it ignores imminent completions that would quickly relieve heavy workers or jobs that are about to enter the system. At the other extreme, taking $H$ very large is also undesirable: the algorithm optimizes the effect of current assignments over a long window without knowing future arrivals or the future assignment decisions that will be made within that window, so its predictions become increasingly speculative and dominated by unmodeled randomness. In practice, a moderate window HH—large enough to see a few steps of meaningful tail behavior, but small enough that predictions remain informative and computational costs stay low—strikes the right balance (See Figure 4). Our theory in Section 5 shows that even the purely myopic case $H=0$ already yields strong worst-case gains over FCFS-type policies, and our experiments in Section 6 confirm that modest $H>0$ can further improve performance when realistic predictive signals are available.
Figure 4: Intuition for selecting the window length $H$, evaluated across values from 0 to $100$. This analysis uses an implementation of the Algorithm BF-IO in the LLM serving context. Further implementation and experimental details are provided in Sections 6.
5 Theoretical Guarantee of the Load Balance Principle in LLM Inference
In Section 4, we introduced the Balance–Future (BF-IO) as a universal principle for load balancing under sticky assignments and barrier synchronization. In this section, we provide theoretical guarantee for its effectiveness by analyzing its performance in the LLM decode serving setting and then discussing extensions to more general workload patterns. Our primary focus is the data-parallel (DP) decoding bottleneck in LLM serving: we show that, under a realistic overloaded regime and for any adversarial arrival instance, BF-IO achieves significantly smaller imbalance than the baseline First-Come-First-Serve (FCFS) policy used in many real systems. A detailed description of the baseline FCFS policy can be found in Appendix B. We then outline how the analysis framework extends beyond LLM serving to a broader family of load balancing problems.
LLM serving-specific workload model.
We specialize the general model of Section 3 to the DP decode stage of LLM serving. Recall that $G$ denotes the number of computational workers (e.g., GPUs) and $B$ the maximum batch size (number of requests that can be processed simultaneously on a single worker). We assume that $\sqrt{G}=\tilde{o}(B)$, which is an extremely mild assumption in practice: typical batch sizes are $B\in[64,128]$, whereas the number of workers $G$ is at most in the hundreds in real deployments.
As in Section 3, each request $i$ is represented by a workload profile
$W_{i}\;=\;\bigl(w_{i}^{(1)},w_{i}^{(2)},\ldots,w_{i}^{(o_{i})}\bigr),$
where $o_{i}\in\mathbb{N}$ is the total number of processing steps for request $i$, and $w_{i}^{(j)}\geq 0$ denotes the workload contributed by $i$ in its $j$-th processing step. In the LLM setting, the initial workload $w_{i}^{(1)}$ corresponds to the prefill size, which we denote by $s_{i}$. We assume that the prefill lengths $\{s_{i}\}$ are i.i.d. draws from an unknown bounded distribution $\mathcal{D}_{\mathrm{prefill}}$ supported on $\{1,\ldots,s_{\max}\}$, and let $\sigma_{s}$ be the standard deviation of $\mathcal{D}_{\mathrm{prefill}}$. We impose a non-degeneracy condition: there exists $\kappa_{0}\in(0,\tfrac{1}{2}]$ such that
$\kappa_{0}\;\leq\;\frac{\sigma_{s}}{s_{\max}}\;\leq\;\frac{1}{2}.$
The upper bound follows from Popoviciu’s inequality, which implies $\sigma_{s}/s_{\max}\leq 1/2$ for any distribution supported on an interval of length $s_{\max}$. The lower bound $\kappa_{0}>0$ simply rules out pathological cases where almost all prompts have essentially identical prefill length.
After prefill, LLM decoding proceeds token by token. Once request $i$ starts decoding, it occupies exactly $o_{i}-1$ consecutive decode steps and its $k$-th decode step ($k\in[o_{i}-1]$) requires workload $s_{i}+k$, reflecting the growth of its KV cache. We assume that the decode lengths $\{o_{i}\}$ are i.i.d. draws from an unknown distribution $\mathcal{D}_{\mathrm{decode}}$ on $\{1,2,\ldots\}$. Thus the full workload profile of request $i$ takes the form
$W_{i}\;=\;\bigl(s_{i},s_{i}+1,\ldots,s_{i}+o_{i}-1\bigr).$
Overloaded arrival instances.
We now formalize the overloaded regime, which is the relevant regime for LLM serving: if the system were lightly loaded, all requests could be processed immediately without sophisticated scheduling. Intuitively, an arrival instance is overloaded if, at every step, there are “many more” pending requests than free slots. Moreover, we assume that no single prefill-length class dominates the pool. The latter condition reflects the empirical fact that real workloads contain a variety of prompt lengths rather than being concentrated on a single fixed prefill size.
Formally, at each step $k$, let $\mathcal{I}_{k}$ denote the set of requests that are pending in the waiting pool, and let
$\mathcal{R}_{k}\;=\;\{(s_{i},o_{i}):i\in\mathcal{I}_{k}\}$
denote the multiset of their prefill and decode lengths. For $\ell\in\{1,\dots,s_{\max}\}$, define the length class
$\mathcal{R}_{k}(\ell)\;:=\;\{\,i\in\mathcal{I}_{k}:s_{i}=\ell\,\}.$
Let $C_{k}=\sum_{g=1}^{G}c_{g}(k)$ denote the total number of free slots at step $k$, where $c_{g}(k)$ is the number of available batch slots on worker $g$.
Definition 1 (Overloaded arrival instance family $\mathcal{I}$).
An arrival instance $I$ is called overloaded if, for every step $k$,
$\min_{\ell\in\{1,\dots,s_{\max}\}}\Bigl(\,\bigl|\mathcal{R}_{k}\setminus\mathcal{R}_{k}(\ell)\bigr|\,\Bigr)\ \geq\ C_{k}.$
Let $\mathcal{I}$ denote the family of all overloaded arrival instances.
In words, even after removing all requests belonging to whichever single prefill-length class is most numerous at step $k$, the remaining pool still contains at least $C_{k}$ requests and can therefore fill all free slots freed in step $k$. This condition is typically satisfied in heavily loaded real systems: prompt lengths exhibit substantial diversity, and the number of pending requests is far larger than the instantaneous capacity.
Specialization of BF-IO to $H=0$.
As discussed in Section 4, BF-IO can, in principle, use a short lookahead window of length $H\geq 0$. From an analytical perspective, however, rigorous control of the dynamics for $H\geq 1$ is extremely involved, because current assignments interact with future arrivals and future decisions within the window. In this section we therefore focus on the case $H=0$, where BF-IO reduces to selecting, at each step, the assignment that minimizes the current-step imbalance. As we will show, even this purely myopic variant achieves substantial improvement over FCFS in the LLM decode model. This is fantastic since we demonstrate that modest $H>0$ yields further gains when realistic predictive signals are available in Section 4.
Imbalance improvement ratio (IIR).
We compare BF-IO and FCFS via an imbalance improvement ratio, which measures, in the worst arrival instance, the long-run factor by which BF-IO reduces average imbalance relative to FCFS. For a given arrival instance $I\in\mathcal{I}$ and policy $\pi\in\{\textsf{BF-IO},\textsf{FCFS}\}$, let $\text{AvgImbalance}(\pi;I)$ denote the time-average imbalance defined in Section 3. We define
$\mathbf{IIR}\;=\;\inf_{I\in\mathcal{I}}\left(\lim_{K\to\infty}\frac{\mathbb{E}\bigl[\text{AvgImbalance}(\textsf{FCFS};I)\bigr]}{\mathbb{E}\bigl[\text{AvgImbalance}(\textsf{BF-IO};I)\bigr]}\right),$
where the expectation is taken over the randomness in the prefill and decode lengths $(s_{i},o_{i})$ drawn from $\mathcal{D}_{\mathrm{prefill}}$ and $\mathcal{D}_{\mathrm{decode}}$. The arrival sequence $I$ itself is chosen adversarially from $\mathcal{I}$; in particular, the adversary can choose any number of arrivals at each step, subject only to the overloaded condition. Larger values of $\mathbf{IIR}$ correspond to larger worst-case improvements of BF-IO over FCFS.
Directly lower bounding $\mathbf{IIR}$ in full generality is challenging. To build intuition, we first study a warm-up setting in which all decode lengths are homogeneous: $o_{i}=o$ for every request $i$. In this case, the dynamics simplify substantially: once $GB$ jobs are admitted, they run in perfect lockstep for $o$ decode steps and then all complete simultaneously. Under the overloaded condition (so that every worker runs at full batch capacity), the per-step imbalance within a round of $o$ steps is constant; optimizing long-run average imbalance reduces to optimizing the imbalance right after each batch admission.
Theorem 1 (Warm-up: Homogeneous decode lengths).
Consider the homogeneous-decode model $o_{i}=o$ for all requests $i$. There exists a universal constant $c>0$ such that
$\mathbf{IIR}\;\geq\;c\,\kappa_{0}\,\sqrt{B\log G}\,\cdot\frac{G}{G-1}\;=\;\Omega\!\big(\sqrt{B\log G}\big).$ (3)
Proof intuition: In the homogeneous-output setting, each admission round is structurally simple: once $GB$ jobs are admitted, they run in perfect lockstep for $o$ token steps and then all finish simultaneously. As a result, all steps in the round share the same inter-device imbalance, and minimizing the long-run time-average imbalance is equivalent to minimizing the imbalance immediately after admission in each round.
On the BF-IO side, this reduction means that BF-IO is effectively solving a static bin-packing problem in every round: partition the $GB$ prompt lengths into $G$ groups of size $B$ to minimize the maximum per-device load. A simple exchange argument shows that any such optimal assignment must be $s_{\max}$-balanced: the difference between the heaviest and lightest devices can never exceed $s_{\max}$. This immediately implies that the expected BF-IO imbalance per round is at most $(G-1)s_{\max}$, independent of the detailed shape of the prefill-length distribution.
On the FCFS side, the story is almost the opposite. Under size-agnostic FCFS admission and an overloaded pool, each device simply sees $B$ i.i.d. prompt lengths. The per-device loads $S_{g}=\sum_{j=1}^{B}s_{g,j}$ are thus i.i.d. with variance $\sigma_{s}^{2}B$. For large $B$, the central limit theorem (and a Berry–Esseen bound) shows that $S_{g}$ behaves approximately Gaussian. Taking the maximum over $G$ such devices amplifies these fluctuations: with constant probability, at least one device deviates above the mean by order $\sigma_{s}\sqrt{B\log G}$. This produces an expected FCFS imbalance of order $G\sigma_{s}\sqrt{B\log G}$, while BF-IO keeps the imbalance of order $(G-1)s_{\max}$. The non-degeneracy condition $\sigma_{s}/s_{\max}\geq\kappa_{0}>0$ then yields an imbalance-reduction factor of order $\sqrt{B\log G}$, uniformly over all overloaded arrival instances. The complete proof is deferred to Appendix C.1. $\square$
Inhomogeneous setting.
Figure 5: Empirical distributions of decode lengths from production LLM traces [wang2023openchat, wang2024burstgpt, zheng2023lmsys, zhao2024wildchat]. Decode lengths exhibit the geometric (discrete-exponential) pattern.
The homogeneous-decode assumption $o_{i}=o$ in Theorem 1 serves as a warm-up: it is analytically convenient but clearly unrealistic for LLM inference, whose responses exhibit substantial variability in length. Empirical studies in Fig 5 on both open-source ([wang2023openchat, wang2024burstgpt, zheng2023lmsys, zhao2024wildchat]) and production traces show that decode lengths typically follow a heavy right-skewed pattern well approximated by a geometric (or discrete exponential) distribution: most responses terminate quickly, while a non-negligible tail runs for many tokens. Motivated by this evidence, we now consider an inhomogeneous setting in which decode lengths are random and independent across requests. Specifically, we assume
$o_{i}\;\sim\;\mathcal{D}_{\mathrm{decode}}\;=\;\mathrm{Geo}(p)\quad\text{i.i.d.\ for each request }i,$
for some $p\in(0,1)$, so that the mean and variance of $o_{i}$ are $1/p$ and $(1-p)/p^{2}$, respectively. Under this more realistic decode model, the next theorem shows that BF-IO still achieves a worst-case imbalance improvement factor of order $\Omega\!\big(\sqrt{B\log G}\big)$, with an explicit dependence on both the prefill and decode variability.
Theorem 2 (Inhomogeneous decode lengths).
Consider the inhomogeneous-decode model in which $o_{i}\sim\mathrm{Geo}(p)$ i.i.d. for each request $i$ with some $p\in(0,1)$. Then, there exists a universal constant $c>0$ such that
$\mathbf{IIR}\;\geq\;c\,\frac{p}{s_{\max}}\,\sqrt{\sigma_{s}^{2}+\frac{1-p}{p^{2}}}\;\frac{G}{G-1}\,\sqrt{B\log G}\;=\;\Omega\!\big(\sqrt{B\log G}\big).$ (4)
Proof intuition
Under geometric output lengths with unit per-step growth, the stochastic dynamics of loads under FCFS and BF-IO differ sharply. On the FCFS side, each slot behaves as a simple renewal process: its size increases deterministically by one each step until a geometric completion occurs, at which point it resets to a fresh prompt length $S$. In stationarity, a snapshot of a slot therefore looks like
$\text{(prompt) }S\;+\;\text{(age) }A,$
where $S$ has variance $\sigma_{s}^{2}$ and $A$ is geometric with mean (1p)/p(1-p)/p and variance $(1-p)/p^{2}$, independent of $S$. This yields a per-slot “snapshot” variance
$\sigma_{\mathrm{snap}}^{2}=\sigma_{s}^{2}+\frac{1-p}{p^{2}}.$
Aggregating $B$ i.i.d. slots on each device gives per-device loads with variance $\sigma_{\mathrm{snap}}^{2}B$, and comparing $G$ such independent devices, standard Gaussian extreme-value theory shows that at least one device sits of order $\sigma_{\mathrm{snap}}\sqrt{B\log G}$ above the mean with constant probability. This creates a typical FCFS imbalance of order $G\,\sigma_{\mathrm{snap}}\sqrt{B\log G}$.
On the BF-IO side, the analysis from Part 1 shows that, despite the asynchronous completions, the myopic balancing rule enforces a contraction of the inter-device gap: in each step, pre-admission loads shrink by a factor roughly (1p)(1-p), and the subsequent assignment step of BF-IO prevents the gap from growing by more than $s_{\max}$, up to small error events. Over time, this yields a steady-state average gap of order $s_{\max}/p$, independent of $(B,G)$ up to lower-order terms. Since instantaneous imbalance is at most (G1)(G-1) times the gap, the long-run average imbalance under BF-IO is at most $(G-1)s_{\max}/p$, while FCFS incurs imbalance of order $G\sigma_{\mathrm{snap}}\sqrt{B\log G}$. The resulting imbalance–reduction ratio is therefore of order
$\frac{G\,\sigma_{\mathrm{snap}}\sqrt{B\log G}}{(G-1)\,s_{\max}/p}\;\asymp\;\frac{p\,\sigma_{\mathrm{snap}}}{s_{\max}}\,\sqrt{B\log G},$
matching the homogeneous-output case in $\sqrt{B\log G}$ scaling, but with the snapshot variance $\sigma_{\mathrm{snap}}^{2}$ capturing the additional variability introduced by geometric completion times. The complete proof can be found in Appendix C.2.
5.1 Theoretical Guarantee of BF-IO in General Cases
The LLM-specific analysis above exploits a particular structure of the workload profiles,
$W_{i}\;=\;\bigl(s_{i},s_{i}+1,\ldots,s_{i}+o_{i}-1\bigr),$
where each alive request gains one unit of KV load per decode step. This captures current DP decoding architectures, but many other load-balancing problems do not obey this linear drift. A classical example is the constant-workload setting, where each job has a fixed processing requirement per step and
$W_{i}\;=\;\bigl(s_{i},s_{i},\ldots,s_{i}\bigr)\quad\text{($o_{i}$ copies of $s_{i}$),}$
which arises naturally in traditional CPU or manufacturing workloads. Even within LLM serving, future systems may reduce or compress KV usage dynamically so that per-step cost does not grow exactly by +1+1; instead, every alive request may experience a common but time-varying increment that reflects changes in the underlying architecture or memory-management policy.
To encompass these scenarios, we consider a more general class of non-decreasing workload drift models. We retain the prefill and decode assumptions from the inhomogeneous LLM model—namely, $s_{i}$ drawn from a bounded prefill distribution with variance $\sigma_{s}^{2}$ and decode lengths $o_{i}\sim\mathrm{Geo}(p)$ i.i.d. for some $p\in(0,1)$—but relax the per-step growth pattern as follows.
Definition 2 (Non-decreasing Workload Drift).
There exists a deterministic sequence $(\delta_{k})_{k\geq 1}$ with
$0\;\leq\;\delta_{k}\;\leq\;\delta_{\max}\qquad\text{for all }k\geq 1,$
where $\delta_{\max}$ is a non-negative constant independent of $(B,G)$, such that for every request $i$,
$w_{i}^{(1)}=s_{i},\qquad w_{i}^{(j)}=s_{i}+\sum_{t=1}^{j-1}\delta_{t},\quad j=2,\ldots,o_{i}.$
In words, all alive requests share the same nonnegative increment $\delta_{k}$ at global step $k$; workloads are non-decreasing over time and their per-step growth is uniformly bounded. The LLM decode model with linear KV growth corresponds to the special case $\delta_{k}\equiv 1$ for all $k$, while the constant-workload model corresponds to $\delta_{k}\equiv 0$. More sophisticated architectures, in which per-step KV work is throttled or sparsified over time, are captured by intermediate patterns with $0<\delta_{k}<1$ or time-varying $\delta_{k}$. The following theorem shows that BF-IO continues to enjoy a worst-case imbalance improvement factor of order $\sqrt{B\log G}$ relative to FCFS in this more general setting.
Theorem 3 (General non-decreasing workload drift).
Consider the inhomogeneous-decode model in which $o_{i}\sim\mathrm{Geo}(p)$ i.i.d. for each request $i$ with some $p\in(0,1)$, and assume the non-decreasing drift structure in Definition 2. Then there exists a universal constant $c>0$, depending only on $(\sigma_{s}^{2},p,\delta_{\max},s_{\max})$, such that
$\mathbf{IIR}\;=\;\inf_{I\in\mathcal{I}}\left(\lim_{K\to\infty}\frac{\mathbb{E}\bigl[\text{AvgImbalance}(\textsf{FCFS};I)\bigr]}{\mathbb{E}\bigl[\text{AvgImbalance}(\textsf{BF-IO};I)\bigr]}\right)\;\geq\;c\,\frac{p\,\sigma_{s}}{s_{\max}}\,\frac{G}{G-1}\,\sqrt{B\log G}\;=\;\Omega\!\big(\sqrt{B\log G}\big).$ (5)
Proof intuition: Compared with the “+1+1 per step” case, the core structure of the BF-IO analysis remains unchanged: the gap contraction still comes entirely from completions, not from the deterministic per-step growth. When every survivor gains a common increment $\delta_{k}$ at step $k$, each device’s load is shifted by exactly $B\delta_{k}$, so the inter-device gap D(k1)D(k{-}1) is unaffected by this drift; only the random removals from completions change the spread. As a result, the separation lemma (which enforces $D(k)\lesssim A^{\mathrm{pre}}(k)-Bp$ on good completion-count events) still holds since it depends only on prompt sizes and binomial completions, not on the specific value of $\delta_{k}$. The main technical adjustments are (i) in the concentration lemma, where the per-slot summands are now bounded by $\Lambda(k)+\delta_{\max}$ instead of $\Lambda(k)$, so the Bernstein bound is applied with a slightly larger envelope, and (ii) in the envelope/age accounting, where we use that an alive job of age $r$ has size at most $s_{\max}+\delta_{\max}r$, so the excess $\Lambda(k)-s_{\max}$ is controlled by $\delta_{\max}$ times the maximum age. The geometric tail of the output length then yields the same exponential decay in age (up to constants depending on $\delta_{\max}$), and hence the same $O(s_{\max}/p)$ bound on the long-run average gap as in the +1+1 case. In other words, as long as the per-step increments are common across jobs and uniformly bounded, the contraction mechanism and the averaging argument that underpin the BF-IO upper bound survive unchanged; only the constants in the concentration and tail estimates are rescaled by $\delta_{\max}$.
In the constant +1+1 model, the slot-level process under FCFS is a homogeneous Markov chain that admits a clean stationary representation: at any snapshot, each slot contains “prompt + geometric age”, and the age has variance $(1-p)/p^{2}$. With a time-varying drift sequence $(\delta_{k})$, the chain for $Y_{g,b}(k)$ becomes non-homogeneous, but the key stochastic structure survives: completions still occur as i.i.d. Bernoulli(p)(p) trials, independent of job sizes and of the drift. This means that at any step $k$, the prompt length of the current job in a slot is still an independent copy of $S$, and the number of drift steps it has accumulated—its age—still has geometric tails with parameter $p$, regardless of the past. The per-slot snapshot can therefore be written as
$\text{(prompt)}\;+\;\text{(sum of the last $A_{k}$ drifts)},$
where $A_{k}$ has uniformly bounded moments and $\delta_{k}$ is uniformly bounded. This immediately implies that the per-slot variance is uniformly bounded below by $\sigma_{s}^{2}$ (from the prompt) and the third moment is uniformly bounded above, independent of time. Summing $B$ such independent snapshots on each device yields loads with variance of order $B$, and a Berry–Esseen bound (in its non-identical version) shows that each device’s load fluctuates on the $\sqrt{B}$ scale and is close to Gaussian. Taking the maximum over $G$ devices then produces an excess of order $\sqrt{B\log G}$ above the average device load, with constant probability, exactly as in the homogeneous case. Thus, even though the per-slot chain is no longer time-homogeneous and lacks a simple stationary law, the bounded and common per-step drift leaves the order of the FCFS imbalance unchanged: its expected size remains $\Theta\!\big(G\sqrt{B\log G}\big)$. This concludes that the imbalance-reduction factor is at least on the order of $\sqrt{B\log G}$. The complete proof can be found in Appendix C.3. $\square$
5.2 Theoretical Guarantee of BF-IO on Energy Consumption
We measure total energy consumption (in Joules) as the time integral of instantaneous GPU power:
$E\;=\;\int_{0}^{T}P\bigl(\mathrm{mfu}(t)\bigr)\,dt,$ (6)
where $\mathrm{mfu}(t)$ denotes the Model FLOPs Utilization (mfu), i.e., the ratio of observed computational throughput to the GPU’s theoretical peak FLOPs capacity. Empirical measurements in [ozcan2025quantifying] validate that instantaneous GPU power is a sublinear function of utilization:
$P(\mathrm{mfu})\;=\;P_{\mathrm{idle}}+(P_{\mathrm{max}}-P_{\mathrm{idle}})\left(\frac{\mathrm{mfu}}{\mathrm{mfu}_{\mathrm{sat}}}\right)^{\gamma},\qquad\gamma\in(0,1),$ (7)
where $P_{\mathrm{idle}}$ is idle power, $P_{\mathrm{max}}$ is peak power, $\mathrm{mfu}_{\mathrm{sat}}$ is a saturation threshold, and $\gamma<1$ captures the inefficiency that GPUs draw substantial power even at low utilization.
Synchronized phase and utilization model.
To isolate the impact of barrier-induced waiting, we focus on the synchronized phase of each decoding step (e.g., the attention computation) during which all workers wait for the slowest one. Other phases (e.g., all-to-all EP/TP communication) can be incorporated additively but are omitted here for clarity.
Recall from Section 3 that at discrete step $k$ the instantaneous workload on worker $g$ under policy $\pi$ is $L_{g}(k)$, and $L_{g^{*}}(k)=\max_{g\in[G]}L_{g}(k)$. The wall-clock duration of the synchronized attention phase in step $k$ is linear in the maximal workload, namely, $\tau_{k}(\pi)\;=\;\kappa_{\mathrm{ATT}}\cdot L_{g^{*}}(k)$, for some constant $\kappa_{\mathrm{ATT}}>0$ (seconds per unit workload). During this synchronized phase, worker $g$ performs useful attention computation for $\kappa_{\mathrm{ATT}}L_{g}(k)$ seconds and waits at the barrier for $\kappa_{\mathrm{ATT}}(L_{g^{*}}(k)-L_{g}(k))$ seconds. Hence its utilization fraction within the synchronized phase is
$u_{g}(k)\;:=\;\frac{L_{g}(k)}{L_{g^{*}}(k)}\in[0,1].$ (8)
As the utilization fraction equals to the instant throughput fraction, $\frac{\mathrm{mfu}_{g}(t)}{\mathrm{mfu}_{\mathrm{sat}}}$, we have
$\frac{\mathrm{mfu}_{g}(t)}{\mathrm{mfu}_{\mathrm{sat}}}\;=\;u_{g}(k),\qquad t\ \text{during the synchronized phase of step }k.$ (9)
Under (7)–(9), the per-worker power in step $k$ equals $P_{\mathrm{idle}}+(P_{\mathrm{max}}-P_{\mathrm{idle}})\,u_{g}(k)^{\gamma}.$
Total synchronized-phase energy.
Fix an arrival instance $\mathcal{I}$ and a policy $\pi$. Let $K_{\pi}(\mathcal{I})$ denote the number of decoding steps until all requests in $\mathcal{I}$ complete under $\pi$. The total energy consumed in the synchronized attention phase across all $G$ workers is
$E(\pi;\mathcal{I})$ $:=\sum_{g=1}^{G}\int_{0}^{T_{\pi}}P\!\left(\mathrm{mfu}_{g}(t)\right)\,dt$
$=\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}\sum_{g=1}^{G}\tau_{k}(\pi)\Bigl(P_{\mathrm{idle}}+(P_{\mathrm{max}}-P_{\mathrm{idle}})\,u_{g}(k)^{\gamma}\Bigr)$
$=\kappa_{\mathrm{ATT}}\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}L_{g^{*}}(k)\sum_{g=1}^{G}\Bigl(P_{\mathrm{idle}}+(P_{\mathrm{max}}-P_{\mathrm{idle}})\,u_{g}(k)^{\gamma}\Bigr),$ (10)
where $T_{\pi}=\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}\tau_{k}(\pi)$ is the corresponding makespan of the synchronized phase.
Define the policy-independent total attention workload of the instance $\mathcal{I}$ as:
$W(\mathcal{I})\;:=\;\sum_{i\in\mathcal{I}}\sum_{j=1}^{o_{i}}w_{i}^{(j)}\;=\;\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}\sum_{g=1}^{G}L_{g}(k),$ (11)
where the second equality holds for every policy $\pi$ because each request contributes its full workload profile exactly once regardless of assignment timing or placement. Also define the cumulative imbalance up to completion:
$\mathrm{ImbTot}(\pi;\mathcal{I})\;:=\;\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}\mathrm{Imbalance}(k;\pi)\;=\;\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}\sum_{g=1}^{G}\bigl(L_{g^{*}}(k)-L_{g}(k)\bigr).$ (12)
Finally, for a baseline policy $\pi_{0}$, define the normalized imbalance level
$\eta_{\mathrm{sum}}(\pi_{0};\mathcal{I})\;:=\;\frac{\mathbb{E}\bigl[\mathrm{ImbTot}(\pi_{0};\mathcal{I})\bigr]}{\mathbb{E}\bigl[W(\mathcal{I})\bigr]}.$ (13)
Energy savings from imbalance reduction.
We now show that any guaranteed improvement in imbalance yields an explicit guaranteed improvement in total synchronized-phase energy (as a percentage), with constants depending only on the power model parameters.
Theorem 4 (Energy saving guarantee from imbalance reduction).
Let $\pi_{0}$ be any baseline policy and $\pi_{1}$ any improved policy. Suppose the imbalance improvement factor satisfies
$\mathbf{IIR}\;=\;\inf_{I\in\mathcal{I}}\left(\lim_{K\to\infty}\frac{\mathbb{E}\bigl[\text{AvgImbalance}(\pi_{0};I)\bigr]}{\mathbb{E}\bigl[\text{AvgImbalance}(\pi_{1};I)\bigr]}\right)\ \geq\ \alpha\qquad\text{for some }\alpha>1.$ (14)
Define
$C_{\gamma}\;:=\;(1-\gamma)P_{\mathrm{max}}+\gamma P_{\mathrm{idle}},\qquad D_{\gamma}\;:=\;(1-\gamma)(P_{\mathrm{max}}-P_{\mathrm{idle}}),$ (15)
then the synchronized-phase energy saving fraction satisfies
$\frac{\mathbb{E}\bigl[E(\pi_{0};\mathcal{I})\bigr]-\mathbb{E}\bigl[E(\pi_{1};\mathcal{I})\bigr]}{\mathbb{E}\bigl[E(\pi_{0};\mathcal{I})\bigr]}\ \geq\ \frac{1}{P_{\mathrm{max}}/\eta_{\mathrm{sum}}(\pi_{0};\mathcal{I})+C_{\gamma}}\left(P_{\mathrm{idle}}\Bigl(1-\frac{1}{\alpha}\Bigr)\;-\;\frac{D_{\gamma}}{\alpha}\right).$ (16)
Proof sketch.
Equation (10) expresses $E(\pi;\mathcal{I})$ as a sum over steps of the barrier duration $\kappa_{\mathrm{ATT}}L_{g^{*}}(k)$ times power. Writing $u_{g}(k)^{\gamma}=u_{g}(k)+(u_{g}(k)^{\gamma}-u_{g}(k))$ and using $L_{g^{*}}(k)u_{g}(k)=L_{g}(k)$ yields an exact decomposition:
$E(\pi;\mathcal{I})=\kappa_{\mathrm{ATT}}P_{\mathrm{max}}W(\mathcal{I})+\kappa_{\mathrm{ATT}}P_{\mathrm{idle}}\mathrm{ImbTot}(\pi;\mathcal{I})+\text{(concavity correction)}.$
Since $\gamma\in(0,1)$, $u^{\gamma}$ is concave on $[0,1]$; the tangent bound at $u=1$ implies that the concavity correction is nonnegative and at most $\kappa_{\mathrm{ATT}}D_{\gamma}\,\mathrm{ImbTot}(\pi;\mathcal{I})$. This shows $E(\pi;\mathcal{I})$ is sandwiched between two affine functions of $\mathrm{ImbTot}(\pi;\mathcal{I})$ with the same policy-independent baseline term $\kappa_{\mathrm{ATT}}P_{\mathrm{max}}W(\mathcal{I})$. Combining these bounds with the imbalance ratio assumption (14) and normalizing by $W(\mathcal{I})$ yields (16). The complete proof can be found in Appendix C.4.
Remark 1 (Instantiation to $\pi_{0}=\textsf{FCFS}$ and $\pi_{1}=\textsf{BF-IO}$ in our LLM model).
In the inhomogeneous LLM decode model of Section 5, if we take $\pi_{0}=\textsf{FCFS}$ and $\pi_{1}=\textsf{BF-IO}$, then (14) holds with $\alpha=c\,\frac{p}{s_{\max}}\,\sqrt{\sigma_{s}^{2}+\frac{1-p}{p^{2}}}\;\frac{G}{G-1}\,\sqrt{B\log G}$, and thus Theorem 4 yields a nontrivial constant lower bound on the synchronized-phase energy saving fraction once $\mathbf{IIR}$ is sufficiently large. Moreover, under the overloaded regime and geometric output lengths, the normalized imbalance level $\eta_{\mathrm{sum}}(\textsf{FCFS};\mathcal{I})$ admits an explicit lower bound:
$\eta_{\mathrm{sum}}(\textsf{FCFS};\mathcal{I})\ \gtrsim\ \frac{\sigma_{\mathrm{snap}}}{\mu_{s}+\frac{1-p}{p}}\sqrt{\frac{\log G}{B}},\qquad\sigma_{\mathrm{snap}}^{2}=\sigma_{s}^{2}+\frac{1-p}{p^{2}}.$ (17)
Plugging (17) and $\alpha=c\,\frac{p}{s_{\max}}\,\sqrt{\sigma_{s}^{2}+\frac{1-p}{p^{2}}}\;\frac{G}{G-1}\,\sqrt{B\log G}$ into (16) yields an explicit energy saving guarantee in terms of $(p,\sigma_{s},s_{\max},B,G)$ and $(P_{\mathrm{idle}},P_{\mathrm{max}},\gamma)$. The complete proof of Inequality (17) can be found in Appendix C.5
Corollary 1 (Asymptotic energy saving at scale).
Under the conditions of Remark 1, as the number of workers $G\to\infty$, the guaranteed energy saving fraction of BF-IO relative to FCFS converges to
$\lim_{G\to\infty}\frac{\mathbb{E}\bigl[E(\textsf{FCFS};\mathcal{I})\bigr]-\mathbb{E}\bigl[E(\textsf{BF-IO};\mathcal{I})\bigr]}{\mathbb{E}\bigl[E(\textsf{FCFS};\mathcal{I})\bigr]}\ \geq\ \frac{P_{\mathrm{idle}}}{(1-\gamma)P_{\mathrm{max}}+\gamma P_{\mathrm{idle}}}.$ (18)
Remark 2 (Concrete instantiation for modern GPUs).
For NVIDIA A100 GPUs, empirical measurements in [ozcan2025quantifying] report $P_{\mathrm{idle}}=100$W, $P_{\mathrm{max}}=400$W, and $\gamma=0.7$. Substituting into (18) in Corollary 1:
$\frac{P_{\mathrm{idle}}}{(1-\gamma)P_{\mathrm{max}}+\gamma P_{\mathrm{idle}}}\;=\;\frac{100}{0.3\times 400+0.7\times 100}\;=\;\frac{100}{190}\;\approx\;52.6\%.$
That is, in large-scale deployments, the principled load balancing introduced in this work can reduce synchronized-phase energy consumption by more than half compared to the widely-deployed FCFS baseline. This asymptotic guarantee underscores the sustainability impact of addressing the decode-stage load imbalance bottleneck at fleet scale.
6 Experiments
We evaluate the proposed routing strategies through discrete-event simulation of GPU-based LLM request processing. This section describes the experimental setup, evaluation metrics, and numerical results.
6.1 Dataset and Request Generation
To simulate realistic LLM inference workloads, we use request traces derived from LongBench [bai2024longbench], a challenging benchmark designed for long-context language models. LongBench encompasses diverse long-text application scenarios—single-document and multi-document question answering, summarization, few-shot learning, synthetic tasks, and code completion—and reflects production-relevant settings in which input prompts are substantially longer than typical benchmarks. Each request is characterized by a pair $(s_{i},o_{i})$, where $s_{i}$ denotes the prefill (input prompt) length and $o_{i}$ the decode (output) length; the resulting distributions are shown in Figure 6. Requests arrive according to a stationary Poisson process at a rate exceeding the system’s processing capacity, ensuring that the system operates in the overloaded regime central to our theoretical analysis. For comparison, we also evaluate on BurstGPT [wang2025burstgpt], a lighter-load trace; these results appear in Appendix D.2.
Figure 6: Prefill and decode length distribution of LongBench.
6.2 Simulation Environment
The router simulator models an online setting with $G=256$ A100 GPU workers, each supporting up to $B=72$ concurrent requests (batch size). Key components include:
  • Undiscovered queue: Requests not yet revealed to the router.
  • Wait queue: Candidate requests available for routing decisions.
  • Active sets $\mathcal{A}_{g}$: Requests currently executing on GPU $g$.
  • Load tracking $L_{g}$: Total token load on GPU $g$, defined in Equation (1).
Time Progression Model.
Each simulation step advances time by:
$\Delta t=C+t_{\ell}\cdot\max_{g\in[G]}L_{g}$ (19)
where $C=9.775\cdot 10^{-3}$ is a fixed overhead constant and $t_{\ell}=1.005\cdot 10^{-7}$ is the per-token generation latency. The values of $C$ and $t_{\ell}$ are derived from a minimum mean square error regression on real-world traces. During each step, all active requests generate one token. A request $i$ completes when $\tau_{i}\geq o_{i}$.
Routing Policies.
We evaluate four routing policies: (i) FCFS, (ii) Join-Shortest-Queue (JSQ), (iii) BF-IO ($H=0$), (iv) BF-IO ($H=20$), (vi) BF-IO ($H=40$), (vii) BF-IO ($H=60$), (viii) BF-IO ($H=80$), (ix) BF-IO ($H=100$).
6.3 Evaluation Metrics
We evaluate routing performance using four metrics. Let $G$ denote the number of GPUs, $\mathcal{A}_{g}^{(k)}$ the active requests on GPU $g$ at step $k$, and $L_{g}(k)=\sum_{i\in\mathcal{A}_{g}^{(k)}}w_{i}^{(k-x_{i}+1)}$ the instantaneous workload on worker $g$, where $w_{i}^{(j)}$ is the workload contributed by request $i$ in its $j$-th processing step.
Average Imbalance.
Following the theoretical framework, we adopt the average imbalance as the primary metric. At each step $k$, the instantaneous imbalance is defined as:
$\text{AvgImbalance}=\frac{1}{K}\sum_{k=1}^{K}\text{Imbalance}(k)$ (20)
In addition to the theoretically motivated imbalance metric, we evaluate two metrics that directly reflect real-world LLM inference efficiency:
Throughput.
Throughput measures the token processing rate (tokens/second):
$\text{Throughput}=\frac{\sum_{k=1}^{K}\left|\mathcal{A}^{(k)}\right|}{\sum_{k=1}^{K}\Delta t^{(k)}}$ (21)
where $\mathcal{A}^{(k)}=\bigcup_{g}\mathcal{A}_{g}^{(k)}$ denotes all active requests at step $k$, and $\Delta t^{(k)}=C+t_{\ell}\cdot\max_{g}L_{g}(k)$ is the wall-clock duration of step $k$. Since step duration is governed by the maximum load across workers, reducing imbalance directly increases throughput.
Time Per Output Token (TPOT).
TPOT captures per-request latency as perceived by users:
$\text{TPOT}=\frac{1}{N}\sum_{i=1}^{N}\frac{t_{i}^{\text{finish}}-t_{i}^{\text{start}}}{o_{i}}$ (22)
where $t_{i}^{\text{start}}$ and $t_{i}^{\text{finish}}$ are the start and completion times of request $i$, and $o_{i}$ is its output length. Requests assigned to heavily loaded GPUs experience longer per-step durations due to the barrier synchronization, resulting in higher TPOT. This metric is widely used in production LLM serving systems to measure user-facing latency.
Energy Consumption.
As is mentioned in Section 5.2, the total energy consumption (in Joules) is the time integral of instantaneous GPU power (Equation (6)), where the instantaneous GPU power can be expressed by Equation (7). Details of MFU computation and parameter calibration are provided in Appendix D.
6.4 Results
We present quantitative comparisons across routing policies, supported by visualization of load dynamics and power consumption patterns.
Load Trajectory Analysis.
Figure 7 presents per-worker load trajectories over the processing interval. To maintain visual clarity with $G=256$ workers, we display a random sample of 16 GPUs. Under the baseline policies FCFS and JSQ, worker loads exhibit pronounced fluctuations ranging from 10M to 35M tokens during stable decoding, with high-frequency oscillations reflecting severe and persistent imbalance across the cluster. BF-IO ($H=0$) substantially compresses the load spread to approximately 12M–23M tokens. With a 40-step lookahead, BF-IO ($H=40$) achieves markedly tighter clustering (16M–17M tokens) despite the challenging workload characteristics, demonstrating the robustness of the lookahead mechanism under high-variance conditions.
Figure 7: Per-worker token load trajectories under four routing policies (16 randomly sampled workers from $G=256$). FCFS and JSQ exhibit extreme fluctuations (10M–35M tokens), whereas BF-IO policies maintain substantially tighter distributions, with BF-IO ($H=40$) achieving near-uniform loads around 16M–17M tokens.
Power Consumption Analysis.
Figure 8 illustrates the temporal dynamics of average GPU power consumption, revealing how routing policies affect both aggregate energy and instantaneous power stability.
Figure 8: Average GPU power consumption over time. BF-IO sustains consistently high power draw (395–400 W), indicating near-full utilization, and completes the workload faster. The baseline FCFS policy exhibits substantial fluctuations (270–360 W) due to load imbalance. Despite higher instantaneous power, BF-IO achieves 30% lower total energy consumption (20.9 MJ vs. 29.1 MJ) by eliminating idle time.
During stable decoding (excluding initial ramp-up), BF-IO maintains consistently high instantaneous power (395–400 W), indicating balanced utilization across all devices. For reference, the NVIDIA A100 GPU has a maximum power draw of $P_{\mathrm{max}}=400$ W (Appendix C.4), implying that BF-IO operates near full computational capacity throughout execution. In contrast, the baseline FCFS policy oscillates between 270–360 W, reflecting periods in which subsets of GPUs remain underutilized due to uneven workload distribution.
Crucially, although BF-IO operates at higher instantaneous power, it achieves substantially lower total energy consumption (20.9 MJ vs. 29.1 MJ for FCFS). This apparent paradox resolves by recognizing that total energy is the time integral of instantaneous power: by maintaining all GPUs at high utilization, BF-IO completes the workload faster, reducing total execution time and thereby the integral. In effect, balanced scheduling eliminates idle power—energy expended without productive computation—yielding significant net savings.
Quantitative Metrics Comparison.
Table 1 summarizes performance across routing policies. BF-IO consistently outperforms baselines across all metrics for every choice of lookahead window $H$ tested.
Table 1: Performance comparison on LongBench. Arrows indicate preferred direction. Best values in each column are highlighted.
[…]
($\times 10^{8}$) ($10^{3}$ tok/s) (s/tok) (MJ)
FCFS 37.0 5.47 2.97 29.1
JSQ 28.5 6.15 2.64 27.8
BF-IO ($H=0$) 9.94 8.55 1.93 23.4
BF-IO ($H=20$) 3.03 10.1 1.70 21.3
BF-IO ($H=40$) 2.49 10.5 1.67 20.6
BF-IO ($H=60$) 2.44 10.4 1.66 20.9
BF-IO ($H=80$) 2.43 10.4 1.70 21.0
BF-IO ($H=100$) 2.51 10.4 1.67 20.9
The optimal lookahead window is $H=40$. Relative to FCFS, BF-IO ($H=40$) reduces average imbalance by a factor of $15\times$, improves throughput by 92%, reduces time-per-output-token (TPOT) by 44%, and lowers energy consumption by 29%. These gains are substantial and consistent across metrics.
Comparison across BF-IO variants reveals a tradeoff in selecting $H$ (Figure 9). When $H$ is too small, the policy optimizes myopically, balancing loads only over the immediate future and missing opportunities to anticipate near-term completions. When $H$ is too large, the optimization horizon extends beyond the range of reliable prediction—since arrivals are online and unpredictable, the objective becomes biased toward stale information. Additionally, larger $H$ incurs greater computational overhead per decision. Empirically, performance saturates around $H=40$; further increases yield no benefit and marginally degrade some metrics.
Figure 9: Effect of lookahead horizon $H$ on key performance metrics. All metrics improve rapidly as $H$ increases from 0 to 40, then plateau. The optimal operating point balances prediction reliability against computational overhead; beyond $H\approx 40$, additional lookahead provides no further benefit.
6.5 Ablation Study
Scalability with Cluster Size.
To understand how BF-IO scales with increasing computational resources, we conduct an ablation study varying the number of GPU workers $G$ from 16 to 224 while keeping the workload fixed (Longbench dataset). Figures 10 and 11 present the scaling behavior of BF-IO compared to FCFS.
Figure 10: Scalability analysis with varying cluster size $G$. Left: Average imbalance grows super-linearly under FCFS but remains bounded under BF-IO. Right: BF-IO achieves near-linear throughput scaling while FCFS exhibits diminishing returns.
The left panel of Figure 10 reveals that the average imbalance under FCFS grows super-linearly with cluster size, increasing from approximately $3\times 10^{8}$ at $G=16$ to over $3.5\times 10^{9}$ at G=224G=224—a $12\times$ increase for a $7\times$ scale-up in resources. In contrast, BF-IO maintains imbalance below $3\times 10^{8}$ across all configurations, demonstrating that workload-aware scheduling effectively contains the coordination overhead even at scale. At $G=224$, BF-IO achieves $14\times$ lower imbalance than FCFS.
The right panel shows that throughput scaling exhibits fundamentally different characteristics between the two strategies. BF-IO achieves near-linear scaling, increasing from approximately 1k tokens/s at $G=16$ to over 9k tokens/s at $G=224$. Meanwhile, FCFS throughput grows sub-linearly, reaching only about 4.5k tokens/s at the largest scale—less than half of BF-IO’s throughput. This divergence stems directly from the imbalance growth: as more workers are added under FCFS, the probability of severe load skew increases, causing stragglers that bottleneck overall progress.
Figure 11: Energy consumption scaling with cluster size. BF-IO (blue) achieves consistently lower energy than FCFS (orange), with energy reduction (green, right axis) increasing from 12% at $G=16$ to 30% at $G=224$.
Figure 11 highlights the operational efficiency gains of BF-IO. While FCFS energy consumption remains relatively flat around 31–34 MJ across different cluster sizes, BF-IO energy decreases steadily from approximately 29 MJ at $G=16$ to 22 MJ at $G=224$. The energy reduction percentage (green line) increases monotonically with cluster size, rising from 12% at small scales to 30% at $G=224$. This widening gap demonstrates that the benefits of intelligent scheduling compound at larger scales: not only does BF-IO complete workloads faster, but it also avoids the energy waste from idle GPUs waiting at synchronization barriers.
7 Discussion and Conclusion
7.1 Summary of contributions
This work develops a universal load-balancing principle for barrier-synchronized systems with sticky assignment decisions. We formalize the principle as BF-IO, which selects assignments by minimizing a short-horizon prediction of the imbalance metric induced by the slowest worker via an integer optimization. We establish worst-case, scale-sensitive guarantees for BF-IO under adversarial arrivals in the LLM decode model and extend the analysis to a broader class of non-decreasing workload drift processes. Crucially, we prove that this imbalance reduction yields rigorous energy savings guarantees: as system scale grows, the guaranteed energy reduction approaches 52% for modern GPU hardware. This result establishes, to our knowledge, the first theoretical link between load balancing policy and provable energy efficiency in LLM serving. On the empirical side, extensive numerical experiments on both public benchmarks and proprietary industrial traces demonstrate that BF-IO substantially reduces imbalance, improves throughput, and lowers energy consumption relative to widely deployed baselines.
7.2 Broader relevance beyond LLM serving
Although the LLM serving bottleneck motivates our study, the abstraction captured by BF-IO is broader: many scientific and engineering workloads evolve in discrete steps, include at least one synchronization point (explicitly or implicitly), and bind work to a device or process for efficiency reasons (data locality, memory residency, or stateful execution). Two representative examples illustrate how the short-lookahead philosophy transfers.
Large-scale linear algebra and matrix computation.
Many high-performance linear algebra kernels (e.g., blocked matrix factorizations, tiled matrix multiplication, and iterative solvers) execute as task graphs over submatrices/tiles. Even when per-kernel costs are relatively stable, effective completion times of “jobs” (usually several hours) are hard to predict far in advance due to cache effects, contention, and dependence constraints. Nevertheless, short lookahead information is often readily available from the runtime: tile-based systems maintain progress counters and DAG readiness queues, and a node can infer whether it will finish a set of outstanding tile tasks within the next few scheduling quanta by tracking (i) the number of remaining tiles on the critical path, (ii) the fraction of tiles already completed in the current panel/update phase, and (iii) convergence indicators for iterative methods (e.g., residual norms approaching tolerance). In such settings, BF-IO can be applied by treating each core/GPU as a worker, using short-horizon completion signals for currently running tasks, and routing or admitting new tasks to reduce load imbalance.
Adaptive multi-physics simulations with mesh refinement.
A second example arises in physics and geoscience simulations that use adaptive mesh refinement (AMR) or adaptive particle resolution [berger1984adaptive, berger1989local]. Here the per-step workload on each rank/accelerator can drift as refinement/coarsening changes the number of cells/particles and as localized physics kernels activate (e.g., shocks, phase boundaries, or localized forcing). Long-run workload is difficult to forecast because refinement decisions depend on evolving solution features, but near-term evolution is often predictable: refinement criteria are computed at regular intervals from local error indicators, gradients, or thresholded physical diagnostics, so it is feasible to estimate whether a region will remain refined (or be coarsened) over the next few time steps. Because time stepping typically includes synchronization (global reductions, halo exchanges, or coordinated stage boundaries), the slowest rank gates progress. Under these conditions, a short-lookahead policy like BF-IO is natural: it uses near-term refinement/coarsening signals of ongoing regions to guide assignment of newly generated work units (e.g., blocks, patches, or particle batches) so as to reduce barrier-induced idle time.
These examples share the key property exploited by BF-IO: while predicting the total remaining runtime of a newly arriving unit of work is unreliable, predicting whether ongoing work will finish within a short window is often feasible from progress signals, diagnostics, or domain structure. This makes short lookahead a practical and portable information interface across scientific workloads.
7.3 Limitations and future directions
Several limitations suggest clear directions for future work.
System interfaces and buffering.
Our strongest theoretical guarantees are developed under a centralized waiting-pool abstraction. Some serving engines bind requests immediately to per-worker queues, which reduces the scheduler’s ability to reshape batches at slot-release times and can weaken future-aware balancing when queueing delays exceed the lookahead window. A promising direction is to develop a full theory for the “instant-dispatch” interface (routing to per-worker FIFO queues) and to quantify how short-horizon signals degrade as activation delays grow.
Model scope.
Our general theory covers a broad class of non-decreasing workload drift processes, motivated by stateful execution where work grows with accumulated state. Some systems exhibit decreasing or highly non-monotone per-step workloads (e.g., aggressive pruning, adaptive compression, or multi-phase jobs with alternating heavy/light stages). Extending worst-case theory to such regimes, and characterizing when monotonicity is essential, remains open.
Multi-objective serving.
Real deployments often optimize multiple objectives simultaneously (tail latency, fairness across tenants, service-level objective compliance, and energy caps). Although the BF-IO objective aligns well with barrier-induced idle reduction, integrating fairness and service-level objective constraints into the short-horizon optimization—while preserving millisecond decision budgets—deserves dedicated study.
Appendix A Supplementary Materials for Section 1
A.1 Literature Review
A.1.1 Load Balancing Problem
In this section, we briefly review classical load balancing algorithm structures [aslam2015load] (except for First-come-first-serve (FCFS) [saeed2018load, rathi2024design, tarandeep2020load], which serves as the benchmark in the paper and is discussed in Section 3) and elucidate their fundamental inefficiency or incompatibility in the LLM inference problem:
  • Join Shortest Queue (JSQ). JSQ [whitt1986deciding, dai2007stability, foley2001join] routes each arriving request to the worker with the smallest queue length (ties arbitrary). In practice, LLM serving systems in widely used (e.g., vLLM [kwon2023efficient], SGLang [zheng2024sglang]) implement a JSQ-style policy, so we take JSQ as the primary baseline. The core difficulty is that, in decode-time DP with sticky assignments, “queue length” is a poor proxy for per-step work: true workloads are unknown at assignment time and evolve as KV caches grow, so production systems typically measure the number of requests rather than their total workload. This makes JSQ brittle. Consider $G$ workers and an adversarial arrival sequence with “heavy” requests of unknown long length $L$ and “short” requests of length $s\ll L$. Because a heavy counts as one request, JSQ will place a heavy on the worker that currently has the fewest requests; subsequent bursts of many short requests inflate the counts on other workers while leaving the heavy worker’s count small, so the next heavy again goes to the same worker. After $k$ such events, that worker holds $k$ heavies while others hold many shorts. Since per-step time under the barrier model is
    $T_{\mathrm{step}}(t)=\max_{g}T^{(g)}_{\mathrm{local}}$,
    we have $T_{\mathrm{step}}^{\mathrm{JSQ}}(t)\geq kc(L)$, where $c(L)$ is the per-step local cost of a single heavy (proportional to its resident KV). A balanced assignment would yield $T_{\mathrm{step}}^{\star}(t)\leq\lceil k/G\rceil c(L)$, so the ratio is $\Omega(G)$. The problem persists because assignments are non-preemptive: as heavies generate more tokens, their per-step cost grows, yet JSQ cannot correct earlier placements. Hence JSQ targets the wrong surrogate and can be highly inefficient in the decode-stage, barrier-synchronized setting.
  • Round-Robin (RR). RR [sran2013comparative, subashini2011survey, swarnkar2013survey] dispatch assigns the $i$-th arriving (post-prefill) request to worker $((i-1)\bmod G)+1$, cycling through the $G$ workers irrespective of job size, resident KV, or drift. Under sticky, non-preemptive assignments, this determinism is exploitable. Consider an online sequence in which requests with indices $i\in{1,1+G,1+2G,\dots}$ have long decode length $L$ (heavy), and all others have short length $s\ll L$. RR places all heavy jobs on worker 1 and none on workers $2,\dots,G$. Let $H$ be the number of heavy jobs that arrive before any heavy completes (this is typical when $L$ is large). Then, for each decode step in that interval, the per-step makespan under RR satisfies
    $T_{\text{step}}^{\text{RR}}(t)\geq H\cdot c(L),$
    where $c(L)$ is the per-step local cost contributed by one heavy (proportional to its resident KV), because the max local time occurs on worker 1 carrying $H$ heavies. By contrast, an assignment that evenly spreads heavies yields
    $T_{\text{step}}^{\star}(t)\leq\Big\lceil\tfrac{H}{G}\Big\rceil\cdot c(L),$
    so the ratio $T_{\text{step}}^{\text{RR}}/T_{\text{step}}^{\star}\geq\tfrac{H}{\lceil H/G\rceil}\geq G$. This gap persists across many steps because jobs are sticky and heavy requests evolve slowly, so RR is not competitive for our LLM inference load balancing problem under adversarial arrivals.
  • Power-of-d. At each arrival, Power-of-d [gardner2017redundancy, hellemans2018power, mukherjee2018universality] samples $d\ll G$ workers uniformly at random and routes the request to the one with the smallest queue length among the $d$ sampled. This reduces coordination overhead from $O(G)$ to $O(d)$ per arrival and, in classical migratable/size-known settings, retains JSQ-like balance with strong tail guarantees. However, in decode-time DP with sticky assignments and barrier synchronization, power-of-d inherits the core JSQ shortcomings: queue length is a poor surrogate for workload when job sizes are unknown and evolve with KV growth; placements are non-preemptive, so early mistakes persist; and minimizing counts does not control the per-step maximum local time that gates progress.
  • Min-min. The Min–Min scheduler [patel2015enhanced, chen2013user, bhoi2013enhanced] builds an “earliest completion time” (ECT) matrix and repeatedly assigns the task that can finish soonest on some worker. Concretely, for each unscheduled task $i$ and worker $g$, compute $\mathrm{ECT}_{ig}=r_{g}+p_{ig}$, where $r_{g}$ is the worker’s ready time and $p_{ig}$ is the estimated processing time of task $i$ on $g$. Choose $i^{\star}=\arg\min_{i}\min_{g}\mathrm{ECT}_{ig}$, assign it to $g^{\star}=\arg\min_{g}\mathrm{ECT}_{i^{\star}g}$, update $r_{g^{\star}}$, and repeat. This strategy presumes reasonably accurate, worker-specific $p_{ig}$ and is effective for static batches on heterogeneous clusters. However, it is incompatible in our decode-time setting it is ill-suited: (i) processing times are unknown and evolve token-by-token with KV growth, so $p_{ig}$ is neither observable nor stationary; (ii) assignments are sticky (KV migration is impractical), so Min–Min’s advantage of revising choices before start or splitting jobs is unavailable; (iii) computing and maintaining a full ECT matrix is computationally prohibitive. Consequently, Min–Min cannot address the constraints of decode-stage load balancing in LLM serving.
  • Opportunistic Load Balancing and Load Balancing Min-Min (OLB+LBMM). This hybrid couples opportunistic load balancing with a load-balancing Min–Min refinement [rewehel2014new, kunwar2017load, kansal2012cloud]. OLB greedily dispatches arrivals to the first idle or least-busy worker (size-agnostic) to keep resources utilized; periodically, LBMM builds an earliest-completion-time table over the waiting pool (often after chunking tasks) and reassigns pending or not-yet-started chunks to heterogeneous workers that minimize predicted completion time. This can reduce transient idleness relative to pure Min–Min. However, it still relies on (i) accurate per-chunk processing-time estimates, (ii) the ability to split work before start, and (iii) periodic ECT recomputation overhead. Crucially, in our decode-time regime requests are indivisible: token generation is sequential and bound to a single KV cache, so pre-chunking across workers is not possible. Therefore, the OLB+LBMM architecture remains incompatible with the constraints in LLM serving.
  • Max–Min. Max–Min [nace2009max, mao2014max, ghosh2012load] is the dual of Min–Min: for each unscheduled task $i$ and worker $g$, compute the earliest–completion time $\mathrm{ECT}_{ig}=r_{g}+p_{ig}$. For each task, take its best machine $g(i)=\arg\min_{g}\mathrm{ECT}_{ig}$ with best time $m_{i}=\min_{g}\mathrm{ECT}_{ig}$. Max–Min then selects $i^{\star}=\arg\max_{i}m_{i}$ (the task whose best predicted finish time is largest), assigns it to $g(i^{\star})$, updates $r_{g(i^{\star})}$, and repeats. This choice favors long tasks early to mitigate starvation on heterogeneous clusters. In decode-time LLM serving, the same incompatibilities as Min–Min apply: $p_{ig}$ is unknown and evolves token by token (KV growth), assignments are sticky and requests are indivisible (no pre-chunking or migration), and maintaining ECT tables within millisecond is impractical. Consequently, Max–Min is not suitable for the decode-stage, barrier-synchronized LLM serving setting.
  • Ant Colony Optimization (ACO). ACO [sim2003ant, nishant2012load, li2011cloud] is a constructive metaheuristic that repeatedly builds task–worker assignments by sampling edges with probability $P_{ig}\propto\tau_{ig}^{\alpha}\eta_{ig}^{\beta}$, where $\tau_{ig}$ (pheromone) summarizes past good solutions and $\eta_{ig}$ is a heuristic desirability. After evaluating an assignment, pheromones evaporate and are reinforced, and the process iterates over many ants/iterations. In our decode-time setting ACO is incompatible for two core reasons. Information: ACO needs informative processing-time estimates $p_{ig}$, but decode workloads are unknown at assignment and drift token by token with KV growth, so the signals required to guide sampling are unreliable. Computation: Even if perfect predictions existed, ACO requires multiple ants and iterations and is typically batch oriented; with online, non-stationary arrivals and a millisecond decision budget, recomputing pheromones and re-sampling assignments per arrival is computationally infeasible. Consequently, ACO and its variants are ill-suited to decode-stage load balancing.
  • Honey Bee Foraging Algorithm (HBFA). HBFA-style schedulers [ld2013honey, thapliyal2022load] emulate bee foraging: “employed” workers advertise their current load and the number of pending priority tasks (a waggle-dance–like signal), while “onlooker” tasks probabilistically select a destination with high advertised fitness (e.g., few priority tasks, low queue length, sufficient CPU/memory). This algorithm aims to raise utilization and cut mean response time on heterogeneous clusters. In decode-time LLM serving, HBFA is incompatible for the same structural reasons as ACO: (i) it relies on queue-length/priority signals, but per-request workloads are unknown at assignment and evolve token by token; (ii) its population-style selection/updates are infeasible to a millisecond scheduling budget under online, non-stationary arrivals. Consequently, HBFA does not satisfy the information, control, and timing constraints in decode-stage load balancing.
  • Throttled Load Balancing (TLB). TLB [domanal2013load, panigrahi2020m] maintains a catalog of workers with capability vectors and enforces a per-worker concurrency threshold $\Theta$; an arriving request is routed to the first worker that is under its threshold and appears “suitable” given current residual capacity. While this capacity-aware gating can outperform RR, it is misaligned with decode-time DP. First, TLB presumes a known, stable requirement vector $d(r)$ per request, but in decode the effective demand is unknown at admission and drifts token by token as the KV cache grows, so eligibility tests based on initial size or queue length are unreliable. Second, assignments are sticky: early misplacements cannot be corrected as $d(r)$ inflates. Third, throttling caps concurrency rather than minimizing the per-step maximum local work; depending on $\Theta$, it can even increase idle time by admitting fewer short requests to otherwise underloaded workers while a heavy request gates progress. So, TLB is ineffective for decode-stage load balancing under unknown, evolving workloads and synchronous barriers.
  • Carton. Carton [hu2010scheduling, singh2015carton, ramesh2018sclba] combines load balancing with distributed rate limiting: a controller spreads incoming work across workers, while each worker enforces a local rate cap so that measured performance levels are equalized with minimal coordination. This can work well when workloads and rate signals are stable. In decode-time DP, however, Carton is structurally incompatible for two reasons. First, the signals it relies on—per-worker rates, queue lengths, or smoothed performance counters—are poor surrogates for decode-time work: per-request workloads are unknown at admission and drift token by token as KV caches grow, so rate-based estimates become stale on the step-by-step timescale. Second, assignments are sticky: once a request is placed on a decode worker, its KV cache is not migrated, so DRL cannot reallocate in-flight work to correct early imbalances. Heavy requests then continue to dominate a worker’s local time while others idle at barriers, leaving Carton’s LB+DRL control loop without the information or actuation needed to repair decode-stage imbalance.
A.1.2 LLM Serving and Energy Consumption
Recent work on the energy footprint of LLM serving falls into two complementary directions: (i) characterization and accounting of energy, and (ii) control/optimization of GPU power under service-level objectives (SLOs). On characterization, [stojkovic2024towards] profile serving across batching and parallelism knobs and show stage-specific energy–latency trade-offs; [ding2024sustainable] frame sustainability as a system-level challenge, highlighting that serving can rival or exceed training in operational energy; [nguyen2024towards] quantify both operational and embodied emissions, showing that hardware generation and batching interact non-trivially with energy efficiency; and [ozcan2025quantifying] model the fundamental power-utilization relationship, establishing that GPU power scales sublinearly with utilization—idle states draw substantial power without contributing throughput. On control, [wilkins2024offline] build workload-based energy/runtime models to derive offline energy-optimal serving policies on heterogeneous systems, while iteration-level dynamic voltage frequency scaling (DVFS) emerges as a runtime lever: [kakolyris2024slo] propose SLO-aware GPU frequency control tailored to autoregressive decoding, and [liu2025greenllm] separate prefill and decode into distinct control loops to reduce energy under trace replays without throughput loss. Together, these studies establish that serving’s energy footprint is large, that idle power is non-negligible, that prefill versus decode differ in performance–energy behavior, and that SLO-aware frequency control and hardware/batching choices are potent—but workload- and system-specific—tools
The studies above optimize per-device power or fleet-level energy/carbon via frequency control, hardware selection, or offline scheduling models. None addresses the decode-stage barrier objective where per-step time is governed by the maximum local workload across data-parallel workers—precisely the imbalance our paper targets. Our universal load-balancing principle is orthogonal and complementary: by provably shrinking the per-step maximum (thus reducing systematic idle at barriers), it converts wasted idle power into useful computation, lowering energy per token regardless of which DVFS or hardware policy is deployed. Moreover, it applies under unknown, drifting decode lengths where prediction-based controllers degrade. In short, prior work characterizes and tunes how fast and how green a given worker runs; we address how evenly the decode-time work is distributed across workers—an upstream lever that tightens the energy/latency envelope that DVFS and carbon-aware schedulers can then exploit.
A.1.3 LLM Serving Optimization
Recent systems work has rearchitected the serving stack to expose structural levers that raise throughput and reduce latency/queuing contention. Engine designs such as vLLM introduce PagedAttention to virtualize and compact the KV cache, eliminating fragmentation and enabling large, flexible batches with near-zero KV waste [kwon2023efficient]. SGLang provides a program/runtime interface for multi-call LLM applications (tools, control flow, structured I/O), co-designing execution with batching and caching to keep the accelerator pipeline busy [zheng2024sglang]. At the cluster level, prefill–decode (PD) disaggregation separates prompt encoding from autoregressive decoding across GPUs to remove cross-phase interference and to co-optimize latency targets for TTFT (prefill) and TPOT (decode) [zhong2024distserve, patel2023splitwise]. These structural choices interact with parallelism: decode commonly uses data parallelism (DP) over requests, while model parallelism mixes tensor and pipeline partitioning for large models and expert parallelism (MoE) for sparse activation [liu2024deepseek, shoeybi2019megatron, narayanan2021efficient, fedus2022switch]. Together, this body of work establishes the modern serving substrate—memory-efficient KV management, program-aware engines, PD disaggregation, and composable DP/TP/PP/EP—as the structural foundation on which scheduling and load-balancing policies operate.
Besides optimization in architecture of LLM serving, many recent works design scheduling algorithms to make LLM serving more efficient. Production LLM engines adopt iteration-level scheduling with continuous batching as the default control primitive: Orca introduced iteration-level scheduling and selective batching to keep accelerators busy despite heterogeneous request lengths [yu2022orca], and vLLM co-designs Paged-Attention with preemptive request scheduling to admit and retire sequences at iteration granularity [kwon2023efficient]. Building on these primitives, recent systems explore complementary schedulers: Sarathi-Serve proposes stall-free scheduling via chunked prefill to relieve the prefill–decode interference and improve the throughput–latency frontier [agrawal2023sarathi, agrawal2024taming]; Fast-Serve adds token-granular preemption with a skip-join multi-level feed back queue (MLFQ) to reduce job completion time under bursty arrivals [wu2023fast]. A parallel line uses learned or proxy signals to approximate shortest-first: proxy-model-based predictors to rank jobs and avoid head-of-line blocking [qiu2024efficient], learned-to-rank policies that plug into vLLM [fu2024efficient], and embedding-based estimators with preemptive schedulers [shahout2024don]. Fairness-aware policies integrate service fairness into continuous batching without preemption [sheng2024fairness]. Beyond single-instance engines, decentralized or cross-replica schedulers have been proposed for heterogeneous or edge settings [srivatsa2024preble].
A growing subset of works tackles load balancing explicitly. ShuffleInfer disaggregates prefill and decode instances and introduces a two-level scheduler to spread mixed workloads across instances, reporting latency gains but relying on instance-level migration [hu2025shuffleinfer]. For MoE inference, HarMoEny balances expert load via dynamic token redistribution and asynchronous expert prefetching across GPUs [doucet2025harmoeny], while DuoServe-MoE separates expert scheduling for prefill versus decode to match distinct activation patterns [zhang2025duoserve]. At the router layer, performance-aware load balancers combine response-length prediction with mixing strategies to steer flows across replicas [jain2025performance]. Systems surveys and empirical analyses further catalog scheduling variants and highlight the gap between academic schedulers and deployable, low-overhead policies [kim2024effect]. Many of these works target expert-parallel (EP) balancing rather than data-parallel (DP) balancing of per-request decode work; the latter is more challenging because job sizes are unknown and evolving, assignments are sticky (KV non-migratable), and step progress is gated by synchronous barriers. Among approaches that could be adapted to DP, most are engine- or trace-specific heuristics without theoretical guarantees; their effectiveness varies across datasets and configurations and thus does not constitute a universal load-balancing principle.
Moreover, a recent theory stream formalizes LLM serving as online scheduling with KV-cache constraints and provides provable policies at the single-worker level. [jaillet2025online] develop a benchmark via a hindsight-optimal integer program, prove impossibility in the fully online adversarial setting (no constant competitive ratio for any algorithm), and design algorithms that are near-optimal in a semi-online model. Following this model, [wang2025llm, chen2025adaptively] relax several assumptions in the model. They study heterogeneous prefill and unknown decode lengths with only interval prediction information. In parallel, [ao2025optimizing] introduce a fluid-guided framework with memory constraints, deriving Nested-WAIT policies that are near-optimal against a fluid benchmark in heavy traffic. These works provide principled scheduling for batching and scheduling under a single worker and do not address problems under multiple workers.
A.1.4 Assignment and Scheduling Problem
Our decode worker selection problem is an instance of the online assignment paradigm: jobs (requests) arrive sequentially, some side information may be observed, and the decision maker must irrevocably assign each job to a resource so as to optimize a system objective. Foundationally, online assignment connects to online bipartite matching and budgeted allocation (AdWords): the classical RANKING algorithm achieves the optimal (11/e)(1-1/e) competitive ratio in adversarial arrival models for bipartite matching, while primal–dual methods underlie many extensions [karp1990optimal, mehta2007adwords, devanur2009adwords, mehta2013online]. Subsequent work generalizes from linear objectives to concave returns and to learning or estimation within the primal–dual framework, as well as fairness-aware variants [jaillet2014online, gallego2015online, ma2017online, wang2017online, ma2023fairness]. Beyond the classic “offline resources fixed, online jobs arrive” setting, fully online models allow both sides to arrive and sharpen competitive limits [huang2020fully]. Parallel lines model task heterogeneity and worker qualities in crowdsourcing markets, formalizing how partial information at arrival influences assignment quality [ho2012online]. Taken together, these literatures provide baseline guarantees for irrevocable decisions under uncertainty, typically in profit/throughput objectives and without per-step synchronization.
A newer strand studies reusable resource allocation, where an assigned resource is occupied for some duration and then returns to the pool; this captures ridesharing, healthcare beds, and inventory with returns [sumita2022online]. Recent results establish near-optimal guarantees in adversarial or stochastic arrivals by exploiting fluid or learning approximations and carefully designed dual controls [goyal2020online, zhang2022online, huo2022online, ao2024two]. These models are closer to our setting (decode workers are reusable) but still differ in two crucial ways: (i) they optimize cumulative reward (linear) while our objective is the load balancing (non-linear); and (ii) they typically assume either known or stationary service-time models, whereas decode workloads are unknown at admission and evolve token-by-token. Thus, classical online assignment and reusable-resource frameworks inform admission/routing under uncertainty, but they do not directly address data-parallel decode with sticky assignments and synchronous barriers that gate progress each step.
The classical online scheduling literature studies how a decision maker allocates processing time to jobs that arrive sequentially, aiming to optimize latency/throughput objectives under limited lookahead [chen1998review, brucker1999resource, xing2000parallel, handbookSched, allahverdi2008survey, kong2013scheduling, mak2015appointment, beyondWorstCase]. Two strands are especially relevant to LLM inference on a single engine. First, batch scheduling groups jobs to amortize setup or memory costs and to raise effective utilization; models and algorithms range from static batch formation to fully online batching with competitive guarantees [brucker1998scheduling, chen2008logistics, lucier2013efficient, im2013online, li2020online]. Second, precedence-constrained scheduling captures intra-job dependencies; here each task must respect a partial order, and objectives include makespan and total weighted completion time [schedPrecedence, precSchedSchabanel, precSchedBen, precSchedAnupam]. LLM decoding on a single worker fits this mold: tokens within a request obey a strict chain precedence (each token depends on its prefix), batching trades latency for throughput, and memory (KV-cache) constraints couple sequencing and admission. Results from these lines provide principled policies for admission, batching, and service order when processing times are known or learnable, but they operate at the level of one engine and do not account for cross-worker synchronization.
Multi-processor scheduling generalizes to identical/related/unrelated machines and mixes preemptive and non-preemptive models [handbookSched, brucker1999resource, xing2000parallel, kong2013scheduling]. Canonical approaches (e.g., list scheduling, longest-processing-time rules, or migration-based balancing) assume either known processing times or allow job preemption/migration to correct early placements; performance is typically analyzed for makespan or flow-time objectives. These models illuminate when per-machine queues equalize and when migration is essential, yet they remain misaligned with LLM decode in two ways: (i) unknown and evolving processing times undermine size-aware or learned dispatching; (ii) sticky assignments remove the main lever used to fix early mistakes. Classic multi-worker scheduling therefore informs local ordering and (where allowed) migration, but it does not address the barrier-synchronized, unknown-size, non-preemptive regime that governs data-parallel decoding; this gap motivates our universal load-balancing principle.
A.1.5 Online Algorithms under Adversarial Arrivals
An online algorithm chooses actions sequentially as inputs arrive, without seeing the future; an adversarial arrival model allows the input sequence to be chosen by an adversary, often either oblivious (fixes the entire sequence before seeing the algorithm’s random bits) or adaptive (may react to the algorithm’s past random choices). Performance is assessed by competitive analysis: finding out the largest gap between the performance of the online algorithm and the hindsight optimal algorithm (which has full access to all arrivals in advance).
Beyond our online assignment domain, a broad literature studies online algorithms under adversarial arrivals. In online resource allocation, booking-limit and protection-level policies [ball2009toward, golrezaei2023online], together with primal–dual methods [agrawal2014dynamic, li2020simple, li2022online, jiang2025online], yield constant or logarithmic competitive ratios for covering/packing and related network-allocation problems [buchbinder2009online]. In online caching and paging, classic results show least-recently-used (LRU) is $k$-competitive, and randomized marking schemes achieve $O(\log k)$ competitiveness against an oblivious adversary [sleator1985amortized, fiat1991competitive, lykouris2021competitive]. The $k$-server problem and metrical task systems capture stateful movement on metric spaces; the work-function algorithm attains (2k1)(2k{-}1)-competitiveness for $k$-server on general metrics [koutsoupias1995k, borodin1992optimal]. In online convex optimization, regret-minimizing methods (e.g., online gradient and mirror descent) guarantee $O(\sqrt{T})$ adversarial regret [zinkevich2003online]. In online facility location, Meyerson’s randomized algorithm is $O(\log n)$-competitive in random-order models, whereas under oblivious adversaries the tight rate is $\Theta(\log n/\log\log n)$ [meyerson2001online, fotakis2008competitive]. Our analysis likewise adopts an adversarial-arrival viewpoint—inputs (arrivals and lengths) are determined by an oblivious adversary—and establishes competitive optimality for decode-stage load balancing under synchronous barriers.
Appendix B Baseline Policy: First-Come-First-Serve (FCFS)
In this subsection, we formalize the baseline policy used throughout the paper: First-Come-First-Serve (FCFS). FCFS is widely deployed in real LLM serving systems’ scheduler [kwon2023efficient, zheng2024sglang].
At each step $k$, the scheduler maintains a waiting queue $R_{\text{wait}}(k)$ that stores all requests that have arrived by time $k$ but have not yet been assigned to any worker. Under FCFS, the scheduler scans workers in increasing index $g=1,2,\ldots,G$ and fills any available slots on each worker up to the batch capacity $B$. Requests are always taken from $R_{\text{wait}}(k)$ in strict arrival order (the head of the queue), and no request that arrived later may be scheduled before an earlier one. The capacity constraint $\lvert\mathcal{A}_{g}(k)\rvert\leq B$ is always respected, and the loop terminates as soon as either all workers are full (no remaining capacity) or the waiting queue is empty.
This policy is extremely fast and deterministic (for a fixed arrival sequence), and it does not rely on any information about the workload profile $W_{i}$ of individual requests. Precisely because it ignores workload heterogeneity, FCFS can easily accumulate many heavy requests on a single worker while others receive only light ones, leading to large per-step imbalance and, consequently, substantial idle time at the barrier. The pseudocode used in our analysis and implementation is given in Algorithm 2.
Input: Waiting queue $R_{\text{wait}}(k)$ (arrival order), worker concurrency $B$, active sets $\{\mathcal{A}_{g}(k)\}_{g\in[G]}$
Output: Assignments at step $k$
1exCompute remaining slots on each worker: $\texttt{cap}[g]\leftarrow B-\lvert\mathcal{A}_{g}(k)\rvert$ for all $g\in[G]$;
if $\sum_{g\in[G]}\texttt{cap}[g]=0\ \mathbf{or}\ R_{\text{wait}}(k)=\emptyset$ then
 return
while $R_{\!\text{wait}}(k)$ is not None do
   Pop the oldest request $i$ from $R_{\!\text{wait}}(k)$;
   Select $g^{*}\in\mathop{\arg\max}_{g}\texttt{cap}[g]$ with maximal free slots;
   Assign $i$ to worker $g^{*}$ : $\mathcal{A}_{g^{*}}(k)\leftarrow\mathcal{A}_{g^{*}}(k)\cup\{i\}$;
 $\texttt{cap}[g^{*}]\leftarrow\texttt{cap}[g^{*}]-1$;
Algorithm 2 FCFS Baseline
Appendix C Supplementary Materials for Section 5
C.1 Proof of Theorem 1
Proof of Theorem 1:.
We first reduce the statement to a single admission round, and then analyze BF-IO and FCFS separately.
Reduction to a single admission round. In the homogeneous-output setting $o_{i}=o$ for all $i$, the evolution naturally decomposes into rounds. At the beginning of a round, the cluster is empty; $GB$ jobs are admitted (up to capacity), with $B$ jobs sent to each worker. All admitted jobs have the same decode length $o$, so they generate tokens in lockstep and terminate simultaneously after $o$ steps. No further jobs are admitted until all $GB$ jobs complete, at which point the next round begins.
Within a round, the instantaneous imbalance is the same at each of the $o$ steps, because all workers advance in parallel, and the per-step workloads merely translate in time. Thus, the time-average imbalance over a long horizon is proportional to the average imbalance per admission round. Formally, let $\mathrm{Imb}^{(r)}(\pi)$ be the (step-wise) imbalance in round $r$ under a policy $\pi$ immediately after the $GB$ jobs of that round have been admitted. As argued above, this quantity also captures the imbalance at each step within the round, up to the fixed multiplicative factor $o$.
Under the prefill-length model in Section 5, and given any arrival instance $I$ that satisfies the overloaded condition at every step, the multiset of prefill lengths entering each round consists of $GB$ i.i.d. draws from the distribution of $s$; different rounds use disjoint sets of jobs, and so the collections of $GB$ prompts across rounds are independent. In particular, the random variables $\{\mathrm{Imb}^{(r)}(\pi)\}_{r\geq 1}$ are i.i.d. for each fixed policy $\pi$, and their law does not depend on the specific overloaded arrival instance $I$.
Let $\text{AvgImbalance}(\pi;I)$ denote the time-average imbalance under policy $\pi$ for arrival instance $I$ as in Section 3. Over a long horizon, the law of large numbers implies that, for any $\pi$ and any $I\in\mathcal{I}$,
$\lim_{K\to\infty}\mathbb{E}\bigl[\text{AvgImbalance}(\pi;I)\bigr]\;=\;\mathbb{E}\big[\mathrm{Imb}^{(1)}(\pi)\big],$
because each round contributes $o$ identical steps, and the rounds form an i.i.d. sequence. Consequently, the imbalance improvement ratio $\mathbf{IIR}$ simplifies to the ratio of one-round expectations:
$\mathbf{IIR}\;=\;\inf_{I\in\mathcal{I}}\frac{\lim_{K\to\infty}\mathbb{E}[\text{AvgImbalance}(\textsf{FCFS};I)]}{\lim_{K\to\infty}\mathbb{E}[\text{AvgImbalance}(\textsf{BF-IO};I)]}\;=\;\frac{\mathbb{E}\big[\mathrm{Imb}^{(1)}(\textsf{FCFS})\big]}{\mathbb{E}\big[\mathrm{Imb}^{(1)}(\textsf{BF-IO})\big]}.$
Thus, it suffices to lower bound the ratio
$\frac{\mathbb{E}[\mathrm{Imb}(\textsf{FCFS})]}{\mathbb{E}[\mathrm{Imb}(\textsf{BF-IO})]}$
in a single admission round, where we write $\mathrm{Imb}(\pi)$ for $\mathrm{Imb}^{(1)}(\pi)$ for brevity.
Step A: BF-IO is $s_{\max}$-balanced. Fix a realization of the $GB$ prefill lengths in the current round. Since prefill lengths are independent across jobs and arrival times do not carry additional information about their values, we may regard these as an unordered multiset of $GB$ numbers in $0,s_{\max}]$. Any assignment that fills all devices in the round corresponds to a partition of this multiset into $G$ disjoint groups $S_{1},\dots,S_{G}$, each of cardinality $|S_{g}|=B$. Define the per-device loads
$L_{g}:=\sum_{s\in S_{g}}s,\qquad M:=\max_{1\leq g\leq G}L_{g},\qquad m:=\min_{1\leq g\leq G}L_{g}.$
Because all jobs have identical decode length $o$, the per-step imbalance within the round is constant and proportional to the deviation of the $L_{g}$ from $M$. In particular, minimizing the sum of future imbalances over the round is equivalent to minimizing the maximum load $M$ across devices in that round. Let $\{S_{g}^{\star}\}_{g=1}^{G}$ be any assignment that minimizes $M$.
Lemma 1.
For any optimal assignment $\{S_{g}^{\star}\}$ minimizing $M$, we have
$M-m\;\leq\;s_{\max}.$
Proof of Lemma 1.
Suppose for contradiction that $M-m>s_{\max}$. Let $p\in\arg\max_{g}L_{g}$ and $q\in\arg\min_{g}L_{g}$, so $L_{p}=M$ and $L_{q}=m$. Because $L_{p}-L_{q}>M-m>s_{\max}$ and every element of $S_{p}^{\star}\cup S_{q}^{\star}$ lies in $0,s_{\max}]$, there must exist $x\in S_{p}^{\star}$ and $y\in S_{q}^{\star}$ with $x>y$; otherwise we would have $\max S_{p}^{\star}\leq\min S_{q}^{\star}$, which together with $|S_{p}^{\star}|=|S_{q}^{\star}|=B$ would imply
$L_{p}\;\leq\;B\max_{s\in S_{p}^{\star}}s\;\leq\;B\min_{s\in S_{q}^{\star}}s\;\leq\;L_{q},$
contradicting $L_{p}>L_{q}$.
Construct a new assignment by swapping $x$ and $y$ between devices $p$ and $q$:
$S_{p}^{\prime}:=(S_{p}^{\star}\setminus\{x\})\cup\{y\},\qquad S_{q}^{\prime}:=(S_{q}^{\star}\setminus\{y\})\cup\{x\},$
and $S_{r}^{\prime}:=S_{r}^{\star}$ for all $r\notin\{p,q\}$. The new loads are
$L_{p}^{\prime}=L_{p}-(x-y)<L_{p},\qquad L_{q}^{\prime}=L_{q}+(x-y)\leq L_{q}+s_{\max}<L_{p},$
and $L_{r}^{\prime}=L_{r}$ for $r\notin\{p,q\}$. Thus every new load $L_{r}^{\prime}$ is strictly less than the old maximum $L_{p}=M$ except possibly those devices $r\neq p$ that already had $L_{r}=M$ before the swap. In particular, if $p$ was the unique maximizer, then
$\max_{g}L_{g}^{\prime}\;<\;\max_{g}L_{g}\;=\;M,$
which contradicts the optimality of $\{S_{g}^{\star}\}$ as a minimizer of $M$.
If instead there were multiple devices achieving the maximum $M$ before the swap, then after the swap, device $p$ falls below $M$, while all other devices $r\neq p$ maintain their previous loads $L_{r}^{\prime}\leq M$. Hence
$\max_{g}L_{g}^{\prime}\;=\;M,$
but the number of devices attaining this maximum strictly decreases. Consequently, starting from any optimal assignment that is also lexicographically minimal in the pair
$\big(\,\max_{g}L_{g},\;\#\{g:L_{g}=M\}\,\big),$
we cannot have $M-m>s_{\max}$, since the above swap produces another assignment with either a smaller maximum load or the same maximum load but fewer maximizers. This contradicts lexicographic minimality. Therefore $M-m\leq s_{\max}$ must hold for every optimal assignment. ∎
For any assignment, the round-level imbalance is
$\mathrm{Imb}\;=\;\sum_{g=1}^{G}(M-L_{g}).$
For the optimal assignment $\{S_{g}^{\star}\}$, choose $p\in\arg\max_{g}L_{g}$ and note that $L_{p}=M$. Then
$\mathrm{Imb}\;=\;\sum_{g\neq p}(M-L_{g})\;\leq\;\sum_{g\neq p}(M-m)\;=\;(G-1)(M-m)\;\leq\;(G-1)\,s_{\max}$
by Lemma 1. Since this bound holds for every realization of the $GB$ prompt lengths in the round, taking expectations yields
$\mathbb{E}\big[\mathrm{Imb}(\textsf{BF-IO})\big]\;\leq\;(G-1)\,s_{\max}.$ (23)
Step B: FCFS lower bound. We now derive a lower bound on the expected imbalance under FCFS in one round. Because of the overloaded assumption and size-agnostic FCFS admission, the $GB$ jobs admitted in a round can be viewed as an unordered multiset of i.i.d. prefill lengths with law $s$, after which FCFS fills the $G$ devices in arrival order. Conditioned on this multiset, the distribution of the $B$ prompts assigned to each device is that of $B$ samples drawn uniformly without replacement from the multiset. Since the underlying pool itself comprises $GB$ i.i.d. draws from $s$, a standard exchangeability argument implies that the prompts $(s_{g,j})_{j=1}^{B}$ on each device $g$ can be represented as $B$ i.i.d. draws from $s$, and that the collections $\{(s_{g,j})_{j=1}^{B}:g\in[G]\}$ are independent across $g$.
Let
$S_{g}:=\sum_{j=1}^{B}s_{g,j},\qquad\mu:=\mathbb{E}[s],\qquad\sigma_{s}^{2}:=\mathrm{Var}(s)>0.$
Then $S_{1},\dots,S_{G}$ are i.i.d. with
$\mathbb{E}[S_{g}]=\mu B,\qquad\mathrm{Var}(S_{g})=\sigma_{s}^{2}B.$
Define the standardized sums
$Z_{g}:=\frac{S_{g}-\mu B}{\sigma_{s}\sqrt{B}},\qquad g=1,\dots,G.$
Because $s\in[0,s_{\max}]$, the random variable $s$ has a bounded third absolute moment, and the Berry–Esseen theorem (for sums of i.i.d. real variables) yields the existence of a universal constant $C_{\mathrm{BE}}$ such that
$\sup_{x\in\mathbb{R}}\Big|\Pr(Z_{g}\leq x)-\Phi(x)\Big|\;\leq\;\frac{C_{\mathrm{BE}}}{\sqrt{B}},$ (24)
where $\Phi$ is the standard normal distribution function.
Set $z_{G}:=\sqrt{\frac{\log G}{2}}$ and consider the threshold $z_{G}$. For a standard normal $Z\sim N(0,1)$ and large $G$, the Gaussian tail bound implies
$\Pr(Z\geq z_{G})\;\geq\;\frac{1}{\sqrt{2\pi}}\,\Big(\frac{1}{z_{G}}-\frac{1}{z_{G}^{3}}\Big)\,\exp\!\Big(-\frac{z_{G}^{2}}{2}\Big)\;\geq\;\frac{1}{2\sqrt{2\pi}\,z_{G}}\,G^{-1/4}.$
Combining this with (24), we obtain
$\Pr(Z_{g}\geq z_{G})\;\geq\;\frac{1}{2\sqrt{2\pi}\,z_{G}}\,G^{-1/4}\;-\;\frac{C_{\mathrm{BE}}}{\sqrt{B}}.$
Since $\sqrt{G}\log G=o(B)$, there exists $B_{0}(G)$ such that for all $B\geq B_{0}(G)$,
$\frac{C_{\mathrm{BE}}}{\sqrt{B}}\;\leq\;\frac{1}{2}\cdot\frac{1}{2\sqrt{2\pi}\,z_{G}}\,G^{-1/4},$
and hence
$\Pr(Z_{g}\geq z_{G})\;\geq\;\frac{1}{4\sqrt{2\pi}\,z_{G}}\,G^{-1/4}\;=:\;p_{G}.$ (25)
By independence across $g$, we have
$\Pr\Big(\max_{1\leq g\leq G}Z_{g}\geq z_{G}\Big)\;\geq\;1-\big(1-p_{G}\big)^{G}.$
Using the inequality $1-x\leq e^{-x}$ for $x\geq 0$, we find
[…]
From (25), $Gp_{G}\geq\frac{1}{4\sqrt{2\pi}\,z_{G}}G^{3/4}$, which is bounded away from zero for every fixed $G\geq 2$. Thus, there exists a universal constant $c_{1}\in(0,1)$ (for instance $c_{1}:=1-e^{-1/(4\sqrt{2\pi})}$) such that
$\Pr\Big(\max_{1\leq g\leq G}Z_{g}\geq z_{G}\Big)\;\geq\;c_{1}\qquad\text{for all }\log G\geq 4\text{ and all }B\geq B_{0}(G).$ (26)
We now translate this into a lower bound on $\mathbb{E}[\max_{g}S_{g}]$. For any $t>0$, an elementary inequality yields
$\mathbb{E}\Big[\max_{1\leq g\leq G}S_{g}\Big]=\mu B+\mathbb{E}\Big[\max_{1\leq g\leq G}(S_{g}-\mu B)\Big]\;\geq\;\mu B+t\;\Pr\Big(\max_{g}(S_{g}-\mu B)\geq t\Big).$
Choose $t:=\sigma_{s}\sqrt{B}z_{G}$, so that the event $\{\max_{g}(S_{g}-\mu B)\geq t\}$ coincides with $\{\max_{g}Z_{g}\geq z_{G}\}$. Combining with (26), we obtain
$\mathbb{E}\Big[\max_{g}S_{g}\Big]\;\geq\;\mu B+\sigma_{s}\sqrt{B}z_{G}\,c_{1}\;=\;\mu B+\frac{c_{1}}{\sqrt{2}}\;\sigma_{s}\,\sqrt{B\log G}.$
Since $\mathbb{E}[\sum_{g=1}^{G}S_{g}]=G\mu B$, it follows that the expected imbalance in the round under FCFS satisfies
$\mathbb{E}\big[\mathrm{Imb}(\textsf{FCFS})\big]=G\,\mathbb{E}\Big[\max_{g}S_{g}\Big]-\mathbb{E}\Big[\sum_{g=1}^{G}S_{g}\Big]\;\geq\;c_{2}\,G\,\sigma_{s}\,\sqrt{B\log G},$ (27)
for all $B\geq B_{0}(G)$, where $c_{2}:=c_{1}/\sqrt{2}>0$ is a universal constant.
Conclusion. Combining the upper bound (23) for BF-IO with the lower bound (27) for FCFS, we find that for all $B\geq B_{0}(G)$,
$\frac{\mathbb{E}[\mathrm{Imb}(\textsf{FCFS})]}{\mathbb{E}[\mathrm{Imb}(\textsf{BF-IO})]}\;\geq\;\frac{c_{2}\,G\,\sigma_{s}\,\sqrt{B\log G}}{(G-1)\,s_{\max}}\;=\;\Big(c_{2}\,\frac{\sigma_{s}}{s_{\max}}\Big)\,\sqrt{B\log G}\cdot\frac{G}{G-1}.$
By the non-degeneracy assumption in Section 5, the ratio $\sigma_{s}/s_{\max}$ is bounded below by $\kappa_{0}>0$. Setting $c:=c_{2}$ and recalling that $\mathbf{IIR}$ equals the ratio of one-round expectations, we conclude that
$\mathbf{IIR}\;\geq\;c\,\kappa_{0}\,\sqrt{B\log G}\cdot\frac{G}{G-1},$
which is the claimed lower bound (3). This completes the proof. ∎
C.2 Proof of Theorem 2
Proof of Theorem 2:.
We work on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ that supports the arrival instance $\mathcal{I}$, the prefill lengths $\mathcal{D}_{\mathrm{prefill}}$, the geometric decode lengths $\mathcal{D}_{\mathrm{decode}}=\text{Geo}(p)$, and the completion indicators. For each step $k\in[K]$, we define the natural filtration
[…]
and write $\mathbb{E}[\cdot\mid\mathcal{F}_{k}]$ for conditional expectations.
At the end of step k1k-1 (after BF-IO admission at that step), each device $g\in\{1,\dots,G\}$ holds exactly $B$ active requests, indexed by $i=1,\dots,B$, with sizes $a_{g,i}(k-1)\in\mathbb{N}$ (including prompt tokens and already generated output tokens). Let
$L_{g}(k-1)\ :=\ \sum_{i=1}^{B}a_{g,i}(k-1),\qquad D(k-1)\ :=\ \max_{g}L_{g}(k-1)\;-\;\min_{g}L_{g}(k-1),$
and
$\Lambda(k)\ :=\ \max_{g,i}a_{g,i}(k-1),$
be, respectively, the device loads, the inter-device gap, and the one-step envelope at the end of step k1k-1. These quantities are $\mathcal{F}_{k-1}$-measurable.
At the beginning of step $k$, every active request grows by exactly one token, then completes independently with probability $p$ at the end of the step. Let $\xi_{g,i}(k)\sim\mathrm{Bernoulli}(p)$ be the completion indicator of slot $(g,i)$ in step $k$, independent of $\mathcal{F}_{k-1}$ and across $(g,i)$. The number of completions (free slots) on device $g$ in step $k$ is
$c_{g}(k)\ :=\ \sum_{i=1}^{B}\xi_{g,i}(k)\ \sim\ \mathrm{Binomial}(B,p),$
independent across $g$ and of $\mathcal{F}_{k-1}$, and we write $C_{k}:=\sum_{g=1}^{G}c_{g}(k)$ for the total number of completions.
Before admitting new requests at step $k$, the total pre-admission load on device $g$ is
$a_{g}^{\mathrm{pre}}(k)\ :=\ L_{g}(k-1)+B-R_{g}(k),\qquad R_{g}(k)\ :=\ \sum_{i=1}^{B}\xi_{g,i}(k)\big(a_{g,i}(k-1)+1\big),$
where $R_{g}(k)$ is the load removed by completions (each surviving job contributes +1+1 token). The corresponding pre-admission spread is
$A^{\mathrm{pre}}(k)\ :=\ \max_{g}a_{g}^{\mathrm{pre}}(k)\;-\;\min_{g}a_{g}^{\mathrm{pre}}(k).$
After BF-IO admission at step $k$, we write $L_{g}(k)$ for the post-admission loads and
$D(k)\ :=\ \max_{g}L_{g}(k)\;-\;\min_{g}L_{g}(k)$
for the post-admission gap. Finally, we set
$\Lambda(k+1)\ :=\ \max_{g,i}a_{g,i}(k),$
the one-step envelope at the end of step $k$. By construction, $L_{g}(k)$, $D(k)$, and $\Lambda(k+1)$ are $\mathcal{F}_{k}$-measurable.
Part 1 (Upper Bound of Imbalance of BF-IO)
Step 1.A: Separation and one-step drift
The first lemma is a geometric statement about the BF-IO assignment at step $k$.
Lemma 2 (Separation and gap reduction).
Fix $\delta\in\big(0,(2s_{\max}-1)^{-1}\big)$ and set
$\gamma\ :=\ 1-\delta(2s_{\max}-1)\ \in\ (0,1).$
Then, for every $k\geq 1$,
$\mathbb{P}\Big(D(k)\ \leq\ \max\big\{s_{\max},\,A^{\mathrm{pre}}(k)-\gamma Bp\big\}\,\Big|\,\mathcal{F}_{k-1}\Big)\ \geq\ 1-2G\exp\Big(-\tfrac{\delta^{2}}{3}Bp\Big).$ (28)
Proof of Lemma 2:.
Fix a step $k$ and condition on $\mathcal{F}_{k-1}$. Given $\mathcal{F}_{k-1}$, the collection of sizes $\{a_{g,i}(k-1)\}$ and loads $\{L_{g}(k-1)\}$ is deterministic, and the randomness at step $k$ comes only from the completion indicators $\{\xi_{g,i}(k)\}$ and the prompt lengths of newly admitted requests. Let
$E_{\mathrm{cnt}}(k)\ :=\ \big\{|c_{g}(k)-Bp|\leq\delta Bp\ \ \text{for all }g\in\{1,\dots,G\}\big\}.$
The event $E_{\mathrm{cnt}}(k)$ depends only on $\{\xi_{g,i}(k)\}$, and thus is independent of $\mathcal{F}_{k-1}$. By the Chernoff bound for binomial variables,
$\mathbb{P}\big(E_{\mathrm{cnt}}(k)^{c}\mid\mathcal{F}_{k-1}\big)=\mathbb{P}\big(E_{\mathrm{cnt}}(k)^{c}\big)\ \leq\ 2G\exp\Big(-\tfrac{\delta^{2}}{3}Bp\Big).$
On the event $E_{\mathrm{cnt}}(k)$, we have
$c_{\min}(k)\ :=\ \min_{g}c_{g}(k)\ \geq\ (1-\delta)Bp,\qquad c_{\max}(k)\ :=\ \max_{g}c_{g}(k)\ \leq\ (1+\delta)Bp.$
Fix an outcome $\omega\in E_{\mathrm{cnt}}(k)$ and consider deterministically the BF-IO assignment at step $k$ on that outcome. Let $S_{g}(k)$ denote the multiset of prompt sizes of the new requests admitted to device $g$ at step $k$; by construction $|S_{g}(k)|=c_{g}(k)$, and
$L_{g}(k)\;=\;a_{g}^{\mathrm{pre}}(k)+\sum_{s\in S_{g}(k)}s.$
We first establish a separation property for the minimizer of $D(k)$. For each feasible family of multisets $\{S_{g}(k)\}_{g=1}^{G}$ with $|S_{g}(k)|=c_{g}(k)$ and using only jobs from the waiting pool, define
$L_{g}(k)\ :=\ a_{g}^{\mathrm{pre}}(k)+\sum_{s\in S_{g}(k)}s,\qquad D(k)\ :=\ \max_{g}L_{g}(k)-\min_{g}L_{g}(k).$
Among all such assignments, BF-IO chooses one that minimizes $D(k)$. Let $\{S_{g}^{\star}(k)\}$ be such a minimizer, and let
$L_{g}^{\star}(k)\ :=a_{g}^{\mathrm{pre}}(k)+\sum_{s\in S_{g}^{\star}(k)}s,\qquad D^{\star}(k)\ :=\max_{g}L_{g}^{\star}(k)-\min_{g}L_{g}^{\star}(k),$
denote the corresponding loads and gap. If $D^{\star}(k)\leq s_{\max}$, then $D(k)=D^{\star}(k)\leq s_{\max}$ and the claimed inequality (28) trivially holds on this outcome. Hence it suffices to analyze the case $D^{\star}(k)>s_{\max}$.
Let $p\in\arg\max_{g}L_{g}^{\star}(k)$ and $q\in\arg\min_{g}L_{g}^{\star}(k)$, so that
$L_{p}^{\star}(k)-L_{q}^{\star}(k)\ =\ D^{\star}(k)\ >\ s_{\max}.$
We first show that, under this strict gap and the optimality of $\{S_{g}^{\star}(k)\}$, the assignment must be “separated” between any such heavy–light pair $(p,q)$. That is, for the sets $S_{p}^{\star}(k)$ and $S_{q}^{\star}(k)$ there exists an integer $\theta\in\{0,\dots,s_{\max}-1\}$ such that
$s\leq\theta\quad\text{for all }s\in S_{p}^{\star}(k),\qquad s\geq\theta+1\quad\text{for all }s\in S_{q}^{\star}(k).$
We prove by contradiction. Suppose not, then there exist $x\in S_{p}^{\star}(k)$ and $y\in S_{q}^{\star}(k)$ with $x>y$. Consider the assignment $\{\widetilde{S}_{g}(k)\}$ obtained by swapping $x$ and $y$ between devices $p$ and $q$,
$\widetilde{S}_{p}(k):=(S_{p}^{\star}(k)\setminus\{x\})\cup\{y\},\qquad\widetilde{S}_{q}(k):=(S_{q}^{\star}(k)\setminus\{y\})\cup\{x\},$
and $\widetilde{S}_{r}(k):=S_{r}^{\star}(k)$ for all $r\notin\{p,q\}$. This swap preserves $|S_{g}(k)|=c_{g}(k)$ for every $g$ and uses only jobs from the waiting pool (by the overloaded definition 1, the pool contains sufficiently many jobs of each size to support such exchanges). The new loads are
$\widetilde{L}_{p}(k)\;=\;a_{p}^{\mathrm{pre}}(k)+\sum_{s\in\widetilde{S}_{p}(k)}s\;=\;L_{p}^{\star}(k)-(x-y)\;<\;L_{p}^{\star}(k),$
and
$\widetilde{L}_{q}(k)\;=\;a_{q}^{\mathrm{pre}}(k)+\sum_{s\in\widetilde{S}_{q}(k)}s\;=\;L_{q}^{\star}(k)+(x-y)\;\leq\;L_{q}^{\star}(k)+s_{\max}\;<\;L_{p}^{\star}(k),$
using $x-y\leq s_{\max}$ and $L_{p}^{\star}(k)-L_{q}^{\star}(k)>s_{\max}$. All other loads $\widetilde{L}_{r}(k)$ equal $L_{r}^{\star}(k)\leq L_{p}^{\star}(k)$.
If $p$ is the unique maximizer of $\{L_{g}^{\star}(k)\}$, then
$\max_{g}\widetilde{L}_{g}(k)<\max_{g}L_{g}^{\star}(k),$
contradicting the minimality of $D^{\star}(k)$. If there are multiple maximizers, then the swap strictly reduces the number of devices attaining the maximum load while keeping the maximum value at most $L_{p}^{\star}(k)$. Repeating this argument finitely many times, we either strictly reduce the maximum load or obtain a contradiction to the choice of $\{S_{g}^{\star}(k)\}$ as a minimizer with minimal number of maximizers. Hence no such pair $(x,y)$ can exist, and the desired threshold $\theta$ exists.
Condition on $E_{\mathrm{cnt}}(k)$ and fix an outcome with $D^{\star}(k)>s_{\max}$. Using the separation property, for device $p$ every $s\in S_{p}^{\star}(k)$ satisfies $s\leq\theta$, while for $q$ every $s\in S_{q}^{\star}(k)$ satisfies $s\geq\theta+1$. Therefore,
$\sum_{s\in S_{p}^{\star}(k)}s\;\leq\;\theta\,c_{p}(k),\qquad\sum_{s\in S_{q}^{\star}(k)}s\;\geq\;(\theta+1)\,c_{q}(k),$
and hence
$D^{\star}(k)$ $=\ L_{p}^{\star}(k)-L_{q}^{\star}(k)$
$=\ \bigg(a_{p}^{\mathrm{pre}}(k)+\sum_{s\in S_{p}^{\star}(k)}s\bigg)-\bigg(a_{q}^{\mathrm{pre}}(k)+\sum_{s\in S_{q}^{\star}(k)}s\bigg)$
$\leq\ a_{p}^{\mathrm{pre}}(k)-a_{q}^{\mathrm{pre}}(k)+\theta c_{p}(k)-(\theta+1)c_{q}(k).$
Using $c_{p}(k)\leq c_{\max}(k)$ and $c_{q}(k)\geq c_{\min}(k)$, we obtain
$\theta c_{p}(k)-(\theta+1)c_{q}(k)\;$ $\leq\;\theta c_{\max}(k)-(\theta+1)c_{\min}(k)\;\leq\;\theta(1+\delta)Bp-(\theta+1)(1-\delta)Bp$
$=Bp\big(\theta(1+\delta)-(\theta+1)(1-\delta)\big)=Bp\big(-1+\delta(2\theta+1)\big)$
$\leq Bp\big(-1+\delta(2s_{\max}-1)\big)=-\gamma Bp,$
where we used $\theta\leq s_{\max}-1$ and the definition of $\gamma$. Consequently,
$D^{\star}(k)\leq a_{p}^{\mathrm{pre}}(k)-a_{q}^{\mathrm{pre}}(k)-\gamma Bp\leq A^{\mathrm{pre}}(k)-\gamma Bp.$
Combining this with the trivial bound $D^{\star}(k)\leq s_{\max}$ in the case $D^{\star}(k)\leq s_{\max}$, and recalling that $D(k)=D^{\star}(k)$ under BF-IO, we obtain, for every outcome in $E_{\mathrm{cnt}}(k)$,
[…]
Since $E_{\mathrm{cnt}}(k)$ is independent of $\mathcal{F}_{k-1}$, taking conditional probability given $\mathcal{F}_{k-1}$ yields
$\mathbb{P}\Big(D(k)\ \leq\ \max\big\{s_{\max},\,A^{\mathrm{pre}}(k)-\gamma Bp\big\}\,\Big|\,\mathcal{F}_{k-1}\Big)\;\geq\;\mathbb{P}\big(E_{\mathrm{cnt}}(k)\,\big|\,\mathcal{F}_{k-1}\big)\;\geq\;1-2G\exp\Big(-\tfrac{\delta^{2}}{3}Bp\Big),$
as claimed. ∎
Step 1.B: Concentration of pre-admission spread
We now control $A^{\mathrm{pre}}(k)$ around its conditional baseline (1p)D(k1)(1-p)D(k-1).
Define the range functional $r(x):=\max_{g}x_{g}-\min_{g}x_{g}$ for $x\in\mathbb{R}^{G}$. For step $k$, recall that
$R_{g}(k)\;=\;\sum_{i=1}^{B}\xi_{g,i}(k)\big(a_{g,i}(k-1)+1\big),\qquad\mu_{g}(k)\ :=\ \mathbb{E}\big[R_{g}(k)\,\big|\,\mathcal{F}_{k-1}\big]=p\sum_{i=1}^{B}(a_{g,i}(k-1)+1),$
and set
$Z_{g}(k)\;:=\ R_{g}(k)-\mu_{g}(k),\qquad Z(k):=(Z_{g}(k))_{g=1}^{G}.$
The one-step envelope
$\Lambda(k):=\max_{g,i}a_{g,i}(k-1)$
is $\mathcal{F}_{k-1}$-measurable and bounds every summand in $R_{g}(k)$.
Lemma 3 (One-step concentration of $A^{\mathrm{pre}}(k)$).
For each step $k\geq 1$ and every $\mathcal{F}_{k-1}$-measurable $\Lambda(k)$, define
$\alpha_{B}:=3\log(GB),\qquad u_{A}(k):=8\,\Lambda(k)\,\max\Big\{\sqrt{p(1-p)\,B\,\alpha_{B}},\ \alpha_{B}\Big\}.$
Then, for all $k\geq 1$,
$\mathbb{P}\Big(A^{\mathrm{pre}}(k)-(1-p)D(k-1)>u_{A}(k)\,\Big|\,\mathcal{F}_{k-1}\Big)\ \leq\ 2G\,e^{-\alpha_{B}}\ \leq\ 2\,G^{-2}B^{-3}.$ (29)
Proof of Lemma 3:.
Given $\mathcal{F}_{k-1}$, the vector of loads $L(k-1)=(L_{g}(k-1))_{g=1}^{G}$ is deterministic, and
$a_{g}^{\mathrm{pre}}(k)=L_{g}(k-1)+B-R_{g}(k)=(1-p)\big(L_{g}(k-1)+B\big)\;-\;Z_{g}(k).$
Let $X_{g}:=(1-p)(L_{g}(k-1)+B)$ and $X:=(X_{g})_{g=1}^{G}$. Then
$A^{\mathrm{pre}}(k)=r\big(a^{\mathrm{pre}}(k)\big)=r(X-Z(k)).$
For any vectors $x,z\in\mathbb{R}^{G}$, comparing coordinates achieving the maximum and minimum of xzx-z gives the elementary inequality
$r(x-z)\;\leq\;r(x)+2\|z\|_{\infty},$
where $\|z\|_{\infty}:=\max_{g}|z_{g}|$. Applying this with $x=X$ and $z=Z(k)$, and using r(X)=(1p)D(k1)r(X)=(1-p)D(k-1), we obtain
$A^{\mathrm{pre}}(k)-(1-p)D(k-1)=r(X-Z(k))-r(X)\ \leq\ 2\|Z(k)\|_{\infty}=2\max_{g}|Z_{g}(k)|.$
Thus, for any $u>0$,
$\mathbb{P}\Big(A^{\mathrm{pre}}(k)-(1-p)D(k-1)>u\,\Big|\,\mathcal{F}_{k-1}\Big)\ \leq\ \sum_{g=1}^{G}\mathbb{P}\Big(|Z_{g}(k)|>u/2\,\Big|\,\mathcal{F}_{k-1}\Big).$
It remains to control each tail $\mathbb{P}(|Z_{g}(k)|>u/2\mid\mathcal{F}_{k-1})$. Conditional on $\mathcal{F}_{k-1}$, the variables $\{(\xi_{g,i}(k)-p)(a_{g,i}(k-1)+1)\}_{i=1}^{B}$ are independent, mean-zero, and bounded in absolute value by $\Lambda(k)+1\leq\,2\Lambda(k)$. Their conditional variance satisfies
$\mathrm{Var}\Big(\sum_{i=1}^{B}(\xi_{g,i}(k)-p)(a_{g,i}(k-1)+1)\,\Big|\,\mathcal{F}_{k-1}\Big)=\sum_{i=1}^{B}p(1-p)\,(a_{g,i}(k-1)+1)^{2}\leq p(1-p)\,B\,4\Lambda(k)^{2}.$
By Bernstein’s inequality for bounded summands, for every $g$ and $u>0$,
$\mathbb{P}\Big(|Z_{g}(k)|>u/2\,\Big|\,\mathcal{F}_{k-1}\Big)\ \leq\ 2\exp\left(-\min\left\{\frac{(u/2)^{2}}{16\,p(1-p)\,B\,\Lambda(k)^{2}},\ \frac{3\,(u/2)}{8\,\Lambda(k)}\right\}\right).$
Summing over $g=1,\dots,G$ and recalling the earlier inequality, we obtain
$\mathbb{P}\Big(A^{\mathrm{pre}}(k)-(1-p)D(k-1)>u\,\Big|\,\mathcal{F}_{k-1}\Big)\ \leq\ 2G\exp\left(-\min\left\{\frac{(u/2)^{2}}{16\,p(1-p)\,B\,\Lambda(k)^{2}},\ \frac{3\,(u/2)}{8\,\Lambda(k)}\right\}\right).$
Choosing
$u=u_{A}(k)=8\,\Lambda(k)\,\max\Big\{\sqrt{p(1-p)\,B\,\alpha_{B}},\ \alpha_{B}\Big\},$
forces
$\frac{(u/2)^{2}}{16\,p(1-p)\,B\,\Lambda(k)^{2}}\ \geq\ \alpha_{B},\qquad\frac{3\,(u/2)}{8\Lambda(k)}\ \geq\ \alpha_{B},$
so that the minimum in the exponent is at least $2\alpha_{B}$. This yields
$\mathbb{P}\Big(A^{\mathrm{pre}}(k)-(1-p)D(k-1)>u_{A}(k)\,\Big|\,\mathcal{F}_{k-1}\Big)\ \leq\ 2Ge^{-\alpha_{B}}.$
With $\alpha_{B}=3\log(GB)$ we have $2Ge^{-\alpha_{B}}\leq 2G^{-2}B^{-3}$, completing the proof. ∎
Step 1.C: Average gap bound for BF-IO
We now combine Lemmas 2 and 3 to obtain a uniform bound on the long-run average of $D(k)$. Recall that at each step, the instantaneous imbalance under any policy satisfies
$\mathrm{Imbalance}(k;\textsf{BF-IO})=\sum_{g=1}^{G}\big(\max_{h}L_{h}(k)-L_{g}(k)\big)\ \leq\ (G-1)\Big(\max_{h}L_{h}(k)-\min_{h}L_{h}(k)\Big)=(G-1)\,D(k).$
Lemma 4 (Average gap under geometric outputs for BF-IO).
Let $\delta:=1/(4s_{\max})$, set $\gamma:=1-\delta(2s_{\max}-1)\in(0,1)$, and let $\alpha_{B}:=3\log(GB)$. Define
$x_{\star}(B,G):=\Big\lceil\frac{3}{p}\,\log(GB)\Big\rceil,\qquad U_{B}:=8\,(s_{\max}+x_{\star}(B,G))\,\max\Big\{\sqrt{p(1-p)\,B\,\alpha_{B}},\ \alpha_{B}\Big\},$
and
$\varepsilon_{B}:=2G\,e^{-(\delta^{2}/3)Bp}+2\,G^{-2}B^{-3}.$
Take $B$ is large enough (depending on $(p,s_{\max},G)$) so that
$\gamma Bp\ \geq\ 2U_{B}.$ (30)
Then, under the overloaded-arrival regime, the BF-IO policy with $H=0$ satisfies
$\limsup_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[D(k)]\ \leq\ \frac{s_{\max}}{p}\ +\ o_{B,G}(1),$ (31)
where $o_{B,G}(1)\geq 0$ depends only on $(B,G,p,s_{\max})$ and $o_{B,G}(1)\to 0$ whenever $B,G\to\infty$ with $\sqrt{G}=o(B)$.
Proof of Lemma 4:.
Fix $B,G$, and simplify the notation $x_{\star}:=x_{\star}(B,G)$. For each step $k$ define the events
$E_{\mathrm{cnt}}(k):=\big\{|c_{g}(k)-Bp|\leq\delta Bp\ \ \forall g\big\},\quad E_{\mathrm{snap}}(k):=\{\Lambda(k)\leq s_{\max}+x_{\star}\},$
and
$E_{\mathrm{dev}}(k):=\big\{A^{\mathrm{pre}}(k)-(1-p)D(k-1)\leq U_{B}\big\}.$
By Lemma 2,
$\mathbb{P}\big(E_{\mathrm{cnt}}(k)^{c}\mid\mathcal{F}_{k-1}\big)\leq 2G\exp\Big(-\tfrac{\delta^{2}}{3}Bp\Big),$
and by Lemma 3 (applied with $x_{\star}$ and using $\Lambda(k)\leq s_{\max}+x_{\star}$ on $E_{\mathrm{snap}}(k)$),
$\mathbb{P}\big(E_{\mathrm{dev}}(k)^{c}\cap E_{\mathrm{snap}}(k)\mid\mathcal{F}_{k-1}\big)\leq 2\,G^{-2}B^{-3}.$
Abbreviate
$\varepsilon_{B}:=2G\exp\Big(-\tfrac{\delta^{2}}{3}Bp\Big)+2\,G^{-2}B^{-3}.$
On the “good” event
$E_{\mathrm{good}}(k):=E_{\mathrm{cnt}}(k)\cap E_{\mathrm{snap}}(k)\cap E_{\mathrm{dev}}(k),$
we combine Lemma 2 and the bound on $A^{\mathrm{pre}}(k)$ to obtain
$D(k)\leq\max\Big\{s_{\max},\,(1-p)D(k-1)+U_{B}-\gamma Bp\Big\}.$
As $\gamma Bp\geq 2U_{B}$, we have
$(1-p)D(k-1)+U_{B}-\gamma Bp\leq(1-p)D(k-1)-U_{B}\leq(1-p)D(k-1),$
and hence
$D(k)\ \leq\ s_{\max}+(1-p)D(k-1)\qquad\text{on }E_{\mathrm{good}}(k).$ (32)
On the complement $E_{\mathrm{good}}(k)^{c}$, we use a uniform bound. At the beginning of step $k$, before completions, every alive request grows by one token and new arrivals have size at most $s_{\max}$, so
$\Lambda(k+1)\leq\max\{\Lambda(k)+1,\ s_{\max}\}\leq s_{\max}+\big(\Lambda(k)+1-s_{\max}\big)_{+}.$
Therefore,
$B\,\Lambda(k+1)\leq B\,(s_{\max}+x_{\star})+B\big(\Lambda(k)+1-s_{\max}-x_{\star}\big)_{+},$
and since each device holds $B$ requests after admission, we have
$D(k)\leq B\,\Lambda(k+1)\qquad\text{on }E_{\mathrm{good}}(k)^{c}.$
Combining the two cases and using the union bound $\mathbf{1}_{E_{\mathrm{good}}(k)^{c}}\leq\mathbf{1}_{E_{\mathrm{cnt}}(k)^{c}}+\mathbf{1}_{E_{\mathrm{snap}}(k)^{c}}+\mathbf{1}_{E_{\mathrm{dev}}(k)^{c}\cap E_{\mathrm{snap}}(k)}$, we obtain the pathwise decomposition
$D(k)$ $\leq\big(s_{\max}+(1-p)D(k-1)\big)\,\mathbf{1}_{E_{\mathrm{good}}(k)}$
$\quad+B\,(s_{\max}+x_{\star})\big(\mathbf{1}_{E_{\mathrm{cnt}}(k)^{c}}+\mathbf{1}_{E_{\mathrm{dev}}(k)^{c}\cap E_{\mathrm{snap}}(k)}+\mathbf{1}_{E_{\mathrm{snap}}(k)^{c}}\big)+B\big(\Lambda(k)+1-s_{\max}-x_{\star}\big)_{+}.$
Taking conditional expectations given $\mathcal{F}_{k-1}$ and using the bounds on event probabilities, we have
$\mathbb{E}\big[D(k)\mid\mathcal{F}_{k-1}\big]$ $\leq\big(s_{\max}+(1-p)D(k-1)\big)$
$\quad+\varepsilon_{B}B\,(s_{\max}+x_{\star})+B\,(s_{\max}+x_{\star})\,\mathbb{P}\big(E_{\mathrm{snap}}(k)^{c}\mid\mathcal{F}_{k-1}\big)$
$\quad+B\,\mathbb{E}\big[(\Lambda(k)+1-s_{\max}-x_{\star})_{+}\mid\mathcal{F}_{k-1}\big].$ (33)
We use the simple bound, $\mathbf{1}_{\{\Lambda(k)>s_{\max}+x_{\star}\}}\leq(\Lambda(k)+1-s_{\max}-x_{\star})_{+}$, and hence
$\mathbb{P}\big(E_{\mathrm{snap}}(k)^{c}\mid\mathcal{F}_{k-1}\big)\leq\mathbb{E}\big[(\Lambda(k)+1-s_{\max}-x_{\star})_{+}\mid\mathcal{F}_{k-1}\big].$
Substituting into (33) and taking unconditional expectations yields the one-step inequality
$\mathbb{E}[D(k)]$ $\leq\Big(s_{\max}+(1-p)\mathbb{E}[D(k-1)]\Big)+\varepsilon_{B}B\,(s_{\max}+x_{\star})$
$+B\,(s_{\max}+x_{\star}+1)\,\mathbb{E}\big[(\Lambda(k)+1-s_{\max}-x_{\star})_{+}\big].$ (34)
To complete the proof, we bound the horizon average of the envelope term. In any step $k$, the size of an alive request equals the sum of its prompt length (at most $s_{\max}$) and the number of tokens it has generated up to step $k$ (its age). Thus $\Lambda(k)-s_{\max}$ is bounded by the maximal age among alive requests at step $k$. Summing over the horizon and exchanging the order of summation (first over requests, then over the steps they are alive), we obtain
$\sum_{k=0}^{K-1}\big(\Lambda(k)+1-s_{\max}-x_{\star}\big)_{+}\ \leq\ \sum_{j}\sum_{r=0}^{o_{j}-1}(r+1-x_{\star})_{+},$
where the outer sum is over all requests $j$ whose lifetime intersects [0:K1][0{:}K-1], and $o_{j}$ is the (geometric) output length of request $j$. For a single geometric random variable $o_{j}\sim\mathrm{Geo}(p)$,
$\mathbb{E}\Big[\sum_{r=0}^{o_{j}-1}(r+1-x_{\star})_{+}\Big]=\sum_{r=x_{\star}+1}^{\infty}(r-x_{\star})\mathbb{P}(o_{j}\geq r)=\sum_{r=x_{\star}+1}^{\infty}(r-x_{\star})(1-p)^{r-1}=\frac{(1-p)^{x_{\star}}}{p^{2}}.$
Let $N_{K}$ be the number of distinct requests whose lifetime intersects the horizon [0:K1][0{:}K-1]. At each step there are exactly $GB$ alive items, and each item counted in $N_{K}$ is alive for at least one step, so deterministically $N_{K}\leq GBK$ and hence $N_{K}/K\leq GB$. By linearity of expectation,
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\big[(\Lambda(k)+1-s_{\max}-x_{\star})_{+}\big]\leq\frac{\mathbb{E}\big[N_{K}\big]}{K}\cdot\frac{(1-p)^{x_{\star}}}{p^{2}}\leq GB\,\frac{(1-p)^{x_{\star}}}{p^{2}}.$
Averaging (34) over $k=0,\dots,K-1$ and rearranging, we obtain
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[D(k)]$ $\leq\Big(s_{\max}+(1-p)\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[D(k-1)]\Big)$
$\quad+\varepsilon_{B}B\,(s_{\max}+x_{\star})+\frac{GB^{2}}{p^{2}}\,(s_{\max}+x_{\star}+1)\,(1-p)^{x_{\star}}.$ (35)
The sequence $\{\mathbb{E}[D(k)]\}_{k\geq 0}$ is nonnegative, so its Cesàro average is bounded, and the right-hand side of (35) is an affine contraction in the horizon average. Letting $K\to\infty$ and using (1p)<1(1-p)<1, we obtain
$\limsup_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[D(k)]\ \leq\ \frac{s_{\max}+\varepsilon_{B}B\,(s_{\max}+x_{\star})+\frac{GB^{2}}{p^{2}}\,(s_{\max}+x_{\star}+1)\,(1-p)^{x_{\star}}}{p}.$
Finally, for $x_{\star}(B,G)=\lceil 3\log(GB)/p\rceil$, the term $(1-p)^{x_{\star}}$ decays fast in $GB$; more precisely,
$(1-p)^{x_{\star}}\leq\exp(-px_{\star})\leq(GB)^{-3},$
so
$\frac{GB^{2}}{p^{2}}\,(s_{\max}+x_{\star}+1)\,(1-p)^{x_{\star}}=o_{B,G}(1)$
whenever $B,G\to\infty$. Similarly, by definition of $\varepsilon_{B}$ and the assumption $\log G=o(B)$,
$\varepsilon_{B}B\,(s_{\max}+x_{\star})=o_{B,G}(1).$
Absorbing these vanishing terms into $o_{B,G}(1)$, we obtain (31). This completes the proof of Part 1. ∎
Part 2 (Lower Bound of Imbalance of FCFS)
We now analyze the lower bound of average imbalance of the FCFS baseline. Under the overloaded arrival definition, FCFS is size-agnostic: whenever a completion occurs in any slot, the slot is immediately refilled by the earliest pending request, whose prompt length is i.i.d. distributed by the distribution $S=\mathcal{D}_{\text{prefill}}$ with $\mathbb{E}[S]=\mu_{s}$ and $\text{Var}(S)=\sigma_{s}^{2}>0$.
Step 2.A: Slot-level recursion and Markov structure.
Index slots by $(g,b)$ with $g\in\{1,\dots,G\}$ and $b\in\{1,\dots,B\}$. Let $Y_{g,b}(k)$ denote the post-admission size in slot $(g,b)$ at the end of step $k$. Under FCFS, we have the slot-level recursion
$Y_{g,b}(k)\;=\;(1-\xi_{g,b}(k))\big(Y_{g,b}(k-1)+1\big)+\xi_{g,b}(k)\,S_{g,b}(k),\qquad k\geq 1,$ (36)
where, for each $(g,b,k)$,
$\xi_{g,b}(k)\sim\mathrm{Bernoulli}(p),\qquad S_{g,b}(k)\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}S=\mathcal{D}_{\text{prefill}},$
and all these random variables are mutually independent across $(g,b,k)$ and independent of the initial configuration $\{Y_{g,b}(0)\}$. Thus, for each fixed slot $(g,b)$, the process $(Y_{g,b}(k))_{k\geq 0}$ is a time-homogeneous Markov chain on $\mathbb{N}$, and the chains corresponding to different slots evolve independently.
For each device $g$, the post-admission load at step $k$ is
$S_{g}(k):=\sum_{b=1}^{B}Y_{g,b}(k),$
and the instantaneous FCFS imbalance at step $k$ is
$\mathrm{Imbalance}(k;\textsf{FCFS}):=G\cdot\max_{1\leq g\leq G}S_{g}(k)-\sum_{g=1}^{G}S_{g}(k).$
By symmetry across devices, for any $k$ and any initial configuration,
$\mathbb{E}\big[\mathrm{Imbalance}(k;\textsf{FCFS})\big]=G\Big(\mathbb{E}\big[\max_{g}S_{g}(k)\big]-\mathbb{E}\big[S_{1}(k)\big]\Big).$
Step 2.B: Single-slot stationary law.
Fix a slot $(g,b)$ and suppress the indices. The chain $(Y_{k})_{k\geq 0}$ defined by (36) satisfies, for every measurable $A\subseteq\mathbb{N}$ and $y\in\mathbb{N}$,
$\mathbb{P}(Y_{1}\in A\mid Y_{0}=y)\ =\ \mathbb{P}\big((1-\xi_{1})(y+1)+\xi_{1}S_{1}\in A\big)\ \geq\ \mathbb{P}(\xi_{1}=1,S_{1}\in A)\ =\ p\,\mathbb{P}(S\in A).$
Thus the transition kernel admits a Doeblin minorization with parameter $p>0$. In particular, $(Y_{k})_{k\geq 0}$ is uniformly ergodic and admits a unique stationary distribution, which we denote by $U$. To identify $U$ explicitly, it is convenient to track the age process:
[…]
Each time a completion occurs ($\xi_{k}=1$), a fresh job with prompt $S_{k}\sim S$ starts in that slot and the age resets to zero; otherwise, the job survives and the age increases by one. The age process alone evolves via
$A_{k+1}=\begin{cases}0,&\text{with probability }p,\\ A_{k}+1,&\text{with probability }1-p.\end{cases}$
This is a Markov chain on $\mathbb{N}$ with transition kernel
$\mathbb{P}(A_{k+1}=0\mid A_{k}=a)=p,\qquad\mathbb{P}(A_{k+1}=a+1\mid A_{k}=a)=1-p.$
A direct calculation shows that the geometric distribution
[…]
is invariant for the age chain:
$\mathbb{P}(A_{k+1}=a)=p\,\mathbb{P}(A_{k}\geq a)=p\,\sum_{j=a}^{\infty}p(1-p)^{j}=p(1-p)^{a}=\mathbb{P}(A=a).$
Since the prompts $\{S_{k}\}$ are i.i.d. and independent of the completions $\{\xi_{k}\}$, in stationarity the prompt of the current job $S_{\mathrm{cur}}$ and the age $A$ are independent, with $S_{\mathrm{cur}}\stackrel{{\scriptstyle d}}{{=}}S$ and $A\sim\mathrm{Geo}(p)-1$. At a stationary snapshot, the slot size $U$ has the representation
$U\ \overset{d}{=}\ S_{\mathrm{cur}}+A,\qquad S_{\mathrm{cur}}\stackrel{{\scriptstyle d}}{{=}}S,\quad A\sim\mathrm{Geo}(p)-1,\quad S_{\mathrm{cur}}\perp\!\!\!\perp A.$
Consequently,
$\mu_{U}:=\mathbb{E}[U]=\mu_{s}+\frac{1-p}{p},\qquad\sigma_{\mathrm{snap}}^{2}:=\text{Var}(U)=\sigma_{s}^{2}+\frac{1-p}{p^{2}}.$ (37)
Because $S$ is bounded and $A$ has geometric tails, $U$ has finite moments of all orders, in particular $\mathbb{E}[|U|^{3}]<\infty$.
Step 2.C: Device-level loads at stationarity.
Since the slot chains are independent and identically distributed, and their transition kernels do not couple different slots, the stationary distribution of the entire collection $\{Y_{g,b}(k)\}_{g,b}$ is the product measure under which all $Y_{g,b}$ are i.i.d. with law $U$. Let $(U_{g,b})_{g,b}$ be i.i.d. copies of $U$ and define the stationary per-device loads
$S_{g}^{\mathrm{st}}:=\sum_{b=1}^{B}U_{g,b},\qquad g=1,\dots,G.$
Then $\{S_{g}^{\mathrm{st}}\}_{g=1}^{G}$ are i.i.d. with
$\mathbb{E}[S_{g}^{\mathrm{st}}]=\mu_{U}B,\qquad\mathrm{Var}(S_{g}^{\mathrm{st}})=\sigma_{\mathrm{snap}}^{2}B.$
We now invoke the same Berry–Esseen and extreme-value argument as in the homogeneous-output case, but with $U$ replacing $s$. Define the standardized sums
$W_{g}:=\frac{S_{g}^{\mathrm{st}}-\mu_{U}B}{\sigma_{\mathrm{snap}}\sqrt{B}},\qquad g=1,\dots,G.$
Because $U$ has finite third absolute moment, the Berry–Esseen theorem (for sums of i.i.d. real variables) guarantees the existence of a universal constant $C_{\mathrm{BE}}$ such that
$\sup_{x\in\mathbb{R}}\Big|\mathbb{P}(W_{g}\leq x)-\Phi(x)\Big|\;\leq\;\frac{C_{\mathrm{BE}}}{\sqrt{B}},$
for each $g$, where $\Phi$ is the standard normal distribution function.
Set $z_{G}:=\sqrt{\frac{\log G}{2}}$ and consider the threshold $z_{G}$. For a standard normal $Z\sim N(0,1)$ and all sufficiently large $G$ (say $\log G\geq 4$), the Gaussian tail bound implies
$\mathbb{P}(Z\geq z_{G})\;\geq\;\frac{1}{\sqrt{2\pi}}\,\Big(\frac{1}{z_{G}}-\frac{1}{z_{G}^{3}}\Big)\,\exp\!\Big(-\frac{z_{G}^{2}}{2}\Big)\;\geq\;\frac{1}{2\sqrt{2\pi}\,z_{G}}\,G^{-1/4}.$
Combining this with Berry–Esseen, we obtain
$\mathbb{P}\big(W_{g}\geq z_{G}\big)\;\geq\;\frac{1}{2\sqrt{2\pi}\,z_{G}}\,G^{-1/4}\;-\;\frac{C_{\mathrm{BE}}}{\sqrt{B}}.$
Since $\sqrt{G}=\tilde{o}(B)$, there exists $B_{0}(G)$ such that, for all $B\geq B_{0}(G)$,
$\frac{C_{\mathrm{BE}}}{\sqrt{B}}\;\leq\;\frac{1}{2}\cdot\frac{1}{2\sqrt{2\pi}\,z_{G}}\,G^{-1/4},$
and hence
$\mathbb{P}\big(W_{g}\geq z_{G}\big)\;\geq\;\frac{1}{4\sqrt{2\pi}\,z_{G}}\,G^{-1/4}\;=:\;p_{G}.$ (38)
By independence across $g$, we have
$\mathbb{P}\Big(\max_{1\leq g\leq G}W_{g}\geq z_{G}\Big)\;\geq\;1-\big(1-p_{G}\big)^{G}.$
Using the inequality $1-x\leq e^{-x}$ for $x\geq 0$, we find
[…]
From (38), $Gp_{G}\geq\frac{1}{4\sqrt{2\pi}\,z_{G}}G^{3/4}$, which is bounded away from zero for every fixed $G\geq 2$. Thus, there exists a universal constant $c_{1}\in(0,1)$ (for instance $c_{1}:=1-e^{-1/(4\sqrt{2\pi})}$) such that
$\mathbb{P}\Big(\max_{1\leq g\leq G}W_{g}\geq z_{G}\Big)\;\geq\;c_{1}\qquad\text{for all }\log G\geq 4\text{ and all }B\geq B_{0}(G).$
Finally, for any $t>0$,
$\mathbb{E}\Big[\max_{1\leq g\leq G}S_{g}^{\mathrm{st}}\Big]=\mu_{U}B+\mathbb{E}\Big[\max_{1\leq g\leq G}(S_{g}^{\mathrm{st}}-\mu_{U}B)\Big]\;\geq\;\mu_{U}B+t\,\mathbb{P}\big(\max_{g}(S_{g}^{\mathrm{st}}-\mu_{U}B)\geq t\big).$
Choosing $t:=\sigma_{\mathrm{snap}}\sqrt{B}\,z_{G}$, so that the event $\{\max_{g}(S_{g}^{\mathrm{st}}-\mu_{U}B)\geq t\}$ coincides with $\{\max_{g}W_{g}\geq z_{G}\}$, we obtain
$\mathbb{E}\Big[\max_{1\leq g\leq G}S_{g}^{\mathrm{st}}\Big]\;\geq\;\mu_{U}B+\sigma_{\mathrm{snap}}\sqrt{B}\,z_{G}\,c_{1}\;=\;\mu_{U}B+c^{\prime}\,\sigma_{\mathrm{snap}}\,\sqrt{B\log G},$
for some absolute constant $c^{\prime}>0$. Since $\mathbb{E}[\sum_{g=1}^{G}S_{g}^{\mathrm{st}}]=G\mu_{U}B$, the stationary expected imbalance under FCFS satisfies
$\mathbb{E}\big[\mathrm{Imbalance}^{\mathrm{st}}\big]:=\mathbb{E}\Big[G\max_{g}S_{g}^{\mathrm{st}}-\sum_{g=1}^{G}S_{g}^{\mathrm{st}}\Big]\;\geq\;c^{\prime}\,G\,\sigma_{\mathrm{snap}}\sqrt{B\log G},$ (39)
for all $G\geq 2$ and all $B\geq B_{0}(G)$, where $B_{0}(G)$ is as above.
Step 2.D: Time-average imbalance under FCFS.
The global collection of slot loads
$Y(k):=\big(Y_{g,b}(k)\big)_{g\in[G],\,b\in[B]}$
is a Markov chain on $\mathbb{N}^{GB}$ with transition kernel induced by (36). By the Doeblin minorization on each coordinate, the product chain is uniformly ergodic and admits the unique stationary distribution $\Pi$, which is precisely the product law of i.i.d. copies of $U$. Moreover, by uniform ergodicity, for any initial configuration $Y(0)$ and any integrable function $f$,
$\lim_{k\to\infty}\mathbb{E}\big[f(Y(k))\big]=\mathbb{E}_{\Pi}[f(Y)].$
Applying this to $f(Y)=\mathrm{Imbalance}(k;\textsf{FCFS})$, which has finite second moment under $\Pi$ in view of (37), we deduce that
$\lim_{k\to\infty}\mathbb{E}\big[\mathrm{Imbalance}(k;\textsf{FCFS})\big]=\mathbb{E}\big[\mathrm{Imbalance}^{\mathrm{st}}\big].$
By Cesàro averaging, it follows that the long-run expected time-average imbalance under FCFS,
$\text{AvgImbalance}(\textsf{FCFS};\mathcal{I}):=\lim_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\big[\mathrm{Imbalance}(k;\textsf{FCFS})\big],$
exists and equals the stationary expectation $\mathbb{E}\big[\mathrm{Imbalance}^{\mathrm{st}}\big]$, independently of the arrival instance $\mathcal{I}$ and the initial configuration. Combining with (39), we obtain
$\text{AvgImbalance}(\textsf{FCFS};\mathcal{I})\;\geq\;c^{\prime}\,G\,\sigma_{\mathrm{snap}}\sqrt{B\log G},\qquad\sigma_{\mathrm{snap}}^{2}=\sigma_{s}^{2}+\frac{1-p}{p^{2}},$ (40)
for all $G\geq 2$ and all $B$ sufficiently large, uniformly over all overloaded arrival instances $\mathcal{I}$.
Part 3: BF-IO vs. FCFS
We now combine the BF-IO average-gap bound from Lemma 4 with the FCFS lower bound (40) to obtain the imbalance–reduction rate in the geometric-output setting. By Lemma 4, for any overloaded arrival instance $\mathcal{I}$ we have
$\limsup_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[D(k)]\ \leq\ \frac{s_{\max}}{p}\ +\ o_{B,G}(1),\qquad\text{with }o_{B,G}(1)\to 0\ \text{as }\ B,G\to\infty.$
Since, for every step $k$ and any policy,
$\mathrm{Imbalance}(k;\textsf{BF-IO})=\sum_{g=1}^{G}\big(\max_{h}L_{h}(k)-L_{g}(k)\big)\ \leq\ (G-1)\,D(k),$
it follows that
$\mathrm{AvgImbalance}(\textsf{BF-IO};\mathcal{I})=\limsup_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\big[\mathrm{Imbalance}(k;\textsf{BF-IO})\big]\ \leq\ (G-1)\,\frac{s_{\max}}{p}\ +\ o_{B,G}(1),$
with the same $o_{B,G}(1)$.
On the other hand, by (40), for every overloaded arrival instance $\mathcal{I}$,
$\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I})\ \geq\ c^{\prime}\,G\,\sigma_{\mathrm{snap}}\sqrt{B\log G},\qquad\sigma_{\mathrm{snap}}^{2}=\sigma_{s}^{2}+\frac{1-p}{p^{2}},$
for some universal constant $c^{\prime}>0$ and all $B$ sufficiently large. Therefore, for every such $\mathcal{I}$,
$\frac{\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I})}{\mathrm{AvgImbalance}(\textsf{BF-IO};\mathcal{I})}\ \gtrsim\ \frac{G\,\sigma_{\mathrm{snap}}\sqrt{B\log G}}{(G-1)\,s_{\max}/p}\ =\ c\,\frac{p\,\sigma_{\mathrm{snap}}}{s_{\max}}\cdot\frac{G}{G-1}\,\sqrt{B\log G},$
for a possibly smaller universal constant $c>0$, where we absorb the $o_{B,G}(1)$ term into $c$ for $B$ large enough. Since $\sigma_{\mathrm{snap}}\geq\sqrt{1-p}/p$, the prefactor $\frac{p\,\sigma_{\mathrm{snap}}}{s_{\max}}$ is strictly positive (up to the fixed scale $s_{\max}$), and the scaling $\sqrt{B\log G}$ is inherited from the FCFS lower bound. Taking the infimum over all overloaded arrival instances $\mathcal{I}$ completes the proof.
C.3 Proof of Theorem 3
Proof of Theorem 3.
We work on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ that supports the arrival instance $\mathcal{I}$, the prefill lengths $\mathcal{D}_{\mathrm{prefill}}$, the geometric decode lengths $\mathcal{D}_{\mathrm{decode}}=\mathrm{Geo}(p)$, and the completion indicators. For each step $k\in\mathbb{N}$ we define the natural filtration
[…]
and write $\mathbb{E}[\cdot\mid\mathcal{F}_{k}]$ for conditional expectations.
At the end of step k1k-1 (after BF-IO admission at that step), each device $g\in\{1,\dots,G\}$ holds exactly $B$ active requests, indexed by $i=1,\dots,B$, with sizes $a_{g,i}(k-1)\in\mathbb{N}$ (including prompt tokens and already generated output tokens). Let
$L_{g}(k-1)\ :=\ \sum_{i=1}^{B}a_{g,i}(k-1),\qquad D(k-1)\ :=\ \max_{g}L_{g}(k-1)\;-\;\min_{g}L_{g}(k-1),$
and
$\Lambda(k)\ :=\ \max_{g,i}a_{g,i}(k-1),$
be, respectively, the device loads, the inter-device gap, and the one-step envelope at the end of step k1k-1. These quantities are $\mathcal{F}_{k-1}$-measurable.
At the beginning of step $k$, every active request grows by exactly $\delta_{k}$ units of workload and then completes independently with probability $p$ at the end of the step. Let $\xi_{g,i}(k)\sim\mathrm{Bernoulli}(p)$ be the completion indicator of slot $(g,i)$ in step $k$, independent of $\mathcal{F}_{k-1}$ and across $(g,i)$. The number of completions (free slots) on device $g$ in step $k$ is
$c_{g}(k)\ :=\ \sum_{i=1}^{B}\xi_{g,i}(k)\ \sim\ \mathrm{Binomial}(B,p),$
independent across $g$ and of $\mathcal{F}_{k-1}$, and we write $C_{k}:=\sum_{g=1}^{G}c_{g}(k)$ for the total number of completions.
Before admitting new requests at step $k$, the total pre-admission load on device $g$ is
$a_{g}^{\mathrm{pre}}(k)\ :=\ L_{g}(k-1)+B\delta_{k}-R_{g}(k),\qquad R_{g}(k)\ :=\ \sum_{i=1}^{B}\xi_{g,i}(k)\big(a_{g,i}(k-1)+\delta_{k}\big),$
where $R_{g}(k)$ is the load removed by completions in step $k$. The corresponding pre-admission spread is
$A^{\mathrm{pre}}(k)\ :=\ \max_{g}a_{g}^{\mathrm{pre}}(k)\;-\;\min_{g}a_{g}^{\mathrm{pre}}(k).$
After BF-IO admission at step $k$, we write $L_{g}(k)$ for the post-admission loads and
$D(k)\ :=\ \max_{g}L_{g}(k)\;-\;\min_{g}L_{g}(k)$
for the post-admission gap. Finally, we set
$\Lambda(k+1)\ :=\ \max_{g,i}a_{g,i}(k),$
the one-step envelope at the end of step $k$. By construction, $L_{g}(k)$, $D(k)$, and $\Lambda(k+1)$ are $\mathcal{F}_{k}$-measurable.
Part 1 (Upper Bound of Imbalance of BF-IO)
Step 1.A: Separation and one-step drift
The first lemma is a geometric statement about the BF-IO assignment at step $k$. Note that its proof does not depend on the specific value of $\delta_{k}$, only on prompt sizes and completion counts.
Lemma 5 (Separation and gap reduction).
Fix $\delta\in\big(0,(2s_{\max}-1)^{-1}\big)$ and set
$\gamma\ :=\ 1-\delta(2s_{\max}-1)\ \in\ (0,1).$
Then, for every $k\geq 1$,
$\mathbb{P}\Big(D(k)\ \leq\ \max\big\{s_{\max},\,A^{\mathrm{pre}}(k)-\gamma Bp\big\}\,\Big|\,\mathcal{F}_{k-1}\Big)\ \geq\ 1-2G\exp\Big(-\tfrac{\delta^{2}}{3}Bp\Big).$ (41)
Proof of Lemma 5.
Fix a step $k$ and condition on $\mathcal{F}_{k-1}$. Given $\mathcal{F}_{k-1}$, the collection of sizes $\{a_{g,i}(k-1)\}$ and loads $\{L_{g}(k-1)\}$ is deterministic, and the randomness at step $k$ comes only from the completion indicators $\{\xi_{g,i}(k)\}$ and the prompt lengths of newly admitted requests. Define
$E_{\mathrm{cnt}}(k)\ :=\ \big\{|c_{g}(k)-Bp|\leq\delta Bp\ \ \text{for all }g\in\{1,\dots,G\}\big\}.$
The event $E_{\mathrm{cnt}}(k)$ depends only on $\{\xi_{g,i}(k)\}$ and thus is independent of $\mathcal{F}_{k-1}$. By a standard Chernoff bound for binomial variables,
$\mathbb{P}\big(E_{\mathrm{cnt}}(k)^{c}\mid\mathcal{F}_{k-1}\big)=\mathbb{P}\big(E_{\mathrm{cnt}}(k)^{c}\big)\ \leq\ 2G\exp\Big(-\tfrac{\delta^{2}}{3}Bp\Big).$
On the event $E_{\mathrm{cnt}}(k)$, we have
$c_{\min}(k)\ :=\ \min_{g}c_{g}(k)\ \geq\ (1-\delta)Bp,\qquad c_{\max}(k)\ :=\ \max_{g}c_{g}(k)\ \leq\ (1+\delta)Bp.$
Fix an outcome $\omega\in E_{\mathrm{cnt}}(k)$ and consider deterministically the BF-IO assignment at step $k$ on that outcome. Let $S_{g}(k)$ denote the multiset of prompt sizes of the new requests admitted to device $g$ at step $k$; by construction $|S_{g}(k)|=c_{g}(k)$, and
$L_{g}(k)\;=\;a_{g}^{\mathrm{pre}}(k)+\sum_{s\in S_{g}(k)}s.$
For each feasible family of multisets $\{S_{g}(k)\}_{g=1}^{G}$ with $|S_{g}(k)|=c_{g}(k)$, define
$L_{g}(k)\ :=\ a_{g}^{\mathrm{pre}}(k)+\sum_{s\in S_{g}(k)}s,\qquad D(k)\ :=\ \max_{g}L_{g}(k)-\min_{g}L_{g}(k).$
Among all such assignments, BF-IO chooses one that minimizes $D(k)$. Let $\{S_{g}^{\star}(k)\}$ be such a minimizer, and let
$L_{g}^{\star}(k)\ :=a_{g}^{\mathrm{pre}}(k)+\sum_{s\in S_{g}^{\star}(k)}s,\qquad D^{\star}(k)\ :=\max_{g}L_{g}^{\star}(k)-\min_{g}L_{g}^{\star}(k),$
denote the corresponding loads and gap. If $D^{\star}(k)\leq s_{\max}$, then $D(k)=D^{\star}(k)\leq s_{\max}$ and the claimed inequality (41) holds trivially on this outcome. Hence it suffices to analyze the case $D^{\star}(k)>s_{\max}$.
Let $p\in\arg\max_{g}L_{g}^{\star}(k)$ and $q\in\arg\min_{g}L_{g}^{\star}(k)$, so that
$L_{p}^{\star}(k)-L_{q}^{\star}(k)\ =\ D^{\star}(k)\ >\ s_{\max}.$
We first show that, under this strict gap and the optimality of $\{S_{g}^{\star}(k)\}$, the assignment must be “separated” between any such heavy–light pair $(p,q)$: there exists an integer $\theta\in\{0,\dots,s_{\max}-1\}$ such that
$s\leq\theta\quad\text{for all }s\in S_{p}^{\star}(k),\qquad s\geq\theta+1\quad\text{for all }s\in S_{q}^{\star}(k).$
Suppose not. Then there exist $x\in S_{p}^{\star}(k)$ and $y\in S_{q}^{\star}(k)$ with $x>y$. Consider the assignment $\{\widetilde{S}_{g}(k)\}$ obtained by swapping $x$ and $y$ between devices $p$ and $q$,
$\widetilde{S}_{p}(k):=(S_{p}^{\star}(k)\setminus\{x\})\cup\{y\},\qquad\widetilde{S}_{q}(k):=(S_{q}^{\star}(k)\setminus\{y\})\cup\{x\},$
and $\widetilde{S}_{r}(k):=S_{r}^{\star}(k)$ for all $r\notin\{p,q\}$. This swap preserves $|S_{g}(k)|=c_{g}(k)$ for every $g$ and uses only jobs from the waiting pool (by the overloaded definition of $\mathcal{I}$, the pool contains sufficiently many jobs of each size to support such exchanges). The new loads are
$\widetilde{L}_{p}(k)\;=\;a_{p}^{\mathrm{pre}}(k)+\sum_{s\in\widetilde{S}_{p}(k)}s\;=\;L_{p}^{\star}(k)-(x-y)\;<\;L_{p}^{\star}(k),$
and
$\widetilde{L}_{q}(k)\;=\;a_{q}^{\mathrm{pre}}(k)+\sum_{s\in\widetilde{S}_{q}(k)}s\;=\;L_{q}^{\star}(k)+(x-y)\;\leq\;L_{q}^{\star}(k)+s_{\max}\;<\;L_{p}^{\star}(k),$
using $x-y\leq s_{\max}$ and $L_{p}^{\star}(k)-L_{q}^{\star}(k)>s_{\max}$. All other loads $\widetilde{L}_{r}(k)$ equal $L_{r}^{\star}(k)\leq L_{p}^{\star}(k)$.
If $p$ is the unique maximizer of $\{L_{g}^{\star}(k)\}$, then
$\max_{g}\widetilde{L}_{g}(k)<\max_{g}L_{g}^{\star}(k),$
contradicting the minimality of $D^{\star}(k)$. If there are multiple maximizers, then the swap strictly reduces the number of devices attaining the maximum load while keeping the maximum value at most $L_{p}^{\star}(k)$. Repeating this argument finitely many times, we either strictly reduce the maximum load or obtain a contradiction to the choice of $\{S_{g}^{\star}(k)\}$ as a minimizer with minimal number of maximizers. Hence no such pair $(x,y)$ can exist, and the desired threshold $\theta$ exists.
Condition on $E_{\mathrm{cnt}}(k)$ and fix an outcome with $D^{\star}(k)>s_{\max}$. Using the separation property, for device $p$ every $s\in S_{p}^{\star}(k)$ satisfies $s\leq\theta$, while for $q$ every $s\in S_{q}^{\star}(k)$ satisfies $s\geq\theta+1$. Therefore,
$\sum_{s\in S_{p}^{\star}(k)}s\;\leq\;\theta\,c_{p}(k),\qquad\sum_{s\in S_{q}^{\star}(k)}s\;\geq\;(\theta+1)\,c_{q}(k),$
and hence
$D^{\star}(k)$ $=\ L_{p}^{\star}(k)-L_{q}^{\star}(k)$
$=\ \bigg(a_{p}^{\mathrm{pre}}(k)+\sum_{s\in S_{p}^{\star}(k)}s\bigg)-\bigg(a_{q}^{\mathrm{pre}}(k)+\sum_{s\in S_{q}^{\star}(k)}s\bigg)$
$\leq\ a_{p}^{\mathrm{pre}}(k)-a_{q}^{\mathrm{pre}}(k)+\theta c_{p}(k)-(\theta+1)c_{q}(k).$
Using $c_{p}(k)\leq c_{\max}(k)$ and $c_{q}(k)\geq c_{\min}(k)$, we obtain
$\theta c_{p}(k)-(\theta+1)c_{q}(k)\;$ $\leq\;\theta c_{\max}(k)-(\theta+1)c_{\min}(k)\;\leq\;\theta(1+\delta)Bp-(\theta+1)(1-\delta)Bp$
$=Bp\big(\theta(1+\delta)-(\theta+1)(1-\delta)\big)=Bp\big(-1+\delta(2\theta+1)\big)$
$\leq Bp\big(-1+\delta(2s_{\max}-1)\big)=-\gamma Bp,$
where we used $\theta\leq s_{\max}-1$ and the definition of $\gamma$. Consequently,
$D^{\star}(k)\leq a_{p}^{\mathrm{pre}}(k)-a_{q}^{\mathrm{pre}}(k)-\gamma Bp\leq A^{\mathrm{pre}}(k)-\gamma Bp.$
Combining this with the trivial bound $D^{\star}(k)\leq s_{\max}$ in the case $D^{\star}(k)\leq s_{\max}$, and recalling that $D(k)=D^{\star}(k)$ under BF-IO, we obtain, for every outcome in $E_{\mathrm{cnt}}(k)$,
[…]
Since $E_{\mathrm{cnt}}(k)$ is independent of $\mathcal{F}_{k-1}$, taking conditional probability given $\mathcal{F}_{k-1}$ yields
$\mathbb{P}\Big(D(k)\ \leq\ \max\big\{s_{\max},\,A^{\mathrm{pre}}(k)-\gamma Bp\big\}\,\Big|\,\mathcal{F}_{k-1}\Big)\;\geq\;\mathbb{P}\big(E_{\mathrm{cnt}}(k)\,\big|\,\mathcal{F}_{k-1}\big)\;\geq\;1-2G\exp\Big(-\tfrac{\delta^{2}}{3}Bp\Big),$
which is (41). ∎
Step 1.B: Concentration of pre-admission spread
We now control $A^{\mathrm{pre}}(k)$ around its conditional baseline (1p)D(k1)(1-p)D(k-1), taking into account the general per-step increments $\{\delta_{k}\}$.
Define the range functional $r(x):=\max_{g}x_{g}-\min_{g}x_{g}$ for $x\in\mathbb{R}^{G}$. For step $k$, recall that
$R_{g}(k)\;=\;\sum_{i=1}^{B}\xi_{g,i}(k)\big(a_{g,i}(k-1)+\delta_{k}\big),\qquad\mu_{g}(k)\ :=\ \mathbb{E}\big[R_{g}(k)\,\big|\,\mathcal{F}_{k-1}\big]=p\sum_{i=1}^{B}(a_{g,i}(k-1)+\delta_{k}),$
and set
$Z_{g}(k)\;:=\ R_{g}(k)-\mu_{g}(k),\qquad Z(k):=(Z_{g}(k))_{g=1}^{G}.$
The one-step envelope
$\Lambda(k):=\max_{g,i}a_{g,i}(k-1)$
[…]
Lemma 6 (One-step concentration of $A^{\mathrm{pre}}(k)$ under general increments).
For each step $k\geq 1$ and every $\mathcal{F}_{k-1}$-measurable $\Lambda(k)$, define
$\alpha_{B}:=3\log(GB),\qquad\widetilde{\Lambda}(k):=\Lambda(k)+\delta_{\max},\qquad u_{A}(k):=C_{0}\,\widetilde{\Lambda}(k)\,\max\Big\{\sqrt{p(1-p)\,B\,\alpha_{B}},\ \alpha_{B}\Big\},$
where $C_{0}>0$ is a universal numerical constant. Then, for all $k\geq 1$,
$\mathbb{P}\Big(A^{\mathrm{pre}}(k)-(1-p)D(k-1)>u_{A}(k)\,\Big|\,\mathcal{F}_{k-1}\Big)\ \leq\ 2G\,e^{-\alpha_{B}}\ \leq\ 2\,G^{-2}B^{-3}.$ (42)
Proof of Lemma 6.
Given $\mathcal{F}_{k-1}$, the vector of loads $L(k-1)=(L_{g}(k-1))_{g=1}^{G}$ is deterministic, and
$a_{g}^{\mathrm{pre}}(k)=L_{g}(k-1)+B\delta_{k}-R_{g}(k)=(1-p)\big(L_{g}(k-1)+B\delta_{k}\big)\;-\;Z_{g}(k).$
Let $X_{g}:=(1-p)(L_{g}(k-1)+B\delta_{k})$ and $X:=(X_{g})_{g=1}^{G}$. Then
$A^{\mathrm{pre}}(k)=r\big(a^{\mathrm{pre}}(k)\big)=r(X-Z(k)).$
For any vectors $x,z\in\mathbb{R}^{G}$, comparing coordinates achieving the maximum and minimum of xzx-z gives the elementary inequality
$r(x-z)\;\leq\;r(x)+2\|z\|_{\infty},$
where $\|z\|_{\infty}:=\max_{g}|z_{g}|$. Applying this with $x=X$ and $z=Z(k)$, and noting that r(X)=(1p)D(k1)r(X)=(1-p)D(k-1) (since adding the common offset $(1-p)B\delta_{k}$ to all coordinates does not change the range), we obtain
$A^{\mathrm{pre}}(k)-(1-p)D(k-1)=r(X-Z(k))-r(X)\ \leq\ 2\|Z(k)\|_{\infty}=2\max_{g}|Z_{g}(k)|.$
Thus, for any $u>0$,
$\mathbb{P}\Big(A^{\mathrm{pre}}(k)-(1-p)D(k-1)>u\,\Big|\,\mathcal{F}_{k-1}\Big)\ \leq\ \sum_{g=1}^{G}\mathbb{P}\Big(|Z_{g}(k)|>u/2\,\Big|\,\mathcal{F}_{k-1}\Big).$
It remains to control each tail $\mathbb{P}(|Z_{g}(k)|>u/2\mid\mathcal{F}_{k-1})$. Conditional on $\mathcal{F}_{k-1}$, the variables
$X_{g,i}(k):=(\xi_{g,i}(k)-p)\big(a_{g,i}(k-1)+\delta_{k}\big),\qquad i=1,\dots,B,$
are independent, mean-zero, and bounded in absolute value by $\widetilde{\Lambda}(k)=\Lambda(k)+\delta_{\max}$. Their conditional variance satisfies
$\mathrm{Var}\Big(\sum_{i=1}^{B}X_{g,i}(k)\,\Big|\,\mathcal{F}_{k-1}\Big)=\sum_{i=1}^{B}p(1-p)\,(a_{g,i}(k-1)+\delta_{k})^{2}\leq p(1-p)\,B\,\widetilde{\Lambda}(k)^{2}.$
By Bernstein’s inequality for bounded summands, there exists a universal constant $C_{1}>0$ such that, for every $g$ and $u>0$,
$\mathbb{P}\Big(|Z_{g}(k)|>u/2\,\Big|\,\mathcal{F}_{k-1}\Big)\ \leq\ 2\exp\left(-C_{1}\,\min\left\{\frac{(u/2)^{2}}{p(1-p)\,B\,\widetilde{\Lambda}(k)^{2}},\ \frac{u/2}{\widetilde{\Lambda}(k)}\right\}\right).$
Summing over $g=1,\dots,G$ and recalling the earlier inequality, we obtain
$\mathbb{P}\Big(A^{\mathrm{pre}}(k)-(1-p)D(k-1)>u\,\Big|\,\mathcal{F}_{k-1}\Big)\ \leq\ 2G\exp\left(-C_{1}\,\min\left\{\frac{(u/2)^{2}}{p(1-p)\,B\,\widetilde{\Lambda}(k)^{2}},\ \frac{u/2}{\widetilde{\Lambda}(k)}\right\}\right).$
Choosing
$u=u_{A}(k)=C_{0}\,\widetilde{\Lambda}(k)\,\max\Big\{\sqrt{p(1-p)\,B\,\alpha_{B}},\ \alpha_{B}\Big\},$
with $C_{0}$ sufficiently large (depending only on $C_{1}$), forces
$\frac{(u/2)^{2}}{p(1-p)\,B\,\widetilde{\Lambda}(k)^{2}}\ \geq\ 2\alpha_{B},\qquad\frac{u}{2\widetilde{\Lambda}(k)}\ \geq\ 2\alpha_{B},$
so that the minimum in the exponent is at least $2\alpha_{B}$. This yields
$\mathbb{P}\Big(A^{\mathrm{pre}}(k)-(1-p)D(k-1)>u_{A}(k)\,\Big|\,\mathcal{F}_{k-1}\Big)\ \leq\ 2Ge^{-\alpha_{B}}.$
With $\alpha_{B}=3\log(GB)$ we have $2Ge^{-\alpha_{B}}\leq 2G^{-2}B^{-3}$, completing the proof. ∎
Step 1.C: Average gap bound for BF-IO under general per-step growth
We now combine Lemmas 5 and 6 to obtain a uniform bound on the long-run average of $D(k)$ when per-step growth is governed by $(\delta_{k})$ so that $0\leq\delta_{k}\leq\delta_{\max}$. As before, at each step the instantaneous imbalance under any policy satisfies
$\mathrm{Imbalance}(k;\textsf{BF-IO})=\sum_{g=1}^{G}\big(\max_{h}L_{h}(k)-L_{g}(k)\big)\ \leq\ (G-1)\Big(\max_{h}L_{h}(k)-\min_{h}L_{h}(k)\Big)=(G-1)\,D(k).$
Lemma 7 (Average gap under geometric outputs with general increments).
Let $\delta:=1/(4s_{\max})$, set $\gamma:=1-\delta(2s_{\max}-1)\in(0,1)$, and let $\alpha_{B}:=3\log(GB)$. Define
$x_{\star}(B,G):=\Big\lceil\frac{3}{p}\,\log(GB)\Big\rceil,\qquad U_{B}:=C_{2}\,(s_{\max}+x_{\star}(B,G)+\delta_{\max})\,\max\Big\{\sqrt{p(1-p)\,B\,\alpha_{B}},\ \alpha_{B}\Big\},$
and
$\varepsilon_{B}:=2G\,e^{-(\delta^{2}/3)Bp}+2\,G^{-2}B^{-3},$
for a universal constant $C_{2}>0$. Assume $B$ is large enough (depending on $(p,s_{\max},\delta_{\max},G)$) so that
$\gamma Bp\ \geq\ 2U_{B}.$ (43)
Then, the BF-IO policy with $H=0$ satisfies
$\limsup_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[D(k)]\ \leq\ \frac{s_{\max}}{p}\ +\ o_{B,G}(1),$ (44)
where $o_{B,G}(1)\geq 0$ depends only on $(B,G,p,s_{\max},\delta_{\max})$ and $o_{B,G}(1)\to 0$ whenever $B,G\to\infty$ with $\sqrt{G}\,\log G=o(B)$.
Proof.
Fix $B,G$, and write $x_{\star}:=x_{\star}(B,G)$. For each step $k$ define the events
$E_{\mathrm{cnt}}(k):=\big\{|c_{g}(k)-Bp|\leq\delta Bp\ \ \forall g\big\},\quad E_{\mathrm{snap}}(k):=\{\Lambda(k)\leq s_{\max}+x_{\star}\},$
and
$E_{\mathrm{dev}}(k):=\big\{A^{\mathrm{pre}}(k)-(1-p)D(k-1)\leq U_{B}\big\}.$
By Lemma 5,
$\mathbb{P}\big(E_{\mathrm{cnt}}(k)^{c}\mid\mathcal{F}_{k-1}\big)\leq 2G\exp\Big(-\tfrac{\delta^{2}}{3}Bp\Big),$
and by Lemma 6 (applied with $\widetilde{\Lambda}(k)=\Lambda(k)+\delta_{\max}$ and using $\Lambda(k)\leq s_{\max}+x_{\star}$ on $E_{\mathrm{snap}}(k)$),
$\mathbb{P}\big(E_{\mathrm{dev}}(k)^{c}\cap E_{\mathrm{snap}}(k)\mid\mathcal{F}_{k-1}\big)\leq 2\,G^{-2}B^{-3},$
provided $C_{2}$ is chosen large enough so that $U_{B}$ dominates the bound $u_{A}(k)$ from Lemma 6 on $E_{\mathrm{snap}}(k)$. Abbreviate
$\varepsilon_{B}:=2G\exp\Big(-\tfrac{\delta^{2}}{3}Bp\Big)+2\,G^{-2}B^{-3}.$
On the “good” event
$E_{\mathrm{good}}(k):=E_{\mathrm{cnt}}(k)\cap E_{\mathrm{snap}}(k)\cap E_{\mathrm{dev}}(k),$
we combine Lemma 5 and the bound on $A^{\mathrm{pre}}(k)$ to obtain
$D(k)\leq\max\Big\{s_{\max},\,(1-p)D(k-1)+U_{B}-\gamma Bp\Big\}.$
Under (43), $\gamma Bp\geq 2U_{B}$, so
$(1-p)D(k-1)+U_{B}-\gamma Bp\leq(1-p)D(k-1)-U_{B}\leq(1-p)D(k-1),$
and hence
$D(k)\ \leq\ s_{\max}+(1-p)D(k-1)\qquad\text{on }E_{\mathrm{good}}(k).$ (45)
On the complement $E_{\mathrm{good}}(k)^{c}$, we use a uniform bound. At the beginning of step $k$, before completions, every alive request grows by $\delta_{k}\leq\delta_{\max}$ and new arrivals have size at most $s_{\max}$, so
$\Lambda(k+1)\leq\max\{\Lambda(k)+\delta_{k},\ s_{\max}\}\leq s_{\max}+\big(\Lambda(k)+\delta_{\max}-s_{\max}\big)_{+}.$
Therefore,
$B\,\Lambda(k+1)\leq B\,(s_{\max}+x_{\star}+\delta_{\max})+B\big(\Lambda(k)+\delta_{\max}-s_{\max}-x_{\star}\big)_{+},$
and since each device holds $B$ requests after admission, we have
$D(k)\leq B\,\Lambda(k+1)\qquad\text{on }E_{\mathrm{good}}(k)^{c}.$
Combining the two cases and using the union bound $\mathbf{1}_{E_{\mathrm{good}}(k)^{c}}\leq\mathbf{1}_{E_{\mathrm{cnt}}(k)^{c}}+\mathbf{1}_{E_{\mathrm{snap}}(k)^{c}}+\mathbf{1}_{E_{\mathrm{dev}}(k)^{c}\cap E_{\mathrm{snap}}(k)}$, we obtain the pathwise decomposition
$D(k)$ $\leq\big(s_{\max}+(1-p)D(k-1)\big)\,\mathbf{1}_{E_{\mathrm{good}}(k)}$
$\quad+B\,(s_{\max}+x_{\star}+\delta_{\max})\big(\mathbf{1}_{E_{\mathrm{cnt}}(k)^{c}}+\mathbf{1}_{E_{\mathrm{dev}}(k)^{c}\cap E_{\mathrm{snap}}(k)}+\mathbf{1}_{E_{\mathrm{snap}}(k)^{c}}\big)$
$\quad+B\big(\Lambda(k)+\delta_{\max}-s_{\max}-x_{\star}\big)_{+}.$
Taking conditional expectations given $\mathcal{F}_{k-1}$ and using the bounds on event probabilities, we have
$\mathbb{E}\big[D(k)\mid\mathcal{F}_{k-1}\big]$ $\leq\big(s_{\max}+(1-p)D(k-1)\big)$
$\quad+\varepsilon_{B}B\,(s_{\max}+x_{\star}+\delta_{\max})+B\,(s_{\max}+x_{\star}+\delta_{\max})\,\mathbb{P}\big(E_{\mathrm{snap}}(k)^{c}\mid\mathcal{F}_{k-1}\big)$
$\quad+B\,\mathbb{E}\big[(\Lambda(k)+\delta_{\max}-s_{\max}-x_{\star})_{+}\mid\mathcal{F}_{k-1}\big].$ (46)
We use the simple bound $\mathbf{1}_{\{\Lambda(k)>s_{\max}+x_{\star}\}}\leq(\Lambda(k)+\delta_{\max}-s_{\max}-x_{\star})_{+}$, and hence
$\mathbb{P}\big(E_{\mathrm{snap}}(k)^{c}\mid\mathcal{F}_{k-1}\big)\leq\mathbb{E}\big[(\Lambda(k)+\delta_{\max}-s_{\max}-x_{\star})_{+}\mid\mathcal{F}_{k-1}\big].$
Substituting into (46) and taking unconditional expectations yields the one-step inequality
$\mathbb{E}[D(k)]$ $\leq s_{\max}+(1-p)\mathbb{E}[D(k-1)]+\varepsilon_{B}B\,(s_{\max}+x_{\star}+\delta_{\max})$
$\quad+B\,(s_{\max}+x_{\star}+\delta_{\max})\,\mathbb{E}\big[(\Lambda(k)+\delta_{\max}-s_{\max}-x_{\star})_{+}\big].$ (47)
To bound the horizon average of the envelope term, observe that in any step $k$, the size of an alive request equals the sum of its prompt length (at most $s_{\max}$) and the cumulative increments $\sum_{r=1}^{\text{age}}\delta_{r}$ over its age. Since $0\leq\delta_{k}\leq\delta_{\max}$, we have
$\text{size}(k)\;\leq\;s_{\max}+\delta_{\max}\cdot\text{age}(k),$
so $\Lambda(k)-s_{\max}$ is bounded by $\delta_{\max}$ times the maximal age among alive requests at step $k$. Summing over the horizon and exchanging the order of summation (first over requests, then over the steps they are alive), we obtain
$\sum_{k=0}^{K-1}\big(\Lambda(k)+\delta_{\max}-s_{\max}-x_{\star}\big)_{+}\ \leq\ \delta_{\max}\sum_{j}\sum_{r=0}^{o_{j}-1}(r+1-y_{\star})_{+},$
where the outer sum is over all requests $j$ whose lifetime intersects [0:K1][0{:}K-1], $o_{j}\sim\mathrm{Geo}(p)$ is the decode length of request $j$, and $y_{\star}:=\lfloor x_{\star}/\delta_{\max}\rfloor$ is an integer threshold on age. For a single geometric random variable $o_{j}\sim\mathrm{Geo}(p)$,
$\mathbb{E}\Big[\sum_{r=0}^{o_{j}-1}(r+1-y_{\star})_{+}\Big]=\sum_{r=y_{\star}+1}^{\infty}(r-y_{\star})\mathbb{P}(o_{j}\geq r)=\sum_{r=y_{\star}+1}^{\infty}(r-y_{\star})(1-p)^{r-1}=\frac{(1-p)^{y_{\star}}}{p^{2}}.$
Let $N_{K}$ be the number of distinct requests whose lifetime intersects the horizon [0:K1][0{:}K-1]. At each step there are exactly $GB$ alive items, and each item counted in $N_{K}$ is alive for at least one step, so deterministically $N_{K}\leq GBK$ and hence $N_{K}/K\leq GB$. By linearity of expectation,
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\big[(\Lambda(k)+\delta_{\max}-s_{\max}-x_{\star})_{+}\big]\leq\frac{\delta_{\max}\,\mathbb{E}[N_{K}]}{K}\cdot\frac{(1-p)^{y_{\star}}}{p^{2}}\leq GB\,\delta_{\max}\,\frac{(1-p)^{y_{\star}}}{p^{2}}.$
Averaging (47) over $k=0,\dots,K-1$ and rearranging, we obtain
$\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[D(k)]$ $\leq s_{\max}+(1-p)\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[D(k-1)]$
$\quad+\varepsilon_{B}B\,(s_{\max}+x_{\star}+\delta_{\max})+\frac{GB^{2}}{p^{2}}\,\delta_{\max}\,(s_{\max}+x_{\star}+\delta_{\max})\,(1-p)^{y_{\star}}.$ (48)
The sequence $\{\mathbb{E}[D(k)]\}_{k\geq 0}$ is nonnegative, so its Cesàro average is bounded, and the right-hand side of (48) is an affine contraction in the horizon average. Letting $K\to\infty$ and using (1p)<1(1-p)<1, we obtain
$\limsup_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[D(k)]\ \leq\ \frac{s_{\max}+\varepsilon_{B}B\,(s_{\max}+x_{\star}+\delta_{\max})+\frac{GB^{2}}{p^{2}}\,\delta_{\max}\,(s_{\max}+x_{\star}+\delta_{\max})\,(1-p)^{y_{\star}}}{p}.$
Finally, for $x_{\star}(B,G)=\lceil 3\log(GB)/p\rceil$, we have $y_{\star}\geq\frac{3}{p\delta_{\max}}\log(GB)$, so
$(1-p)^{y_{\star}}\leq\exp(-py_{\star})\leq(GB)^{-3},$
and thus
$\frac{GB^{2}}{p^{2}}\,\delta_{\max}\,(s_{\max}+x_{\star}+\delta_{\max})\,(1-p)^{y_{\star}}=o_{B,G}(1)$
whenever $B,G\to\infty$ with $\sqrt{G}\,\log G=o(B)$. Similarly, by the definition of $\varepsilon_{B}$ and the same growth condition,
$\varepsilon_{B}B\,(s_{\max}+x_{\star}+\delta_{\max})=o_{B,G}(1).$
Absorbing these vanishing terms into $o_{B,G}(1)$ and dividing by $p$ yields (44). This completes the proof. ∎
Part 2 (Lower bound of Imbalance of FCFS)
We now extend the FCFS lower-bound analysis to the setting with general per-step drift: at each step $k$, every alive request gains a common increment $\delta_{k}\geq 0$, with $\delta_{k}$ independent of $(B,G)$ and of the stochastic primitives. As before, each active request has an output length $o\sim\mathrm{Geo}(p)$, independent across requests and of prompt lengths $S\sim\mathcal{D}_{\mathrm{prefill}}$ supported on $\{0,1,\dots,s_{\max}\}$, with
$\mathbb{E}[S]=\mu_{s},\qquad\mathrm{Var}(S)=\sigma_{s}^{2}>0.$
Step 2.A: Slot-level dynamics under FCFS.
Index slots by $(g,b)$ with $g\in\{1,\dots,G\}$ and $b\in\{1,\dots,B\}$. Let $Y_{g,b}(k)$ denote the post-admission size in slot $(g,b)$ at the end of step $k$. Under FCFS, when the pool is overloaded and size-agnostic, each completion in any slot is immediately refilled by the earliest pending request whose prompt length is an i.i.d. draw from $S$. Thus, for each $(g,b,k)$, we have the recursion
$Y_{g,b}(k)\;=\;(1-\xi_{g,b}(k))\big(Y_{g,b}(k-1)+\delta_{k}\big)+\xi_{g,b}(k)\,S_{g,b}(k),\qquad k\geq 1,$ (49)
where
$\xi_{g,b}(k)\sim\mathrm{Bernoulli}(p),\qquad S_{g,b}(k)\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}S,$
and all $\{\xi_{g,b}(k),S_{g,b}(k)\}_{g,b,k}$ are mutually independent and independent of the initial configuration $\{Y_{g,b}(0)\}$. The completion probability $p$ and the drift $\delta_{k}$ do not depend on the current size $Y_{g,b}(k-1)$.
For device $g$, define the post-admission load and the instantaneous FCFS imbalance at step $k$ by
$S_{g}(k):=\sum_{b=1}^{B}Y_{g,b}(k),\qquad\mathrm{Imbalance}(k;\textsf{FCFS}):=G\cdot\max_{1\leq g\leq G}S_{g}(k)-\sum_{g=1}^{G}S_{g}(k).$
By symmetry of the FCFS rule across devices, for any arrival instance $\mathcal{I}$ and any $k$,
$\mathbb{E}\big[\mathrm{Imbalance}(k;\textsf{FCFS})\big]=G\Big(\mathbb{E}\big[\max_{g}S_{g}(k)\big]-\frac{1}{G}\,\mathbb{E}\big[{\textstyle\sum_{g}S_{g}(k)}\big]\Big).$
We now show that, for every step $k$, the $G$ device loads have fluctuations of order $\sqrt{B}$ with variance uniformly bounded away from zero, and that the maximum across devices therefore exhibits an excess of order $\sqrt{B\log G}$.
Step 2.B: Per-slot decomposition and moment bounds.
Fix a slot $(g,b)$ and suppress its indices. For each step $k$, let $A(k)$ denote the age of the current job in this slot at the end of step $k$: the number of steps elapsed since the last completion in this slot (and since time 0 if no completion has occurred). The completion indicators $\{\xi(k)\}_{k\geq 1}$ form an i.i.d. Bernoulli(p)(p) sequence. Therefore:
  • At the end of step $k$, the event $\{A(k)\geq n\}$ implies that no completion occurred in the last $n$ steps, hence
    $\mathbb{P}\big(A(k)\geq n\big)\;\leq\;(1-p)^{n},\qquad n\geq 0,\;k\geq 1.$ (50)
  • Every time a completion occurs in this slot, the new job’s prompt length is drawn as $S_{j}\sim S$, independently of the completion history and the past prompts. Consequently, at any fixed time $k$, the prompt length $S_{\mathrm{cur}}(k)$ of the job currently in this slot is independent of the completion history up to time $k$, and in particular independent of $A(k)$, with
    $S_{\mathrm{cur}}(k)\ \stackrel{{\scriptstyle d}}{{=}}\ S,\qquad S_{\mathrm{cur}}(k)\perp\!\!\!\perp A(k).$
Given the age $A(k)$ and the sequence $(\delta_{t})$, the total increment accumulated since the last completion is deterministic:
$H_{k}(A(k))\;:=\;\sum_{r=0}^{A(k)-1}\delta_{k-r},$
with the convention that an empty sum is zero when $A(k)=0$. Thus, at the end of step $k$,
$Y(k)\;=\;S_{\mathrm{cur}}(k)+H_{k}\big(A(k)\big),$
and $S_{\mathrm{cur}}(k)$ and $H_{k}(A(k))$ are independent.
From this representation we can immediately extract uniform moment bounds.
Lemma 8 (Per-slot variance and third moment).
For any slot and any step $k\geq 1$,
$\mathrm{Var}\big(Y(k)\big)\;\geq\;\sigma_{s}^{2},$ (51)
and there exists a constant $C_{3}<\infty$ depending only on $(p,\sigma_{s}^{2},\delta_{\max},s_{\max})$ such that
$\mathbb{E}\big[\,\big|Y(k)-\mathbb{E}[Y(k)]\big|^{3}\,\big]\;\leq\;C_{3}\qquad\text{for all }k\geq 1.$ (52)
Proof of Lemma 8.
Because $S_{\mathrm{cur}}(k)$ and $H_{k}(A(k))$ are independent,
$\mathrm{Var}\big(Y(k)\big)=\mathrm{Var}\big(S_{\mathrm{cur}}(k)\big)+\mathrm{Var}\big(H_{k}(A(k))\big)\;\geq\;\mathrm{Var}(S)=\sigma_{s}^{2},$
establishing (51).
For the third moment, we use that $S$ is bounded and that $H_{k}(A(k))$ has uniformly bounded moments. From (50),
$\mathbb{P}\big(A(k)\geq n\big)\;\leq\;(1-p)^{n},\qquad n\geq 0.$
Hence, by the tail-sum formula,
$\mathbb{E}\big[A(k)^{3}\big]\;\leq\;3\sum_{n=0}^{\infty}n^{2}\,\mathbb{P}\big(A(k)\geq n\big)\;\leq\;3\sum_{n=0}^{\infty}n^{2}(1-p)^{n}\;<\;\infty,$
and this bound is uniform in $k$. As $0\leq\delta_{k}\leq\delta_{\max}$, we have $|H_{k}(A(k))|\leq\delta_{\max}A(k)$ and therefore
$\mathbb{E}\big[|H_{k}(A(k))|^{3}\big]\;\leq\;\delta_{\max}^{3}\,\mathbb{E}\big[A(k)^{3}\big]\;\leq\;C^{\prime}_{H},$
for some $C^{\prime}_{H}$ depending only on $(p,\delta_{\max})$. Since $S_{\mathrm{cur}}(k)$ is supported in $\{0,\dots,s_{\max}\}$, its centered third absolute moment is bounded by a constant $C^{\prime}_{S}$ depending only on $s_{\max}$. Writing
$Y(k)-\mathbb{E}[Y(k)]=\big(S_{\mathrm{cur}}(k)-\mu_{s}\big)+\Big(H_{k}(A(k))-\mathbb{E}[H_{k}(A(k))]\Big)+\big(\mu_{s}-\mathbb{E}[Y(k)]\big),$
and using the triangle inequality and the elementary bound $(x+y+z)^{3}\leq 27(x^{3}+y^{3}+z^{3})$ for $x,y,z\geq 0$, we obtain
$\mathbb{E}\big[|Y(k)-\mathbb{E}[Y(k)]|^{3}\big]\;\leq\;C^{\prime\prime}\Big(\mathbb{E}\big[|S_{\mathrm{cur}}(k)-\mu_{s}|^{3}\big]+\mathbb{E}\big[|H_{k}(A(k))|^{3}\big]+1\Big)\;\leq\;C_{3},$
for some finite constant $C_{3}$ depending only on $(p,\sigma_{s}^{2},\delta_{\max},s_{\max})$. This proves (52). ∎
Step 2.C: Berry–Esseen for device-level loads.
Fix a device $g$ and a step $k$. The $B$ slot sizes $\{Y_{g,b}(k)\}_{b=1}^{B}$ are independent (across $b$) by construction: completions $\{\xi_{g,b}(k^{\prime})\}$ and prompts $\{S_{g,b}(k^{\prime})\}$ are independent across slots, and the FCFS assignment is size-agnostic under the overloaded regime. Their marginal distributions may not be identical, but Lemma 8 provides uniform variance and third-moment bounds.
Let
$\mu_{g,b}(k):=\mathbb{E}\big[Y_{g,b}(k)\big],\qquad X_{g,b}(k):=Y_{g,b}(k)-\mu_{g,b}(k),$
and
$V_{g}(k):=\mathrm{Var}\big(S_{g}(k)\big)=\mathrm{Var}\Big(\sum_{b=1}^{B}Y_{g,b}(k)\Big)=\sum_{b=1}^{B}\mathrm{Var}\big(Y_{g,b}(k)\big).$
By (51),
$V_{g}(k)\;\geq\;B\,\sigma_{s}^{2}\qquad\text{for all }g,k.$ (53)
Moreover,
$\sum_{b=1}^{B}\mathbb{E}\big[|X_{g,b}(k)|^{3}\big]\;\leq\;BC_{3}.$
Define the normalized sum
$W_{g}(k)\;:=\;\frac{S_{g}(k)-\mathbb{E}[S_{g}(k)]}{\sqrt{V_{g}(k)}}.$
By the Berry–Esseen theorem for sums of independent, non-identically distributed random variables, there exists an absolute constant $C_{\mathrm{BE}}\in(0,\infty)$ such that, for each $g,k$,
$\sup_{x\in\mathbb{R}}\Big|\mathbb{P}\big(W_{g}(k)\leq x\big)-\Phi(x)\Big|\ \leq\ \frac{C_{\mathrm{BE}}\,\sum_{b=1}^{B}\mathbb{E}[|X_{g,b}(k)|^{3}]}{V_{g}(k)^{3/2}}\ \leq\ \frac{C_{\mathrm{BE}}\,C_{3}}{\sigma_{s}^{3}}\,\frac{1}{\sqrt{B}}\;=:\;\frac{C_{*}}{\sqrt{B}},$ (54)
where $\Phi$ denotes the standard normal distribution function and $C_{*}:=C_{\mathrm{BE}}C_{3}/\sigma_{s}^{3}$ depends only on $(p,\sigma_{s}^{2},\delta_{\max},s_{\max})$.
We now derive a uniform lower bound on the right-tail probabilities of $W_{g}(k)$. Let
$z_{G}:=\sqrt{\frac{\log G}{2}},$
and assume $\log G\geq 4$ (the case of smaller $G$ can be absorbed into constants). For $Z\sim N(0,1)$, the Gaussian tail bound gives
$\mathbb{P}(Z\geq z_{G})\;\geq\;\frac{1}{\sqrt{2\pi}}\Big(\frac{1}{z_{G}}-\frac{1}{z_{G}^{3}}\Big)\exp\Big(-\frac{z_{G}^{2}}{2}\Big)\;\geq\;\frac{1}{2\sqrt{2\pi}\,z_{G}}\,G^{-1/4}$ (55)
for all $\log G\geq 4$. Combining (54) and (55) yields
$\mathbb{P}\big(W_{g}(k)\geq z_{G}\big)\;\geq\;\frac{1}{2\sqrt{2\pi}\,z_{G}}\,G^{-1/4}\;-\;\frac{C_{*}}{\sqrt{B}}.$ (56)
Since we assume $\sqrt{G}=\tilde{o}(B)$, there exists $B_{0}<\infty$ and $G_{0}\geq 2$ such that, for all $G\geq G_{0}$ and all $B\geq B_{0}$ satisfying this growth condition,
$\frac{C_{*}}{\sqrt{B}}\;\leq\;\frac{1}{4\sqrt{2\pi}\,z_{G}}\,G^{-1/4}.$
Hence, for all such $(B,G)$ and for all devices $g$ and steps $k$,
$\mathbb{P}\big(W_{g}(k)\geq z_{G}\big)\;\geq\;\frac{1}{4\sqrt{2\pi}\,z_{G}}\,G^{-1/4}\;=:\;p_{G}.$ (57)
Step 2.D: Extreme-value lower bound across devices.
For a fixed step $k$, the collections $\{Y_{g,b}(k)\}_{b=1}^{B}$ are independent across $g$, because the slot primitives $(\xi_{g,b}(k),S_{g,b}(k))$ are independent across both $g$ and $b$ and FCFS acts independently on each device under the overloaded regime. Therefore, the device loads $\{S_{g}(k)\}_{g=1}^{G}$ are independent (though not necessarily identically distributed). By (57), for each $g$,
$\mathbb{P}\big(W_{g}(k)\geq z_{G}\big)\geq p_{G},$
so by independence across $g$,
$\mathbb{P}\Big(\max_{1\leq g\leq G}W_{g}(k)\geq z_{G}\Big)\;=\;1-\prod_{g=1}^{G}\mathbb{P}\big(W_{g}(k)<z_{G}\big)\;\geq\;1-(1-p_{G})^{G}\;\geq\;1-\exp(-Gp_{G}),$ (58)
where we used $1-x\leq e^{-x}$ for $x\geq 0$. From (57),
$Gp_{G}\;\geq\;\frac{1}{4\sqrt{2\pi}\,z_{G}}\,G^{3/4},$
which is bounded away from zero for every fixed $G\geq 2$. Thus there exists an absolute $c_{1}\in(0,1)$ such that, for all $G\geq 2$ and all $B\geq B_{0}$,
$\mathbb{P}\Big(\max_{1\leq g\leq G}W_{g}(k)\geq z_{G}\Big)\;\geq\;c_{1},\qquad\forall k\geq 1.$ (59)
Next, define
$t_{B,G}:=\sigma_{s}\,\sqrt{B}\,z_{G}.$
By (53), for every $g,k$,
$\sqrt{V_{g}(k)}\;\geq\;\sigma_{s}\sqrt{B},$
so on the event $\{W_{g}(k)\geq z_{G}\}$ we have
$S_{g}(k)-\mathbb{E}[S_{g}(k)]\;=\;W_{g}(k)\,\sqrt{V_{g}(k)}\;\geq\;z_{G}\sqrt{V_{g}(k)}\;\geq\;t_{B,G}.$
Therefore,
$\big\{\max_{1\leq g\leq G}W_{g}(k)\geq z_{G}\big\}\;\subseteq\;\Big\{\max_{1\leq g\leq G}\big(S_{g}(k)-\mathbb{E}[S_{g}(k)]\big)\geq t_{B,G}\Big\}.$
Using the standard inequality $\mathbb{E}[\max_{i}X_{i}]\geq\max_{i}a_{i}+t\,\mathbb{P}(\max_{i}(X_{i}-a_{i})\geq t)$ for any family $\{X_{i}\}$ and any $t>0$, with $X_{i}=S_{i}(k)$ and $a_{i}=\mathbb{E}[S_{i}(k)]$, we obtain
$\mathbb{E}\Big[\max_{1\leq g\leq G}S_{g}(k)\Big]$ $\geq\max_{1\leq g\leq G}\mathbb{E}[S_{g}(k)]+t_{B,G}\,\mathbb{P}\Big(\max_{1\leq g\leq G}\big(S_{g}(k)-\mathbb{E}[S_{g}(k)]\big)\geq t_{B,G}\Big)$
$\geq\max_{1\leq g\leq G}\mathbb{E}[S_{g}(k)]+t_{B,G}\,\mathbb{P}\Big(\max_{1\leq g\leq G}W_{g}(k)\geq z_{G}\Big).$ (60)
Combining (60) with (59), we obtain
$\mathbb{E}\Big[\max_{1\leq g\leq G}S_{g}(k)\Big]\;\geq\;\max_{1\leq g\leq G}\mathbb{E}[S_{g}(k)]+c_{1}\,t_{B,G}=\max_{1\leq g\leq G}\mathbb{E}[S_{g}(k)]+c_{1}\,\sigma_{s}\,\sqrt{B}\,z_{G}.$ (61)
Step 2.E: Lower bound on FCFS imbalance and its time average.
For any step $k$, the expected imbalance under FCFS is
$\mathbb{E}\big[\mathrm{Imbalance}(k;\textsf{FCFS})\big]=G\,\mathbb{E}\Big[\max_{1\leq g\leq G}S_{g}(k)\Big]-\sum_{g=1}^{G}\mathbb{E}[S_{g}(k)].$
Writing
$M_{k}:=\max_{1\leq g\leq G}\mathbb{E}[S_{g}(k)],\qquad\overline{m}_{k}:=\frac{1}{G}\sum_{g=1}^{G}\mathbb{E}[S_{g}(k)],$
we may rewrite
$\mathbb{E}\big[\mathrm{Imbalance}(k;\textsf{FCFS})\big]=G\Big(\mathbb{E}\big[\max_{g}S_{g}(k)\big]-\overline{m}_{k}\Big).$
Using (61) and $M_{k}\geq\overline{m}_{k}$, we obtain
$\mathbb{E}\big[\mathrm{Imbalance}(k;\textsf{FCFS})\big]$ $\geq G\Big(M_{k}+c_{1}\sigma_{s}\sqrt{B}\,z_{G}-\overline{m}_{k}\Big)$
$=G\big(M_{k}-\overline{m}_{k}\big)+c_{1}\,G\,\sigma_{s}\sqrt{B}\,z_{G}$
$\geq c_{1}\,G\,\sigma_{s}\,\sqrt{B}\,z_{G}.$ (62)
Recalling $z_{G}=\sqrt{(\log G)/2}$, we conclude that there exists a constant $c>0$, depending only on $(p,\sigma_{s}^{2},\delta_{\max},s_{\max})$, such that, for all $G\geq 2$ and all $B$ sufficiently large with $\sqrt{G}=\tilde{o}(B)$,
$\mathbb{E}\big[\mathrm{Imbalance}(k;\textsf{FCFS})\big]\;\geq\;c\,G\,\sigma_{s}\,\sqrt{B\log G},\qquad\forall k\geq 1.$ (63)
Define the long-run expected time-average imbalance under FCFS and arrival instance $\mathcal{I}$ by
$\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I}):=\liminf_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\big[\mathrm{Imbalance}(k;\textsf{FCFS})\big].$
Since the lower bound (63) holds uniformly for all $k\geq 1$, we have
$\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I})\;\geq\;c\,G\,\sigma_{s}\,\sqrt{B\log G},$
uniformly over all overloaded arrival instances $\mathcal{I}$.
In particular, as in the homogeneous-drift case, FCFS incurs imbalance of order $G\sqrt{B\log G}$, with the constant depending only on the prompt-length dispersion $\sigma_{s}^{2}$, the geometric parameter $p$, and the drift bound $\delta_{\max}$, but not on $(B,G)$.
Part 3: BF-IO vs. FCFS
We now combine the BF-IO average-gap bound from Lemma 7 with the FCFS lower bound from Part 2 to obtain the imbalance–reduction rate under geometric outputs with general per-step growth.
Recall the definition of the long-run expected average imbalance for a policy $\pi\in\{\textsf{BF-IO},\textsf{FCFS}\}$ and an arrival instance $\mathcal{I}$:
$\mathrm{AvgImbalance}(\pi;\mathcal{I}):=\limsup_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\big[\mathrm{Imbalance}(k;\pi)\big].$
As in the homogeneous-output case, we define the imbalance–reduction ratio of BF-IO against FCFS by
$\mathbf{IIR}\;:=\;\inf_{\mathcal{I}}\;\frac{\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I})}{\mathrm{AvgImbalance}(\textsf{BF-IO};\mathcal{I})},$
where the infimum ranges over all overloaded arrival instances $\mathcal{I}$.
From Lemma 7, under our growth assumption $\sqrt{G}=\tilde{o}(B)$ we have, for any overloaded arrival instance $\mathcal{I}$,
$\limsup_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}[D(k)]\ \leq\ \frac{s_{\max}}{p}\ +\ o_{B,G}(1),$
where $o_{B,G}(1)\to 0$ as $B,G\to\infty$ with $\sqrt{G}=\tilde{o}(B)$. Since for every step $k$,
$\mathrm{Imbalance}(k;\textsf{BF-IO})=\sum_{g=1}^{G}\big(\max_{h}L_{h}(k)-L_{g}(k)\big)\ \leq\ (G-1)\,D(k),$
we obtain
$\mathrm{AvgImbalance}(\textsf{BF-IO};\mathcal{I})\ =\ \limsup_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\mathbb{E}\big[\mathrm{Imbalance}(k;\textsf{BF-IO})\big]\ \leq\ (G-1)\,\frac{s_{\max}}{p}\ +\ (G-1)\,o_{B,G}(1).$ (64)
On the other hand, from the FCFS analysis in Part 2 we showed that, for all $G\geq 2$ and all $B$ sufficiently large satisfying $\sqrt{G}\,\log G=o(B)$, there exists a constant $c_{\mathrm{F}}>0$ (depending only on $\sigma_{s}^{2},p,\delta_{\max},s_{\max}$) such that, uniformly over all overloaded arrival instances $\mathcal{I}$,
$\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I})\ \geq\ c_{\mathrm{F}}\,G\,\sigma_{s}\,\sqrt{B\log G}.$ (65)
Combining (64) and (65), we obtain, for every such $\mathcal{I}$,
$\frac{\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I})}{\mathrm{AvgImbalance}(\textsf{BF-IO};\mathcal{I})}$ $\ \geq\ \frac{c_{\mathrm{F}}\,G\,\sigma_{s}\,\sqrt{B\log G}}{(G-1)\,\frac{s_{\max}}{p}+(G-1)\,o_{B,G}(1)}$
$\ =\ \Bigg(\frac{c_{\mathrm{F}}\,p\,\sigma_{s}}{s_{\max}}\,\frac{G}{G-1}\Bigg)\,\frac{\sqrt{B\log G}}{1+\frac{p}{s_{\max}}\,o_{B,G}(1)}.$
Since $o_{B,G}(1)\to 0$, we can choose $B$ large enough (as a function of $G$) so that $\frac{p}{s_{\max}}\,o_{B,G}(1)\leq\tfrac{1}{2}$. For all such $(B,G)$,
$\frac{1}{1+\frac{p}{s_{\max}}\,o_{B,G}(1)}\;\geq\;\frac{2}{3},$
and we obtain the uniform lower bound
$\frac{\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I})}{\mathrm{AvgImbalance}(\textsf{BF-IO};\mathcal{I})}\ \geq\ c\,\frac{p\,\sigma_{s}}{s_{\max}}\,\frac{G}{G-1}\,\sqrt{B\log G},$
for some universal constant $c>0$ depending only on $(\sigma_{s}^{2},p,\delta_{\max},s_{\max})$. Taking the infimum over all overloaded arrival instances $\mathcal{I}$ yields
$\mathbf{IIR}\;=\;\inf_{\mathcal{I}}\;\frac{\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I})}{\mathrm{AvgImbalance}(\textsf{BF-IO};\mathcal{I})}\;\geq\;c\,\frac{p\,\sigma_{s}}{s_{\max}}\,\frac{G}{G-1}\,\sqrt{B\log G}\;=\;\Omega\!\big(\sqrt{B\log G}\big),$
which shows that BF-IO achieves the same $\sqrt{B\log G}$-order improvement over FCFS in the general per-step drift setting as in the homogeneous +1+1 model.
C.4 Proof of Theorem 4
Proof of Theorem 4.
Fix $\pi$ and $\mathcal{I}$. For each step $k$, recall $u_{g}(k)=L_{g}(k)/L_{g^{*}}(k)$. By (10),
$E(\pi;\mathcal{I})$ $=\kappa_{\mathrm{ATT}}\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}L_{g^{*}}(k)\sum_{g=1}^{G}\Bigl(P_{\mathrm{idle}}+(P_{\mathrm{max}}-P_{\mathrm{idle}})\,u_{g}(k)^{\gamma}\Bigr)$
$=\kappa_{\mathrm{ATT}}\sum_{k,g}\Bigl(P_{\mathrm{idle}}L_{g^{*}}(k)+(P_{\mathrm{max}}-P_{\mathrm{idle}})L_{g^{*}}(k)u_{g}(k)\Bigr)\;$ (66)
$+\;\kappa_{\mathrm{ATT}}(P_{\mathrm{max}}-P_{\mathrm{idle}})\sum_{k,g}L_{g^{*}}(k)\bigl(u_{g}(k)^{\gamma}-u_{g}(k)\bigr),$ (67)
where $\sum_{k,g}$ abbreviates $\sum_{k=0}^{K_{\pi}(\mathcal{I})-1}\sum_{g=1}^{G}$. Since $L_{g^{*}}(k)u_{g}(k)=L_{g}(k)$, the first sum in (67) becomes
$\kappa_{\mathrm{ATT}}\sum_{k,g}\Bigl(P_{\mathrm{idle}}L_{g^{*}}(k)+(P_{\mathrm{max}}-P_{\mathrm{idle}})L_{g}(k)\Bigr)$ $=\kappa_{\mathrm{ATT}}\sum_{k,g}\Bigl(P_{\mathrm{max}}L_{g}(k)+P_{\mathrm{idle}}(L_{g^{*}}(k)-L_{g}(k))\Bigr)$
$=\kappa_{\mathrm{ATT}}P_{\mathrm{max}}W(\mathcal{I})+\kappa_{\mathrm{ATT}}P_{\mathrm{idle}}\mathrm{ImbTot}(\pi;\mathcal{I}),$ (68)
using (11) and (12). Substituting (68) into (67) yields the exact identity
$E(\pi;\mathcal{I})=\kappa_{\mathrm{ATT}}P_{\mathrm{max}}W(\mathcal{I})+\kappa_{\mathrm{ATT}}P_{\mathrm{idle}}\mathrm{ImbTot}(\pi;\mathcal{I})+\kappa_{\mathrm{ATT}}(P_{\mathrm{max}}-P_{\mathrm{idle}})\sum_{k,g}L_{g^{*}}(k)\bigl(u_{g}(k)^{\gamma}-u_{g}(k)\bigr).$ (69)
We next bound the last term in (69). Since $\gamma\in(0,1)$ and $u\in[0,1]$, we have $u^{\gamma}\geq u$, hence the last term is nonnegative. For an upper bound, concavity of $u^{\gamma}$ on $[0,1]$ implies the tangent inequality at $u=1$:
$u^{\gamma}\;\leq\;1+\gamma(u-1)\;=\;(1-\gamma)+\gamma u,$
so $u^{\gamma}-u\leq(1-\gamma)(1-u)$. Therefore,
$L_{g^{*}}(k)\bigl(u_{g}(k)^{\gamma}-u_{g}(k)\bigr)\;\leq\;(1-\gamma)\,L_{g^{*}}(k)\bigl(1-u_{g}(k)\bigr)=(1-\gamma)\bigl(L_{g^{*}}(k)-L_{g}(k)\bigr).$
Summing over $k,g$ and using (12),
[…]
Plugging (70) into (69) yields the sandwich bound
$\kappa_{\mathrm{ATT}}P_{\mathrm{max}}W(\mathcal{I})+\kappa_{\mathrm{ATT}}P_{\mathrm{idle}}\mathrm{ImbTot}(\pi;\mathcal{I})\ \leq\ E(\pi;\mathcal{I})\ \leq\ \kappa_{\mathrm{ATT}}P_{\mathrm{max}}W(\mathcal{I})+\kappa_{\mathrm{ATT}}C_{\gamma}\,\mathrm{ImbTot}(\pi;\mathcal{I}),$ (71)
where $C_{\gamma}$ is defined in (15).
Now apply (71) with $\pi=\pi_{0}$ and $\pi=\pi_{1}$. Using the upper bound for $E(\pi_{0};\mathcal{I})$ and the lower bound for $E(\pi_{1};\mathcal{I})$ gives
$E(\pi_{0};\mathcal{I})\ \leq\ \kappa_{\mathrm{ATT}}W(\mathcal{I})\Bigl(P_{\mathrm{max}}+C_{\gamma}\,\eta_{\mathrm{sum}}(\pi_{0};\mathcal{I})\Bigr),$ (72)
where we used $\mathrm{ImbTot}(\pi_{0};\mathcal{I})=\eta_{\mathrm{sum}}(\pi_{0};\mathcal{I})W(\mathcal{I})$. For the energy difference, use (69) and (70) to obtain
$E(\pi_{0};\mathcal{I})-E(\pi_{1};\mathcal{I})$ $=\kappa_{\mathrm{ATT}}P_{\mathrm{idle}}\bigl(\mathrm{ImbTot}(\pi_{0};\mathcal{I})-\mathrm{ImbTot}(\pi_{1};\mathcal{I})\bigr)$
$+\kappa_{\mathrm{ATT}}(P_{\mathrm{max}}-P_{\mathrm{idle}})\bigl(X(\pi_{0})-X(\pi_{1})\bigr),$ (73)
where $X(\pi):=\sum_{k,g}L_{g^{*}}(k)\bigl(u_{g}(k)^{\gamma}-u_{g}(k)\bigr)$. By (70), $X(\pi_{0})\geq 0$ and $X(\pi_{1})\leq(1-\gamma)\mathrm{ImbTot}(\pi_{1};\mathcal{I})$. Hence
$E(\pi_{0};\mathcal{I})-E(\pi_{1};\mathcal{I})$ $\geq\kappa_{\mathrm{ATT}}P_{\mathrm{idle}}\bigl(\mathrm{ImbTot}(\pi_{0};\mathcal{I})-\mathrm{ImbTot}(\pi_{1};\mathcal{I})\bigr)$
$-\kappa_{\mathrm{ATT}}(P_{\mathrm{max}}-P_{\mathrm{idle}})(1-\gamma)\,\mathrm{ImbTot}(\pi_{1};\mathcal{I})$
$=\kappa_{\mathrm{ATT}}\left(P_{\mathrm{idle}}\bigl(\mathrm{ImbTot}(\pi_{0};\mathcal{I})-\mathrm{ImbTot}(\pi_{1};\mathcal{I})\bigr)-D_{\gamma}\,\mathrm{ImbTot}(\pi_{1};\mathcal{I})\right),$ (74)
where $D_{\gamma}$ is defined in (15). Under the assumption (14), $\mathrm{ImbTot}(\pi_{1};\mathcal{I})\leq\mathrm{ImbTot}(\pi_{0};\mathcal{I})/\alpha$, so (74) implies
$E(\pi_{0};\mathcal{I})-E(\pi_{1};\mathcal{I})\ \geq\ \kappa_{\mathrm{ATT}}\mathrm{ImbTot}(\pi_{0};\mathcal{I})\left(P_{\mathrm{idle}}\Bigl(1-\frac{1}{\alpha}\Bigr)-\frac{D_{\gamma}}{\alpha}\right).$ (75)
Combining (75) with (72) and using $\mathbb{E}\bigl[\mathrm{ImbTot}(\pi_{0};\mathcal{I})\bigr]=\eta_{\mathrm{sum}}(\pi_{0};\mathcal{I})\mathbb{E}\bigl[W(\mathcal{I})\bigr]$ gives (16). Taking the expectation completes the proof. ∎
C.5 Proof of Inequality (17)
Proof of Inequality (17).
Under FCFS in the overloaded regime, the slot-level Markov recursion (36) (Section 5) is uniformly ergodic with unique stationary slot law $U\overset{d}{=}S_{\mathrm{cur}}+A$ where $S_{\mathrm{cur}}\stackrel{{\scriptstyle d}}{{=}}S$ and $A\sim\mathrm{Geo}(p)-1$ independent, hence
$\mu_{U}:=\mathbb{E}[U]=\mu_{s}+\frac{1-p}{p},\qquad\mathrm{Var}(U)=\sigma_{\mathrm{snap}}^{2}.$
At stationarity, the total system workload in a step equals $\sum_{g=1}^{G}\sum_{b=1}^{B}U_{g,b}$, so its expectation is $GB\mu_{U}$. Therefore the long-run average total workload per step under FCFS equals $GB\mu_{U}$ by uniform ergodicity and Cesàro averaging. On the other hand, Part 2 of Appendix C.2 establishes the lower bound
$\mathbb{E}\bigl[\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I})\bigr]\ \geq\ c^{\prime}\,G\,\sigma_{\mathrm{snap}}\sqrt{B\log G}$
for a universal constant $c^{\prime}>0$ and all $B$ sufficiently large. Dividing by $GB\mu_{U}$ yields
$\frac{\mathbb{E}\bigl[\mathrm{AvgImbalance}(\textsf{FCFS};\mathcal{I})\bigr]}{GB\mu_{U}}\ \geq\ c^{\prime}\,\frac{\sigma_{\mathrm{snap}}}{\mu_{U}}\sqrt{\frac{\log G}{B}}.$
Interpreting $\eta_{\mathrm{sum}}(\textsf{FCFS};\mathcal{I})$ as the corresponding normalized imbalance level (cumulative or long-run averaged; these coincide under stationarity), we obtain (17) (absorbing constants into $\gtrsim$). ∎
Appendix D Supplementary Materials for Section 6
D.1 Energy and MFU Computation
This appendix provides details on the energy consumption metric used in our experiments.
Model FLOPs Utilization (MFU).
MFU measures how efficiently the GPU’s computational capacity is utilized, defined as the ratio of actual floating-point operations performed to the theoretical peak:
$\mathrm{mfu}=\frac{\text{Observed FLOPs/second}}{\text{Peak FLOPs/second}}$ (76)
For LLM inference, the FLOPs required to generate one token is approximately $6\times N_{\mathrm{params}}$, arising from the matrix multiplications in the forward pass (approximately $2\times N_{\mathrm{params}}$ multiply-accumulate operations, each counting as 2 FLOPs, with an additional factor from attention computation). Thus, for a throughput of $\mathcal{T}$ tokens per second on a model with $N_{\mathrm{params}}$ parameters:
$\mathrm{mfu}\approx\frac{\mathcal{T}\times 6\times N_{\mathrm{params}}}{\mathrm{FLOPs}_{\mathrm{peak}}}$ (77)
where $\mathrm{FLOPs}_{\mathrm{peak}}$ is the GPU’s theoretical peak throughput (e.g., 312 TFLOPs for NVIDIA A100 at FP16/BF16 precision).
Power Model and Parameters.
We adopt the power–utilization model from [ozcan2025quantifying]:
$P(\mathrm{mfu})=P_{\mathrm{idle}}+(P_{\mathrm{max}}-P_{\mathrm{idle}})\cdot\left(\frac{\mathrm{mfu}}{\mathrm{mfu}_{\mathrm{sat}}}\right)^{\gamma}$ (78)
The parameters are set as follows:
  • $P_{\mathrm{idle}}=100$W: baseline power draw when the GPU is powered but not computing, reflecting memory refresh, control logic, and leakage current.
  • $P_{\mathrm{max}}=400$W: peak power draw at full utilization, consistent with NVIDIA A100/H100 thermal design power specifications.
  • $\mathrm{mfu}_{\mathrm{sat}}=0.45$: the utilization level at which power consumption saturates. In practice, inference workloads rarely exceed this threshold due to memory bandwidth limitations.
  • $\gamma=0.7$: the sublinear scaling exponent, empirically calibrated in [ozcan2025quantifying] for inference workloads. The value $\gamma<1$ implies diminishing power increase per unit utilization gain, but also substantial power draw ($P_{\mathrm{idle}}$) even at zero useful throughput.
Implications for Load Balancing.
The sublinear power model has a direct implication for barrier-synchronized systems. At each decode step, the wall-clock duration is $\Delta t=C+t_{\ell}\cdot\max_{g}L_{g}$. Workers with load $L_{g}<\max_{g^{\prime}}L_{g^{\prime}}$ complete their local computation early and idle until the barrier clears. During this idle period, these workers draw power near $P_{\mathrm{idle}}$ while generating zero tokens. Therefore, the total energy consumed is the discrete approximation of the integral (Equation (6)) :
$E=\sum_{k=1}^{K}\sum_{g=1}^{G}\left[P\bigl(\mathrm{mfu}_{g}^{(k)}\bigr)\cdot\Delta t^{(k)}\right]$ (79)
where $\mathrm{mfu}_{g}^{(k)}$ reflects worker $g$’s effective utilization at step $k$. Workers idling at barriers have $\mathrm{mfu}_{g}^{(k)}\approx 0$ for a fraction of the step, drawing $P_{\mathrm{idle}}$ without contributing to throughput. By reducing load imbalance, BF-IO ensures that all workers remain productively utilized throughout each step, converting idle power into useful computation and thereby reducing total energy consumption.
D.2 Results on the Lighter-Load Dataset BurstGPT
We additionally evaluate our routing policy on BurstGPT [wang2024burstgpt], a lighter-load benchmark with shorter request lengths than LongBench. The prefill and decode length distributions are shown in Figure 12. Given the reduced per-request workload, 16 GPUs suffice to process this dataset.
Figure 12: Distribution of prefill (input) and decode (output) lengths in the BurstGPT dataset.
Load Trajectory Analysis.
Figure 13 presents per-worker load trajectories for the BurstGPT workload. Under FCFS, worker loads spread across 250k–400k tokens, with persistent inter-worker variance throughout execution. JSQ exhibits even greater imbalance, confirming that queue length is a poor proxy for actual token load when request sizes vary substantially. In contrast, BF-IO ($H=0$) achieves tighter clustering around 270k–320k tokens. With lookahead enabled, BF-IO ($H=20$) produces the most balanced distribution: all workers maintain nearly identical loads throughout execution, reducing average imbalance by a factor of $17\times$ compared to FCFS.
Figure 13: Per-worker token load trajectories under four routing policies (BurstGPT, $G=16$ GPUs). FCFS exhibits persistent load spread (250k–400k tokens); JSQ shows even greater imbalance due to its size-agnostic assignment. BF-IO ($H=0$) achieves substantially tighter clustering, and BF-IO ($H=20$) produces near-uniform loads across all workers throughout execution.
Power Consumption Analysis.
Figure 14 illustrates average GPU power consumption over time. The baseline FCFS policy oscillates between 330–380 W, reflecting periods in which subsets of GPUs remain underutilized. BF-IO ($H=20$) maintains more stable power draw at higher sustained utilization, completing the workload faster despite drawing similar peak power—consistent with the energy-saving mechanism established in the main text.
Figure 14: Average GPU power consumption over time (BurstGPT, $G=16$ GPUs). FCFS exhibits fluctuations (330–380 W) due to load imbalance, while BF-IO maintains higher sustained utilization and completes the workload in less time, yielding lower total energy consumption.
Quantitative Metrics Comparison.
Table 2 summarizes performance across routing policies. Even in this lighter-load regime, BF-IO ($H=20$) reduces average imbalance by $17\times$ relative to FCFS, improves throughput by 14%, reduces TPOT by 12%, and lowers energy consumption by 3.3%. The more modest energy savings compared to LongBench (3.3% vs. 29%) reflect the lower baseline imbalance in this easier workload; nevertheless, the improvements remain consistent and statistically significant across all metrics.
Notably, FCFS and JSQ exhibit nearly identical performance on this dataset, confirming that both policies are effectively size-agnostic: JSQ balances queue lengths rather than token loads, providing no advantage when request sizes vary.
Table 2: Performance comparison on BurstGPT ($G=16$ GPUs, $B=72$ batch size). Arrows indicate preferred direction.
[…]
($\times 10^{4}$) ($10^{4}$ tok/s) ($10^{-2}$ s/tok) (kJ)
FCFS 27.9 8.00 1.42 396
JSQ 28.2 7.99 1.42 396
BF-IO ($H=0$) 2.92 9.03 1.26 386
BF-IO ($H=20$) 1.65 9.13 1.25 383